This page: http://monitor.mikes.fi/ionclock/

TOC/References - Notes - Comb - Ion - Cavity - DAQ

Archived notes: index_2019 index_2020 index_2021

News and notable results

  • 2022-06-18: New OPA818-based PDH photodetector with lower (shot noise limited?) noise
  • 2022-05-21: first search for Autler-Townes splitting
  • 2022-05-02: first AOM chirp measurement (measurements 2022-04-28/29)
  • 2022-03-04 (59642) to 2022-04-03 (59672) ROCIT-campaign
  • 2022-01-09: coil-characterization
  • 2021-12-21: secular frequencies vs. bias voltages
  • 2021-10-28: clock-laser frequency with new 250 MHz Menlo comb
  • 2021-04-30: trapping again after IR-camera BBR-characterization. New machined alu C-beam mount
  • 2021-01-25: first ions trapped after x-mas break. Intensifier+PCO needed for beam-alignment
  • 2021-01-22: new mu-shields delivered. inner and mid-shields installed and fit ok
  • 2020-12-03: Initial measurement of repumper-shift: -2600 (C1) to -2900 Hz (C4)
  • 2020-11-18: New Ion page for clock-run data
  • 2020-11-04: clock run with new HelRes (now 14.424 MHz) and new B-beam direction, but only 800us probe
  • 2020-09-24: Move 2019 results to index_2019
  • 2020-09-24: Secular frequencies with tickler + photon-correlation
  • 2020-09-23: new 14.404 MHz Magic-RF Helical-Resonator
  • 2020-07-02: Fourier limited 4 Hz linewidth with 180ms probe-pulse. Night, no metro.
  • 2020-06-11: Second clock run, OD2 with 4ms probe, only C1, C2 peaks.
  • 2020-05-17: D5/2 lifetime, random-telegraph method with 405nm and fit
  • 2020-04-24: First clock run! 2.5ms probe, stability 2e-13/sqrt(tau_s)
  • 2020-04-01: trapping single ion again, after 405 laser broke
  • 2020-01-31: Move comb-plots to comb
  • 2020-01-15: For temperatures, see DAQ
  • 2019-12-19: Micromotion map, 3D pictures
  • 2019-11-29: Cavity mode number, Comb mode number. Cavity absolute frequency.
  • 2019-11-08: -C2 Zeeman component linewidth estimate of 800 Hz.
  • 2019-11-05: Zeeman peaks with B-shield (1.6 uT estimate).
  • 2019-11-04: Now with first B-shield. Laminated plywood + metglas.
  • 2019-07-12: Zeeman peaks in ambient (66uT estimate) B-field. First clock-spectrum!
  • 2019-07-10: Initial clock-transition search...
  • 2019-07-10: Page created.
  • 2019-06-18: ion photo and secular frequencies

  • 2022-08-12

    New AOM-chirp measurement setup, now with mini-circ programmable attenuator after the AOM PA.
    Old PDH-detector was modified with 3k9 transimpedance to give bandwidth suitable for measuring double-pass beat note at 150 MHz
    Setup is similar to 2022-04 (2022-04-28) chirp-measurement, with added adjustable mirror to align beam onto detector
    PDH-detector is followed by mini-circ amplifier (ca +23 dB) and splitter to SDR and Siglent-SA.
    SNR for beat-note with maximal AOM-drive is >65 dB - clear improvement from old setup with 60MHz detector
    142 Hz sideband @ -51 dBc
    1.04 kHz sideband @ -44 dBc (this could be mechanical resonance of some component - sideband seems to jump up when hitting on the trap-breadboard?)

    2022-08-08

    Note 0.4 nW floor of power-meter restricts measurable range to ca 0-20 dB attenuation.
    Measured power seems to agree with model, using optimal RF-power P0 = 1.3*Pmax_used
    single-pass model is I/I0 = sin**2 [ (pi/2) * sqrt(P/P0) ]
    
    small-angle approximation sin(x) ~ x for small x.
    -> I/I0 ~ P single-pass, and I/I0 ~ P**2 double-pass
    
    When this scaling holds, we expect 10 dB of reduction in optical power (OD1-filter) for each 5 dB of reduced RF power.
    
    

    With RCDAT inserted after PA, 0dB setting, now ARTIQ attenuation ca 15 dB seems OK.
    
    RCDAT  P674
    0      3.8 uW
    3      1.78
    6      605 uW
    9      177 nW
    10     112
    11     71.8
    12     49
    13     31
    14     19.8 nW
    15     12.9
    16     9.1
    17     6.0
    18     4.0
    19     2.6
    20     1.9
    21     1.3
    22     1.0
    23     0.8
    24     0.6
    30     0.4
    40     0.4
    50     0.4
    off    0.4  (artiq DDS off)
    
    Test 674nm power vs. AOM RF drive power, old configuration (no adjustable attenuator)
    674nm power-servo set-point to 2.0 V, lab lights dark, detector-area covered with holy-blanket
    Thorlabs slim detector after 461-cube, before B-fiber
    
    Detector dark: 0.2 nW
    18 dB DDS att   4.03 uW
    19  3.45 uW
    20  2.75
    21  2.15
    22  1.58
    23  1.13
    24  0.77
    25  0.53
    26  0.36
    27  0.24
    28  0.16
    29 105 nW
    30 67 nW
    DDS off 0.5 nW
    18 4.00 uW
    17 4.35
    16 4.42
    15 4.24
    14 4.08
    
    AA Optoelectronics whitepapre on diffraction efficiency:
    I1/I0 = sin**2 [ (pi/2) * sqrt(P/P0) ]
    this formula for single-pass, we use double pass.
    Note RF PA may go into saturation at high power-levels!
    
    


    Earlier measurement of the RCDAT programmable attenuator: RCDAT-3000-63W2
    Measured over typical AOM frequency range of 60-90 MHz.

    Lowest dataset is SA floor

    2022-08-03

    Migration to ARTIQ-7 done.

    2022-08-02

    Starting migration to ARTIQ-7

    Summer break 2022


    2022-07-01

    09:37 ion still stored
    uMotion check: 1.4%, 1.0%, 0.3%
    
    Optical disturbance from Comb-computer screen at 60Hz is significantly less when brightness set to 100%
    Some 50 Hz optical disturbance still from first room emergency-light and normal lab-lights
    
    Drive fiber-noise AOM with fixed 178.5 MHz requency from R&S generator
    this allows turning off LDS12B PSU that was powering fiber-noise detector
    (but what could be sensitive to AC magnetic field in the cavity or PDH-lock?)
    
    Check fiber-coupling of 422nm laser to ion breadboard, only ca 2-3 uW in B-beam before B-beam fiber, with slider pos1 (no filter)
    
    



    50ms spectra with fiber-noise PSU off (no fiber noise cancellation)
    stop scan during +C4

    ion went dark when measuring with magcheck - 422nm rb-unlock, relocked
    ca14 ion lost?

    With Rigol scole, AC line trigger, averaging mode, Magcheck shows ca 1 mVpp on top of outer shield
    Try turning off trap bias-voltage and coil-current PSU? Maybe smoother waveform, but still 1 mVpp (triangular shaped, not sine)
    With top-shelf Siglent PSU, magcheck reading reduced to 0.4 mVpp

    2022-06-30

    ion dark
    stop clock-run, all beams on, clearout on
    ion recovers
    
    uMotion: 1.7% 2.8% 0.4%
    
    


    ID23752 peak-search, 50ms:


    Comb control software crashed - out of memory

    
    Measure AC-magnetic field with Magcheck-95e probe
    Sensitivity 1 mV AC = 1 mG = 0.001 * 0.1 mT = 0.1 uT 
    (1 G = 1e-4 T = 0.1 mT )
    Sensor vertically on top of outer shield, above ion
    Horizontal orientations give significantly smaller measurements, so most of AC B-field vertical?
    Measure with Rigol RF-monitor scope, DC-coupling, 1 MOhm
    
    7 mVpp with Thorlabs LDS1212 on (this caused severe sidebands previoulsy)
    3.2 mVpp with LDS1212 off
    no visible change with PMT off/on
    123 mVpp when sensor placed at back of BOA laser-driver (Thorlabs ITC510)
    3.8 mVpp with sensor on top of outer shield, helres end
    7.8 mVpp top of outer shield, ion-pump end
    1.4 mVpp with sensor on trap-breadboard, next to helres mount
    1.98 mVpp measured with black Keysight scope (sensor on top shield above ion)
    1.96 mVpp with Rigol scope off
    0.56 mVpp main optical table, next to helres mount
    0.11 mVpp lab floor
    1.9 mVpp back to top of outer shield, above ion
    disconnecting 11mA I2 current has no effect
    disconnecting brick PSU for 461nm doubler temperature-controller -> no effect
    5.7 mVpp with LDS1212 on
    
    mis-aligned B-beam while playing with magcheck sensor??
    14:46 after some searching, found B-beam alignment again.
    
    17:20 check uMotion: 0.97%, 2.1%, 1.9%
    Fluorescence-level seems low, but this is because 422nm power is low. ca 2 uW in A-beam befre fiber (usually 10% of this in B-beam after objective, towards ion)
    
    
    Magcheck on top of outer shield, at ion position:
    Normal situation:
    LDS1212 on, distance to sensor ca 1m: Note ca 3-fold larger amplitude
    ca 39 Vpp when magcheck placed on top of LDS1212!! (0.8 mVpp when LDS1212 is off)
    Magcheck on lab floor, close to ion-trap table-leg: ca 0.16 mVpp, 10-fold smaller compared to top of outer shield
    Similar reading on trap-breadboard (back left corner), close to ion-pump wire (dispenser wire-terminal): 0.17 mVpp
    Close to PMT on upper breadbord, even better value 0.08 mVpp
    2.2 mVpp on top of B-beam post (similar height to top of outer shield)
    Top of B-beam post: 1.5 mVpp with BOA current zero (ITC510 on), 2.2 mVpp with BOA current at 450 mA, 0.84 mVpp with ITC510 off!!
    Magcheck on B-pillar: ITC510 completely off, device switched on, TEC on, TEC+LD 450mA on:
    Magcheck on top of outer shield: ITC510 on (shelf above 461nm laser), ITC510 off, ITC510 on floor and on:
    Moving ITC510 to the floor (below 461nm breadboard) removes AC-signal seen on top of outer shield)
    1.13 mVpp signal remains
    Turn off black Keysight scope and Rigol RF-monitor scope fore measurements



    50ms spectra with reduced AC magnetic field?

    Try clock-run to find correct B-field? 4.8279 uT
    100ms spectra with this B-field?

    2022-06-29

    09:20 ion still stored
    uMotion: 1.4% 1.6% 0.5%


    On optical table (C-beam side): Turn off Thorlabs photodiode, HP computer (for phasemeter) - could these cause AC B-field?
    ID23723 50ms spectra, 150 Hz AOM span


    for fun, try 50ms clock-run?
    from short run of C1,C2 pairs, estimate B = 4.8201 uT

    2022-06-28

    09:25 ion still stored. fluorescence quite weak, B-beam counts 29.4 / 0.7 in 10ms (slider pos3?)
    30kphoton uMotion A 0.9%, B 5.6%, C 12.6%
    (uMotion program hangs when creating figure?)
    
    Block extra light from reaching PDH-detector
    - Green LED on EOM (indicating temperature stability) blocked with Cu-tape
    - Lights on cavity ion-pump controller blocked with black cardboard flaps
    
    Try slide pos6 for more fluorescence!?
    
    ID23698 Try running 20 iterations of u-motion minimizer, 40 kphotons/pt
    Set old bias-voltages V12cm , V12diff, V34cm, V34diff = [ -6.47, 0.26, 23.43, 59.5]
    (out-of-repo works?)
    optimized voltages V = [-6.4700 0.2371 23.3268 60.4094 ]
    
    verify after optimization: 0.8%, 1.4%, 1.8%
    
    



    60ms spectra with darkened PDH-detector?
    Driftcomp program was restarted so it may take some time for drift to settle..
    ID23705 is OD5/18dB/60ms spectra, 1Hz step, 150 Hz AOM span

    Repeat with less drift?
    Initialized with comb-linecenter 75798064 Hz
    Drift is ca 25 Hz (674nm) during past 4 h
    Strange results, looks like B-field effect with 1/2/5 scaling for C1/2/4?

    Check uMotion:


    ID23714 Larger 1.5 Hz step, wider 250 Hz AOM span, 60ms probe

    ID23716 Repeat with 50ms probe (non-integer number of 20ms cycles?)

    2022-06-27

    clock-laser on, coolinglaser on
    ion is still trapped! (was without cooling from Friday ca 16:30 local, so >2.5 days!?)
    With comb rep-rate at 250 MHz we get beat-notes:
     fb1 109.71 MHz
     fb2 140.19 MHz
    These could both be mixed down to <30 MHz for SDR with e.g. LO = 120 MHz
    giving
     fb1_mix = -10.29 MHz
     fb2_mix = +20.19 MHz
    
    Usual fb1 = -15.5 MHz, ncomb = 889558, f0 = -35 MHz, frep = 249999896 Hz
     f_ir = f0 + ncomb*frep + fb1 = SRS/2 - 165 MHz
    
    With 250 MHz frep and fb1=-109.71 MHz we get: SRS/2 - 166.7 MHz (two cavity modes off from correct mode?)
    (check with rep-rate dds at 19 999 584 Hz -> beat-note at 17.4 MHz)
     - higher piezo voltage, next mode 16.8 MHz
     - next mode 16.1 MHz
     - next 15.5 MHz - this is the correct one!
    back to 250 MHz rep-rate: beat note at 108.0 MHz (this gives same SRS-165 MHz result for the IR frequency)
    
    
    

    Try measuring phase-noise of Keysight 33500B signal generator (trap drive) with Microsemi phase-meter
    No obvious large spur at 50 Hz seen at 1Vpp output (Microsemi measures +4.3 dBm). AM and PN level ca -120 dBc
    Similar result with siggen at 10 Vpp and 20 dB attenuator on Microsemi input. AM spur at 100 Hz offset ca -120 dBc/Hz
    Noise-floor with REF=op7 and DUT=op3 of 10 MHz FDA - seems ok?
    Compare to ARTIQ DDS output, ch3 6 dBatt measures +6.7 dBm by Microsemi - shows larger 100 Hz spur than Keysight - not good??
    (see also 2022-01-27 measurement of Urukul DDS)
    R&S SML-01 is not that good either..


    Check 50Hz and 100Hz background light with Thorlabs amplified detector (AC coupled to scope)
    - sine-shaped disturbance from emergency-exit sign? (block this, or block window between lab and first room?)
    - square shaped disturbance from comb computer screen!!?? (must close door between lab and first room!?)
    

    uMotion check (20kphotons): 1.2%, 2.6%, 12% (!)
    (program fails after measurement, when creating image??)
    slider pos3, 30kphotons: 2.7%, 3.1%, 14%

    Cavity inner shield temperature slope ca -0.2 mK/day
    Cavity drift now ca +28 mHz/s
    OD5/18dB/60ms:
    Gaussian fit to +C1 gives FWHM 7.85 Hz AOM = 15.7 Hz optical (Fourier limit 0.8/60ms = 13.3 Hz) - some residual laser drift during measurement is probable cause
    Fit to -C1 FWHM 7.2 Hz AOM
    B1 estimate = 4.82136950686864 uT

    OD5/23dB/120ms/0.7Hz_AOM_step

    2022-06-24

    16:30 turn off lasers, leave ion trapped

    2022-06-23

    09:45 ion dark
     - all beams, clearout on
     - 405nm on
     - slider pos1
     - (some time later) ion has recovered!?
     - back to B-beam, slider pos2
     - went dark again?? slider unlocks Rb-lock??
     - relock to Rb-line
     - after some minutes ion re-appears..
     - PMT counts low even with slider pos1 -> low 422nm power?
     - was locked to wrong Rb-line, after another re-lock fluroescence is much brighter again
    
    uMotion: 0.9%, 2.2%, 1.3% (slider pos1, locked to wrong Rb-line)


    Clock laser:
    intra-cavity power stabilization to 3.5 VDC on cavity output photodetector
    seems to stabilize 4 digits of photodetector reading: 3.5086(1) VDC (Keysight DMM statistics)

    Cavity cooldown rate now -2.2 mK/day - cavity drift ca +45 mHz/s
    674nm power-monitor on ion-breadboard slightly lower than before, set-point now ca 2 VDC
    Initial spectrum with OD4/18dB/20ms
    B1 estimate =  4.821564486979048
    (with same coil currents we now get significantly lower B-fields?)
    2022-06-20: 4.8302 uT
    2022-05-17: 4.8404 uT
    
    

    OD4/23dB/40ms: (fit to +C1 fails due to dark moment?)

    OD4/28dB/80ms:


    Switch to OD5 for clock-laser, this should give tau_pi of 60ms, 120ms, and 240ms with usual AOM attenuations of 18dB, 23dB, and 28dB
    OD5/18dB/60ms

    OD5/23dB/120ms

    OD5/28dB/240ms

    2022-06-22

    Ion still stored.
    uMotion: 1.7%, 1.5%, 2.7%

    Cavity inner shield temperature now -3.9 mK/day

    2022-06-21

    Cavity continues to cool down. Slope -7 mK/day, drift ca +72 mHz/s


    Comparing recent 80ms spectra to those measured just after ROCIT campaign we see little or no improvement?
    Changes made: 
    - ECDL, BOA, doubler now on main optical table (cable and fiber-lengths very similar to old setup)
    - EOM drive PA now new 1 W mini-circuits
    - more filters in PDH-loop to suppress 2*25 MHz from PDH-detector and 25 MHz through FALC servo
    - new PDH detector based on OPA818 with (much) better noise floor - seems shot-noise limited at 25 MHz
    


    ion still stored
    check uMotion: 3.5%, 4.7%, 3.9%


    ID23573, Repeat OD4/28dB/80ms spectra
    (Stop run during C2-)

    2022-06-20

    461nm temperature set-point was off, set to old value of 4.642 kOhm (notes 2021-12-09)
    461nm current to 194.7 mA based on helper-dispenser bright spot

    15:38 load with 1.75A
    15:49:30 ion
    Initial uMotion: 17%, 19%, 7%

    Run uMotion minimizer with 20 kphotons/pt, 20 iteration

    Final contrast C = -0.0271 -0.0262 -0.0430 RMS = 0.0571
    Final voltages V = [-6.4700 0.2612 23.4365 60.0925 ]
    voltage change dV = 0.0559 0.0530 0.6492
    Verify minimized uMotion: 2.8%, 2.8%, 3.2%s


    ID23562 peak-search with OD0/18dB/200us
    +C1 FWHM 1955 Hz AOM, height 34
    -C1 FWHM 1990 Hz AOM, height 57.1
    B1 estimate =  4.837593013831374
    +C2 FWHM 2110 Hz AOM, height 45.4
    -C2 FWHM 2042 Hz AOM, height 20.89 ??
    B2 estimate =  4.840632485343382
    +C4 FWHM 2365 Hz AOM, height 30.4
    -C4 FWHM 2103 Hz AOM, height 10.7
    B4 estimate =  4.835183060655167
    
    NOTE: these measured with C-beam, this explains unequal peak-heights..

    ID23564 repeat with B-beam
    +C1 FWHM 2004 Hz, height 31.4
    -C1 FWHM 2128 Hz, height 40.5
    B1 estimate =  4.830021041684657
    +C2 FWHM 1914, height 50.7
    -C2 FWHM 1998, height 36.7
    B2 estimate =  4.825156783993923
    +C4 FWHM 2022, height 44.7
    -C4 FWHM 2127, height 36.1
    B4 estimate =  4.830219685373228
    

    Switch to OD4
    ID23566 OD4/18dB/20ms spectra. 4 Hz AOM step

    OD4/23dB/40ms, 2Hz step

    ID23568 OD4/28dB/80ms, 1Hz step (compare to ID23512 (100ms,2022-06-01) or ID23493(80ms, 2022-05-31)


    Cavity drift slightly below 100 mHz/s - drift-compensation reduces this to maybe <2 mHz/s - good enough for ion spectra?
    11:30 allow driftcomp up to 100 mHz/s

    Cavity temperature continues to stabilize slowly. Slope now -11 mK/day

    2022-06-18 kiky Saturday

    Old PDH-detector ca 1.44 VDC laser unlocked, 1.0 V laser locked.
    New PDH-detector with OPA818, 33 kOhm transimpedance, and 20 dB postgain
    Old LO-angle 135 degrees, need to change for new detector: 115.7?
    PDH-signal spectra with old detector: PNG9, CSV9
    new detector PNG10,11, CSV 10,11
    amplitude zoomed view PNG12,13, CSV 12,13
    new detector ca 1.16 VDC laser unlocked
    PNG14,15 wide view of PDH-signal (amplitude overdrive for laser-locked?)
    20local close cavity doors

    Old detector:
    Green SA floor, Blue detector dark, Red laser unlocked, Yellow laser locked. (wide span from 2022-06-15)

    New detector:
    Note slight rise in noise from dark to laser-unlocked, indicating detector is shot noise limited (!?)
    servo-bump height above noise-floor ca 40 dB with old detector, now 45 dB with new detector.
    Note Toptica mixer-input sensitivity changed for new detector

    TODO: check sideband-strength, optimal should be Ps/Pc = 0.5, i.e. half of power in carrier
    check FALC-settings

    2022-06-17

    16:15 turn off clock-laser, leave comb-oscillator

    Changed AOM amplifier to Minicircuits ZX60-100VH+ (max +30 dBm, 36 dB gain), reduced drive level to -18.2 dBm, phase setting = 135 deg.

    Temperature slope -53 mK/day. Frequency drift +318 mHz/s

    2022-06-16

    Closing of cavity doors seen as bump in inner-shield temperature?
    Slope as of 2022-06-16 10local ca -82 mK/day

    2022-06-15

    25 MHz peak in PDH-signal:
    low-pass filter on FALC input?
    There is a clear 25 MHz peak in the downmixed PDH errorsignal also, BLP-15 reduces this peak
    
    no-filter  -12 dBm
    BLP-15     -13 dBm on PDH-detector (ca 30 dB insertion loss at 25 MHz)
    NOTE: with BLP-15 filter need to change LO-angle by almost 180 degrees! now 96.4 deg
    
    With AA 34 dB PA, -6.2 dB EOM drive-level seems optimal?
    transmitted photodiode 
    Carrier 1.8 V
    lower sideband 1.15V
    cant find the other sideband: this may be the problem causing 25MHz on PDH-detector: + and - sidebands not equally strong?
    
    Try mini-circ PA instead?
    Signals improve significantly with lower EOM-drive level! Previously high drive-level caused high-order sidebands!?
    Up to 3.9V on cavity-transmission detector
    
    Now -13.2 dBm (Toptica source), +34 dB (AA 2W PA), BLP-30 = +20.8 dBm
    674nm power:
    ca 31 uW at corner, before EOM
    ca 21 uW towards cavity, after final mirror
    ca 17 uW at PDH-detector
    
    PDH detector
    1.55 VDC when laser unlocked
    1.1 VDC when laser locked
    3.4 VDC on cavity-transmission detector
    -14.2 dBm from Toptica source might give slightly more cavity-transmission
    PDH-mixer input level at +10 dBm
    LO-angle at 135 degrees
    
    
    
    PDH-errorsignal (input of FALC), laser unlocked, without filter and with BLP-15

    FALC monitor output.
    Yellow: no filter on FALC-input: PDH-errorsignal spectrum around 25MHz. Same spectrum + strong 50MHz carrier around 50 MHz
    Red: BLP-15 filter on FALC-input
    Blue: BLP-15 filter response


    Improved EOM-sideband strength (above figures have too strong sidebands!?)
    yellow: laser locked
    red: laser unlocked
    blue: detector blocked/dark - note we are not shot-noise limited since this trace is not lower than previous

    Zoomed view of same signals with 50 kHz span.

    Comb beat-note, rep-rate DDS at 20 MHz, rep-rate P=-6 dB, I=10 Hz
    Note PI gains suggested in Menlo test-report give worse looking beat-note?
    SNR up to 50 dB. Note use SA-settings that do not limit noise floor (noise floor should fall well below photodetector noise with cable disconnected)

    PDH signal spectrum
    Yellow: laser locked
    red: laser unlocked, detector bright
    blue: detector dark
    green: SA floor
    PDH detector 1.51 VDC unlocked, 1.02 VDC locked - so 32% reduction when locked (50% would be ideal if half of power is in sidebands and half in carrier)
    Note ca -15 dB lower noise floor of detector at 5 MHz compared to 25 MHz - shot-noise from unlocked laser starts to rise above dark-noise around here?
    see 2022-06-15_PDHsignal.csv for data

    Final optimized settings from PDH-tuning today:
    EOM-drive -14.2 dBm from Toptica, with AA 34dB PA and BLP-30 filter before EOM resonance-circuit
    +10 dBm input level on Toptica PDH-mixer
    135 degree LO-angle
    BLP-15 low-pass filter on FALC input
    20dB attenuator on FALC output
    FALC SLI: pos 6 
    FALC FLD: pos 6
    FALC FLI: pos 7
    

    cavity inner shield temperature continues to recover slowly:

    ca16:15 local, close cavity doors
    initial drift quite high 700 mHz/s

    2022-06-14

    PDH-spectrum via 10dB coupler on PDH-detector
    unlocked:
    * 50MHz -55.5 dBm (ca +20 dB above noise-floor)
    * 25 MHz ca -60 dBm (ca 10 dB above noise-floor)
    locked:
    * 50MHz ca -45 dBm (up +10 dB from unlocked)
    * 25MHz ca -14 dBm 
    
    Increase drift-comp AOM drive-level, ARTIQ-attenuation now 18 dB (script: drift_comp_v4.py)
    Cavity-transmission photodiode now up to 1.1-1.2 V
    
    
    Locked and unlocked PDH-spectra via 10dB coupler after PDH-detector:

    After increased AOM-drive to driftcomp-AOM, (need to reduce FALC main-gain trimpot)

    Compare to e.g. 2022-03-18 PDH-signal with no visible 25 MHz central peak!?

    After fixing kavi-computer script inner-shield temperature recovers slowly:

    2022-06-13

    Fix issues with cavity temp-control computer writing points to database (stopped working 9 June at some point?)

    Continue with clock-laser
    PDH-mixer now external ZAD-6 mixer with LPF-30 filters on both LO and RF(PDH-detector)
    
    RF-drive -15.2 dBm (seems optimal?) from Toptica to 2W AA-amplifier
    Nominal gain 34 dB -> -15.2+34 = +18.8 dBm to EOM/resonance-circuit
    LO-angle 270 deg
    
    PDH-detector ca 550-560 mVDC when locked - same voltage when unlcoked? (very little coupled into cavity?)
    cavity transmission detector ca 350 mV (300mV with lab lights off)
    
    Cavity temperature control was not working for about 4 days. Now should work again?
    Inner shield temperature ca 32.56 C, higher than normal
    
    ca 14:30 local cavity-transmission up to 0.45 V
    PDH detector, locked 530 mVDC, unlocked 540 mVDC
    674nm power
    - at corner before EOM ca 13 uW
    - towards cavity after last mirror ca 8.2 uW
    - at PDH-detector ca 6 uW
    
    FAIL: reason for problems was FALC main-output connected to laser-head AC-mod, correct input is DC-mod!!
    with correct cabling laser locks similar to earlier.
    large 25 MHz spike in PDH-detector signal (RAM, or RF-pickup from EOM-drive?)
    
    

    2022-06-10

    Continue with clock-laser
    comb rep-rate at 20 MHz DDS -> 250 MHz rep-rate
    
    laser locks to a mode at 109.1 MHz beat-note, with LO-angle around 0 degrees
    mode around 116.6 MHz with LO-angle 212deg
     
    PDH-signal spectrum ugly with spikes at +/-780 kHz?
    
    
    
    

    2022-06-09

    674nm power measurements. ECDL at usual 170 mA setting, BOA at 500 mA.
    (comparison measurements e.g. 2022-01-31 and 2022-02-03)
    
    to cavity-system (fiber output, before isolator): 1 mW
    to ion (fiber output, before fibernoise retroreflector): 107 uW
    
    Cavity input-end corner (before EOM): 14 uW
    PDH-detector: 7 uW
    
    674nm doubler temperature display: 48.7 C
    
    Got laser locked to cavity, but with large servo-oscillation peaks?
    Seems to work only in narrow range of main-gain trimpot
    
    

    2022-06-08

    2022-06-06,7,8: move clock-laser ECDL to main optical table

    Check EOM drive-level at end of long Aircell SMA-cable, close to cavity-box
    
    Toptica dBm    SA dBm
    -12.2          -12.2   seems OK
    (see notes from 2022-02-11 where Toptica source -> PA -> dummyload was verified)
    
    Toptica (-12.2dBm) -> (10dB att) -> PA -> dummyload -50dB port measures -25.6 dBm - seems OK
    (compare to -17 dBm dummyload(-50 dB) port measured without 10 dB attenuator)
    
    With 10 dB attenuator Toptica setting for drive-level probably -15...0 dBm OK?
    
    Overall PDH-lock gain seems weak compared to previous configuration.
    With 20dB attenuator removed from FALC-output (to ECDL current-mod in) get a weak fluctuating lock.
    EOM-drive from Toptica ca -6...+0 dBm, LO-phase maybe 60 deg.
    
    

    2022-06-03

    After yesterdays power-cut:
    - comb f0 lock unlocked, servo at limit?
    - comb frep unlocked, servo at limit
    -- comb auto-control has servoed comb-temperature down to 21.5C from usual 25.5C setting
    -- manually set usual setting 25.5 C
    -- adjust stepper-motor to get frep within locking-range
    -- temperature seems to settle slowly - need to wait 1h or more?
    
    - clock-laser seems still locked
    -- 15.5 MHz beat note had jumped one cavity-mode lower, ca 14.8 MHz now
    - coolinglaser seems locked to Rb-line
    - ion dark or lost?
    - bias-voltage and coil-current drivers seem to have rebooted!?? all values zero?
    - Siglent PSU above trap-setup has rebooted and outputs are off
    - Siglent at back of rack (Zeeman AOM PA supply etc) seems on - not rebooted
    - 
    
    set usual bias voltages and coil currents:
    V12cm , V12diff, V34cm, V34diff = [ -6.47, 0.2053, 23.3835, 59.4433]
    - coil currents fail, no connection to driver at 192.168.10.7 ?
    - rebooting coil-driver fixes problem with Ethernet connection
    I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    
    To get beat-note measurement ch2 to work, need to set ARTIQ DDS ch2 again to 215 MHz
    This means ARTIQ-box has also rebooted to some extent during power-cut??
    
    
    T110 (referece sensor on optical table, post-mounted) reaches ca 22.7..8 C)
    Vaisala ambient sensor from 21.5 C up to ca 24.8 C

    cT4 bottom of Bosch-profile, ca +30mK, -30 mK
    cT3 top of cavity-box, ca +70 mK, -40 mK
    T2 outer shield ca +10mK, -20 mK
    T3 middle shield ca +2 mK, -1 mK
    Cavity T4 inner shield disturbance is ca -1 mK, + 1 mK


    11:50 Comb frep to 250 MHz, stop beat-note datacollection
    15.5 MHz beat-note moves up to ca 108.1 MHz
    high beat-note moves down to ca 141.9 MHz

    More realistic clock-cycle image for slides?


    16:45 turn off lasers - no ongoing experiments

    2022-06-02

    09:50 ion still stored
    for fun, try 100ms clock-run..
    comb set to frep 250 MHz - comb meausrement stopped - clockrun continues
    15:10 stop clockrun for lab-tour
    ion went dark? recovers with all beams and full power
    15:48 continue clockrun after lab-tour?
    15:54 restart with better abs(B) estimate (C4+ counts were low)

    expect power-cut at 16:00 ?? lasting for 90 minutes
    connection to artiq-box lost at power-cut?
    
    computer did not reboot:
    $ uptime
     19:05:53 up 176 days,  1:46, 26 users,  load average: 2,53, 2,12, 2,07
    
    ca 19 ion dark.
    

    2022-06-01

    09:50 ion still stored
    clock laser unlocked!? (Toptica temperature-loop script had crashed, loss of connection)
    LO-phase from Toptica software was 12.9deg - but this gives FALC-output at lower limit -4V?
    Try BLP-30 low-pass filter on PDH-signal, just before Toptica Mixer input. This works with LO-angle 276.4 deg
    EOM drive-level -15.2 dBm now, PDH input-level -10 dBm
    ARC-factor is -0.5 V/V, Fine-in1 controls Piezo voltage

    Try clockrun over lunch while laser-drift settles?
    12:20 start 100ms clockrun. C1 counts ok?, C4 counts low, C2 counts lowish
    abs(B) higher now? restart.

    100ms spectra with BLP-30 filter on PDH-signal (attenuates 2nd harmonic of EOM-signal at 50 MHz)
    RID23512

    2022-05-31

    08:30 stop clockrun, measure 80ms spectra (yesterday spectra with 100ms)
    RID23493


    13:10 continue with 100ms clock-run
    17:20 stop clockrun
    uMotion check: 3%, 1%, 0.7%


    Connect clock-laser equipment to Green/UPS electricity (ahead of Thursday 2022-06-02 power-cut)
    NOTE: Toptica EOM drive-level goes back to default +12 dBm after reboot! This may be too high for our EOM!?
    Set drive level -21.2 dBm - we now get up to +3.2 VDC on the cavity through-transmission detector
    This might be a lower EOM-drive level compared to previously?

    18:20 continue with 100ms clock-run
    20:50 stop clockrun and measure 100ms spectra
    RID23506

    2022-05-30

    Menlo CombWatch program takes 1.4 Gb of memory -> crash 'out of memory'

    Try driving 1348nm BOA with Vescent current driver (Vescent D2-105) - Instead of usual Thorlabs ITC 510
    Vescent is Common-Cathode condiguration, current output from SMA-connector.
    Adapter to Thorlabs Butterfly-mount 9-pin connector: pin-7: LDC = GND/Shield of SMA, pin-3: LDA = middle-conductor of SMA
    15:42 load with 1.75A
    12:58 ion! 
    
    check initial uMotion RIDs 23472,3,4
    3.1%, 1.9%, 3.5% - OK
    
    voltages are low-EQS setting from last run: [ -6.47, 0.2053, 23.3835, 59.4433]
    
    16:09 try to start 100ms/OD4/28dB clock-run, to see laser-drift and abs(B) for today
    16:39 set slider pos3, OD0.8? to get more fluorescence, with pos2=OD1 got ca 30 counts/10ms - leads to false-dark events with threshold=10?
    17:00 ion went dark - recover manually with slider pos1, 405nm, all beams
    
    


    100ms spectra with Vescent driving BOA!?
    RID23489

    2022-05-27

    06 ion dark, stop clockrun, recover?
     ion lost?
    
    14 local turn off lasers, comb
    

    Seculars from yesterday
    Note meaasurements start and end with -6.5 V12cm, change in Z during measurements ca -80 Hz /1.949 MHz = -4e-5


    Interleaved EQS run, only 100 cycles before ion went dark
    -6.47 V12cm gave low-EQS closer to zero compared to previous runs

    2022-05-26

    ca 21:20 continue with interleaved lo-hi EQS clockrun
     OD4/28dB/100ms, g1=1.0, g2=0.01, n_shots=30
     cycle time 104.9s (12 servos)
     lo-eqs servos seem locked, ccounts ok
     hi-eqs C4+ counts low or zero?
    
    21:41 restart clockrun
    
    

    jump in clock-laser/cavity caused servos to lock to 50 Hz sidebands!?
    uMotion check: 0.9%, 4.1%, 4.1%

    X Y Z

    Automated tickler runs for changing V12cm
    V12cm  X         Y        Z          RIDs
    -6.5   907711.6  955162.2 1950047.8  23297,8,9
    -7.5   875848.8  924916.1 1980681.1  23301,2,3
    
    new schedule with wider scans to find resonance
           X         Y         Z 
    -6.5   907683.4  955142.6  1950014.4  23342,3,4
    -7.5   875850.4  924967.1  1980706.3  46,7,8
    -8.5   842822.7  893740.8  2010837.4  50,1,2
    -5.5   938505.4  984338.2  1918779.9  54,5,6
    -4.5   968255.2  1012785.5 1887121.2  58,59,60
    -3.5   997088.9  1040426.2 1854927.9  62,3,4
    -2.5   out-of-range        1822235.2  66,7,8
    -1.5   out-of-range        1788866.0  70,1,2
    -0.5   out-of-range        1755077.9  23376
    +1.5   not run
    -6.5   not run
    
    Fit 2nd-degree polynomial to above results, and center scan around this?
    once more:
    V12cm X         Y        Z          RIDs
    -6.5  907646.8  955065.3 1949894.6  23387,8,9
    -7.5  875804.7  924939.1 1980592.4  91,2,3
    -8.5  842777.9  893697.1 2010644.3  95,6,7
    -5.5  938396.6  984326.8 1918689.0  99,23400,23401
    -4.5  968178.6 1012741.3 1887046.6  403,4,5
    -3.5  997103.2 1040381.1 1854858.1  407,8,9
    -2.5 1025147.1 1067295.0 1822143.2  411,2,3
    -1.5 1052521.0 1093522.4 1788870.8  415,6,7
    -0.5 1079167.4 1119204.7 1754848.3  419,20,21
    +0.5 --------- 1144265.2 1720316.1  423,4,5
    -6.5  907639.9  955042.5 1949814.7  427,8,9
    

    2nd run, scans centered based on fit to first run
    -6.5V
    -7.5V
    -8.5V
    -5.5V
    -4.5V
    -3.5V
    -2.5V
    -1.5V
    -0.5V
    +0.5V
    -6.5V

    first run, x and y go out of range at -2.5 V12cm
    -6.5 V
    -7.5 V
    restart scans -6.5 V
    -7.5 V
    -8.5 V
    -5.5 V
    -4.5 V
    -3.5 V
    -2.5 V
    -1.5 V

    2022-05-25

    09:46 still dark
    relock to Rb-line

    15:09 load with 1.75A
    15:21 ion, soon after tuning 422nm to 185 MHz detuning, so maybe this improves loading (although loading works at 208 MHz detuning also)
    ID23241 pmt-counts during loading
    
    uMotion 2.5%, 6.3%, 4.2%
    
    bias voltages at [ -6.5, 0.2053, 23.3445, 59.4433] 
    last secular freqs are 2022-04-05: 912.6, 959.7, 1957 kHz x,y,z
    (Z**2)/(X**2+Y**2) = 2.18369
    
    X-search: 909-916 kHz, 15 dB att (20 dB fixed attenuator on DDS), 5 kphotons/pt
    Y also with 15 dB
    Z with 24 dB
    
    
    V12cm V34cm    X        Y        Z          R=Z**2/(X**2+Y**2)
    -6.0  23.9945  923570.1 970254.5 1935144.2  2.0870 (+15.6, +14.9, -15.3 kHz)
    -6.5  23.3445  908011.5 955401.8 1950461.1  2.1898
    -7.0  22.6945  892321.0 940501.8 1965930.8  2.2994  (change from -6.5V delta X,Y,Z: -15.7, -14.9, +15.5 kHz)
    
    rough model for seculars (from single-voltage secular scan experiment)
    (coefficients for V34cm are 50-fold smaller, not relevant for finding resonance here)
    X + 30 kHz*V12cm
    Y + 28 kHz*V12cm
    Z - 30 kHz*V12cm 
    
    


    search for seculars:
    -6.5 V12cm: X Y Z Zrepeat
    -7.0 V12cm X Y Z Zrepeat
    -6.0 V12cm: X Y Z


    17:20: try clockrun, to see that clock-laser ok
    zero-eqs from line-intersections ca -6.47 V12cm
    19:24 set [ -6.47, 0.2053, 23.3835, 59.4433]
    19:55 switch to lo-hi EQS run
    low-EQS [-6.47, 0.2053, 23.3835, 59.4433 ]
    hi-EQS [ 6.5 , 0.2053, 40.2445, 59.4433] - second set of servos start at low-EQS and ramp to these voltages during first 30 servo-cycles

    2022-05-24

    Low PMT-count maybe helps with keeping ion fluorescence alive - short dark moments seen but clock-run continues without problems.
    g1=0.2 doesn't seem to work very well, QPN is lower but most quantities show oscillations with various periods

    test with n_shots=60 and g1=1.0
    15:02 stop clock-run
    uMotion check:



    Nothing better to do, so run Vbias vs. seculars measurements again? Hope that old secular-prediction might still hold?
    tickler scan-range is +/-4 kHz from prediction
    ion went dark!?
    20:30 still dark - lost?

    2022-05-23

    For fun, try a clock-run with lower g1? Our clock-laser is better than QPN at 1 servo-cycle in any case?
    100ms clock-run with g1=0.2, g2=0.01, n_shots = 30
    
    

    EQS run result


    07:30 stop clockrun
    check uMotion: 5.6%, 8.7%, 8.2%

    RID22983 100ms/OD4/28dB spectra. 150 Hz span, 0.8 Hz step, 187 pts/peak
    Note lab activities and temperature changes lead to laser drift during last spectrum scan (C4-)


    after spectra run uMotion minimization:

    C = -0.0159 0.0435 0.0189 RMS = 0.0500
    V = [-6.5000 0.2053 23.3445 59.4433 ]
    dV = 0.0207  -0.1350  0.3892
    
    after optimization:

    2022-05-21 evening session

    configure for clock-run: coils to current-driver as usual
    calibrator shows 0.628V and 1.24576 A for I2
    turn off calibrator (on optical table, has internal fan)
    18:19 local set clockrun currents I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    initial B-guess might be 4.88 uT now, if high coil-current has magnetized shields a bit?
    do 80ms peak search to find initial B

    20ms peak-search
    C1+ Center =  75.83798798965832, FWHM 95 Hz AOM
    C1- Center =  75.81066871743741, 103 Hz
    B1 estimate =  4.8679716376506414
    
    C2+ Center =  75.85155768902926 (FWHM 104 Hz AOM)
    C2- Center =  75.79709159130498 (109 Hz)
    C2 Linecenter =  75.82432464016712
    B2 estimate =  4.868021776190824
    
    uMotion check: 5.7%, 9.5%, 5.4%

    18:57 start 90ms clock-run
    now at last uMotion-optimizer voltages of V12cm , V12diff, V34cm, V34diff = [ -6.5, 0.1846, 23.4795, 59.0541]
    19:03 restart with better line-center
    19:18 restart with B 4.86816 uT, probe to 100ms (better p_exc?)
    -- cycle time 52.24 s, dead-time 45.1 ms as usual
    -- abs(B) seems to settle around 4.86813 uT
    
    EQS-schedule
    22:00 to -8.5 V ( 2V steps, 4h data per step )
    02:00 to -6.5 V
    06:00 to -4.5 V
    10:00 to -2.5 V
    14:00 to -0.5 V
    18:00 to +1.5 V
    22:00 to +0.5 V (back to -6.5V in 1V/1h steps)
    23:00 to -0.5 V
    00:00 to -1.5 V
    01:00 to -2.5 V
    02:00 to -3.5 V
    03:00 to -4.5 V
    04:00 to -5.5 V
    05:00 to -6.5 V
    
    p_exc is quite low, especially for C4. B-direction has changed as AT-tests have magnetized parts (shields?)
    plots show average of 10 cycles, each 30 shots, so 50/300 = 17% is reasobable level

    2022-05-21

    Summary of AT-search:
    - no clear AT-split seen
    - I2 current stepped in 0.1 mA steps from 1.24495 A to 1.24615 A 
    - A2 coil sensitivity estimated at ca 0.41307 uT/mA, in agreement with previous results
    - Clock laser stable during run, line-center ca 75.824326(2) MHz AOM
    - C2- component measured, because it was strongest
    -- smallest shift C2- at  72.9469 MHz -> 514.35 uT
    -- largest  shift C2- at 72.9441 MHz -> 514.85 uT
    
    AT-split expected when ground-state split equals RF frequency
    omega_S = g_S mu_B B0 / hbar
    -> B0 = omega_RF hbar / (g_S mu_B) = 514.6398 uT with electron g_S 2.002319
    
    

    09:10 ion still stored
    OD4/18dB/20ms spectra, 6Hz step 600 Hz span
    C1+ Center =  77.26829496893637 -> B = 514.5967 uT, FWHM 98 Hz AOM (Fourier limit 20 Hz AOM)
    C1 Linecenter =  75.8243250019277
    B1 estimate =  514.5967863403096
    
    C2+ Center =  78.70314602662296
    B2 estimate =  514.6005831797813
    

    scan only C2-?:

    wider 10 kHz AOM span, 20 Hz step, 499 pts:

    A = 514.5 uT/ 1245.55 mA = 0.41307 uT/mA
    Z2 = 11.1885 kHz/uT, peak moves out of 10 kHz AOM span at deltaB = 10 kHz/11.1885 kHz/uT = 0.9 uT
    0.1 uT step corresponds to deltaI = 0.1 uT / A = 0.242 mA
    
    deltaI 0.2mA should move C2 peak 0.924 kHz optical
    we measure C2- so smaller current (smaller B) moves peak up in frequency
    
    I2 (mA)   C2-
    1245.55   72.9455
    1245.35   72.946
    
    Calibrator set to: 1.24535000
    Calibrator set to: 1.24545000
    Calibrator set to: 1.24565000
    Calibrator set to: 1.24575000
    Calibrator set to: 1.24525000
    Calibrator set to: 1.24585000
    Calibrator set to: 1.24515000
    Calibrator set to: 1.24595000
    Calibrator set to: 1.24505000
    Calibrator set to: 1.24605000
    Calibrator set to: 1.24495000
    Calibrator set to: 1.24615000

    1.2462 A search16, search17, esitimate 514.868783
    this gives A= 0.4131510 uT/mA


    For 514.69uT try 1245.76 mA

    OD4/23dB/40ms, 10Hz step

    2022-05-20

    09:30local stop comb-measurement (comb to be locked to clock-laser)
    shortly after this lock to ion-peaks lost

    First eqs clock-run had large laser drift initially, intersection of (V12cm, EQS) lines not so good

    New measurement shows better agreement


    10:40 local uMotion:

    uMotion minimization, 20 iterations, 50 kphotons/pt
    ion went dark on first try
    try again
    Final contrasts C = -0.0036 -0.0323 0.0365 RMS = 0.0488
    V = [-6.5000 0.1846 23.4795 59.0541 ]
    dV = -0.0404  -0.1082  -0.9492
    

    uMotion after optimization:


    Nominal currents were: I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    Try I2= -450 mA from calibrator (Fluke 5500A)
    OD0/18dB/200us (yesterday maybe used 400us probe with OD0/18dB and only I2 current?)
    C1+ Center =  76.3462834288485
    C1- Center =  75.3025162529413
    C1 Linecenter =  75.8243998408949
    B1 estimate =  185.9869826524296
    
    C2+ Center =  76.86428069815916
    C2- Center =  74.7845072404588
    C2 Linecenter =  75.82439396930897
    B2 estimate =  185.88411699475427
    


    Double to -900 mA, estimate is 371.76823398950853
    voltage over coil 468 mV giving 0.9*0.468 W = 421 mW dissipiation
    +C1 Center =  76.86781562179468 (not so great fit)
    -C1 Center =  74.78168881233164 (not good fit)
    C1 Linecenter =  75.82475221706315
    B1 estimate =  371.7231578825482
    
    +C2 Center =  77.90455998261733, FWHM 2178 Hz AOM (Fourier limit 2 kHz AOM)
    -C2 Center =  73.74426133849755, FWHM 1955 Hz
    C2 Linecenter =  75.82441066055745
    B2 estimate =  371.8354213212059
    


    For 514.6 uT, current estimate based on last measurement 1245.5510 mA
    set I2 = -1245.55 mA. Dissipation 1.24555*0.649 W = 808 mW
    +C1 Center =  77.26860020582984 
    -C1 Center =  74.38059887069626
    C1 Linecenter =  75.82459953826304
    B1 estimate =  514.6077272940057
    
    +C2 Center =  78.7030822575593, FWHM 1889 Hz
    -C2 Center =  72.94554541626057, FWHM 1941 Hz
    C2 Linecenter =  75.82431383690994
    B2 estimate =  514.5919368511186
    


    Keep same current setting, switch to OD2 and 10-fold resolution
    OD2/18dB/2ms spectra
    +C1 Center =  77.26828296565073, FWHM 201 Hz
    -C1 Center =  74.38036757719466
    C1 Linecenter =  75.8243252714227
    B1 estimate =  514.5924126112019
    
    +C2 Center =  78.70311178598891, FWHM 213 Hz
    -C2 Center =  72.94554256322625. FWHM 224 Hz
    C2 Linecenter =  75.82432717460759
    B2 estimate =  514.5948310124063
    


    Switch to OD4 for 10-fold resolution
    OD4/18dB/20ms spectra. 400 Hz span, 4 Hz step, 101 pts/peak
    B2 estimate =  514.5976829710373
    

    OD4/28dB/80ms

    2022-05-19

    ion still stored
    -1000 mA


    15:50 set -450 mA, estimate 186.38 uT
    rough estimate from C1+ position 186.9301 uT

    Compare this to I2 = -450 mA from Tomberg current-driver:

    no peak, try lower B-estimate:
    This looks Fourier-limited, FWHM ca 5-6 points at 20 Hz sampling for 4 ms probe - OK
    estimate from C1+ center visually ca 185.872 uT (for same current setting Tomberg-box gives ca -1 uT lower field)

    Back to usual coil currents I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    
    test clock-run with 4ms/OD2/18dB as sanity-check, C1 linecenter ca 75.82426
    
    17:07 start 80ms/OD4/28dB clock-run?
    
    OD4/28dB/80ms peak search?
    Initial B-estimate 4.8399 uT

    Try OD4/18dB/20 ms search, wider span
    B1 estimate = 4.879539178118089

    18:30 start 80ms clock-run with initial B from 20ms spectra
    19:00 restart
    19:20 restart again (C1, C2 B-estimates ok)
    19:40 seems locked and OK
    
    schedule for voltages
    19:40 - 22:45  -6.5 V
    22:45 - 01:45  -8.5 V (V34cm 20.9877)
    01:45 - 04:45  -4.5 V (26.1877)
    04:45 - 07:45  -2.5 V (28.7877)
    07:45 - 10:45  -0.5 V (31.3877)
    

    2022-05-18

    11:00 local lock seismometer mass
    Unlocking causes a -20 Hz jump in the cavity frequency
    This caused unlocking of some servos, and Dsum was stuck at 100 Hz
    11:26 restart clockrun with initial comb-linecenter and B-estimate from clockrun (note no EQS in initial estimate)
    after around 10 servocycles seems to lock OK, with Dsum low and EQS-slope at previous value. B-estimates agree
    11:44 to V12cm = -6.5 V
    11:57 restart clockrun (C4+ not locking previously)
    13:06 stop 80ms clockrun

    coil-currents nominal setting is I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    set I1=0, I3=0, I2=+11 mA
    
    200us spectra:
    C1+ at ca 75.8385, height only 10/100 with 200us, Center =  75.83849932779522
    C1- at ca 75.8105, even lower height due to S-balance, Center =  75.81046256356215
    B1 estimate =  4.995820239808815
    
    Switch to Siglent power-supply for I2 = 11 mA, try spectra with 400us
    C1+ ca 75.840
    C1- ca 75.808
    B guess ca 5.7 uT
    
    double current to 22 mA, guess is 11.4 uT. PSU reads 22 mA and 13 mV
    peak-search with 800us probe
    C1+ ca 75.8585, corresponds to ca 12.2 uT ? only 4/100 counts with 1000us probe
    try rabiflop to find good probe length? 
    
    back to Tomberg current driver, I2 = -11.4939, other currents 0
    200us gives good counts 38/100
    Center =  75.837844309115 -> ca B 4.82687 uT
    
    double to -23mA, guess is 9.6 uT
    200us C1+ gives lower counts 20/100
    Center =  75.85089231997432 -> ca 9.4768 uT
    rabiflop gives longer tau_pi ca 400 us
    
    double to -46 mA, guess is 18.9536 uT
    400us spectrum, 35/100 counts
    Center =  75.87744390626217 -> ca 18.93923 uT
    
    double to -92 mA, guess is 37.87846 uT
    400us spectrum 22/100 counts
    Center =  75.93069195621851 -> 37.9259 uT
    
    double to -184 mA, guess is 75.85182976 uT
    400us spectrum 24/100 counts
    Center =  76.03731458101723 -> 75.923699693 uT
    
    double to -368 mA, guess 151.8473993874 uT
    400us spectrum 23/100 counts
    Center =  76.25073779014917 -> 151.98268987980717 uT
    
    Switch to Siglent PSU and set same current -368 mA. Siglent shows power 72 mW
    400us spectrum
    ion went dark :(
    ID22805 try again, C1+ peak at Center =  76.25044713947645 -> 151.8791088458876 uT
    C1- at Center =  75.39736589491767
    peak-separation gives 152.0090019 uT
    
    Double to -736 mA, guess is 304 uT. Siglent shows dissipation 289 mW
    24/100 counts, C1+ at Center =  76.68031272359639 -> 305.0741177 uT
    C1- at Center =  74.97154384933688
    peak-separation gives B1 estimate =  304.48243086173903 uT
    
    target is 514.6 uT, for this we want ca 1241.487 mA
    set 1241 mA, Siglent shows dissipation 825 mW initially, dropping to 804 mW in 1-2 min
    400us spectrum, counts 30/100
    C1+ at Center =  77.26628267443004 -> 513.899966 uT
    C1- at Center =  74.38209094926103
    B1 estimate =  513.9288997942084 uT
    current for 514.6 uT would be 1242.620 mA
    
    set 1243 mA, dissipation 808 mW
    C1+ at Center =  77.26889938168418 -> 514.8324989065194 uT
    C1- at Center =  74.38197661585203 (seems wrong? gives line-center with large offset to comb prediction?)
    B1 estimate =  514.415538983639
    
    siglent IP for remote control 192.168.10.58
    
    I1=0,I2=-1241 mA, I3=0, 400us/OD0/18dB spectrum
    B-estimate from only C1+ is 513.9 uT, from C1+/C1- pair 513.93 uT
    peak-separation 5 768 383 Hz


    I2=-1243 mA, 400us/OD0/18dB spectrum
    From C1+ B= 514.83 uT, from C1-pair 514.42 uT


    OD2/18dB/4ms spectrum. reduce span to 1 kHz and step to 20 Hz. 51 pts/peak

    larger span, better centered on C1+
    C1+ Center =  77.26880417156454
    C1- visually ca 74.3796 (this gives line-center within 70 Hz AOM of comb-prediction - ok)
    C1- Center =  74.37959783673728
    B1 estimate =  514.822444006984
    C1 Linecenter =  75.82420100415091
    comb pred.    =  75.824270
    
    C2+ Center =  78.70431166691162 giving estimate B2+ 514.8198924 uT
    C2- print:Center =  72.94418412527843
    C2 Linecenter =  75.82424789609502
    B2 estimate =  514.8234861298571
    
    Ingoring B0, this gives A(2) = 0.414178 uT/mA roughly in agreement with previous measurements
    
    next try a lower current, possibly non-resonant?
    I2=-1240 mA with prediction 513.580951569608 uT
    B1 estimate =  513.590341541764
    B2 estimate =  513.5569082755302
    
    I2=-1230 mA, prediction 509.43917 uT
    C1+ Center =  77.25365320214294
    B2 estimate =  509.3861133798284
    
    I2=-1200 mA, prediction 497.01382409962065 uT
    C1+ Center =  77.21869030452979
    B1 estimate =  496.9322033358911
    B2 estimate =  496.9305133715476
    
    I2=-1100 mA, pred. 455.5960054246523 uT
    C1+ Center =  77.10261091315833 -> 455.5705 uT
    B1 estimate =  455.61677559155976
    B2 estimate =  455.63955267289623
    

    first test of A-T splitting with -1243 mA, estimate is 514.82 uT
    C2 pair:

    test with -1240 mA


    -1230 mA


    -1200 mA


    -1100 mA


    2022-05-17

    13:52 local: close the dog-house
    13:42 local: increase minus-K frequency, 1/2 turns counter-clockwise. Re-center load.
    minus-K load indicator was way too high! Very difficult to adjust with current setting.
    11:17 load with 1.75A
    11:26 ion! ID22653 is pmt-counts during loading
    slider to position 2, OD1 filter for 422nm (rb-lock stays locked)
    uMotion 100kphotons: A 9.7%, B 21 % (!), C 19%
    


    Run uMotion-minimizer, 20 iterations, 50 kphotons/pt
    ion went dark at some point, recovers with slider to pos1, then back to pos2

    clock-laser piezo voltage now quite high 63.95 V, to find the usual 15.5MHz beat-note (comb at usual rep-rate setting)
    This is probably because ecdl temp-servo is at 20.375 C set-point, higher than usual starting value of 20.00 C

    some problems with Rb-lock, ion went dark again. Repeat uMotion run:

    V = [-6.5000 0.2250 23.5877 59.0033 ]
    final contrasts C = 0.0070 -0.0077 0.0152 RMS = 0.0185
    
    check uMotion after optimization: 1.25%, 1.7%, 1.6%

    Check validity of EMM-preserving voltage vector [1, 0, 1.3, 0] V
    V12cm  V34cm     uMotion
    -6.5   23.5877   [1.3, 1.7, 1.6] % (initial point after optimization, see above)
    -4.5   26.1877   [2.4, 2.1, 2.1] %
    -2.5   28.7877   [1.4, 0.7, 2.4] %
    -0.5   31.3877   [2.8, 2.7, 3.2] %
    +1.5   33.9877   [2.9, 1.6, 1.7] %
    +3.5   36.5877   [0.9, 1.7, 3.0] %
    +5.5   39.1877   [2.1, 1.9, 4.7] %
    +6.5   40.4877   [2.1, 1.8, 3.3] %
    -6.5   23.5877   [0.7, 2.0, 2.0] % (IDs 22707,8,9) (back to initial point in 3 steps)
    -7.5   22.2877   [1.0, 0.9, 1.4] % (ID 22711,2,3) (this point has EQS with opposite sign to others?)
    
    -4.5 V12cm:
    -2.5 V12cm:
    -0.5 V12cm:
    +1.5 V12cm:
    +3.5 V12cm:
    +5.5 V12cm:
    +6.5 V12cm:

    674nm ion-breadboard intensity stabilizing photodiode shows slightly less signal compared to previously
    (photodiode was removed for AOM-chirp tests)
    ID22716 Initial clock-transition search
    OD0, 300us probe
    ion went dark during -C1 scan?

    relock to Rb-line and try again. slight adjustment of TPA-input focusing maybe improves Rb-signal problems?
    From last clock run initial estimate was B= 4.843 uT
    B1 estimate =  4.816310740308461
    C1 Linecenter =  75.8236721290117
    B2 estimate =  4.838068367818707
    C2 Linecenter =  75.8236972583399
    B4 estimate =  4.840377332692647
    C4 Linecenter =  75.82369446017181
    

    ID22724 try rabiflop on last peak measured i.e. C4- at 75.755955. Visually tau_pi ca 230 us (above spectra with 300us)


    OD4 filter for probe. 70ms/28dB/OD4 spectra?
    C1+ peak at 75.8373146
    C1- peak at 75.8101594
    this gives B1 estimate 4.838735
    

    try clock-run with 80ms probe. now at -7.5 V12cm
      initially Dsum did not converge, was stuck at -50 Hz
    16:36 restart with B-estimate from C1 and C2 pairs 4.8399 uT
    17:40 one hour of OK clock-run, Dsum OK, EQS-slope as expected slightly negative
      clock-laser shows periodic oscillations with period ca 60/5 = 12 min? amplitude ca 5 Hz at 674 nm
    
    16:45-19:45 at V12cm -7.5 V
    19:45 to V12cm -6.5 V 
    22:45 to V12cm -5.5 V
    01:45 to V12cm -3.5 V
    04:45 to V12cm -1.5 V
    07:45 to V12cm +0.5 V
    

    2022-05-16

    10:32 decrease resonance frequency of minus-K by 1/4 turn (clockwise). Center minus-K load.
    10:06 local open dog-house

    all lasers on and try to trap again after (long?) break? (last trapping experiments 2022-04-20/21)
    15:22 load with 1.75A
    12:42 loaded an ion (maybe two?) - some problem with cooling-laser AOM and Rb-signal...
    various adjustments to 422nm AOM and Rb-beams - maybe slight improvements..

    2022-05-10

    Install seismometer on cavity minus-K.
    Close the dog-house at around 15:25 local

    2022-05-03

    Pinkert2015 thermal expansion and refractive index temperature coefficients
    alpha_L = 5.6e-7 /C
    alpha_n = 1.06e-5 /C
    
    d(fi) / dT = (omega0/c) [ L*alpha_n + L*n*alpha_L]
    n(1550nm) = 1.4682 for telecom-wavelength
    
    For 674nm this gives prediction of 106 rad / degC phase change per 1m of fiber (?)
    For 1348nm prediction is 53 rad / degC per m of fiber.
    
    Fiber from ECDL to comb is P3-1310PM-FC-10
    from 674nm doubler to cavity P3-630PM-FC-5
    from cavity to ion-breadboard also P3-630PM-FC-5, length is 5m (+short fiber from breadboard-corner to 674nm launch below helres)
    
    measured phase/temperature correlation ca 500 rad/C, corresponding to 50% of cavity->ion 5m fiber reacting to ambient lab temperature - plausible result.
    single-pass frequency shift: 0.5*500 rad/C / (2*pi 1/s) = 40 Hz / (C/s)
    
    For the ECDL->comb fiber 10m * 53 rad/degC/m = 530 rad/degC = 530/(2pi) Hz / (C/s) = 84 Hz / (C/s)
    
    If a clock-run starts at T0, duration is 24h, and ends at T0 + 1 degC we get:
    f_shift = [84 Hz / (C/s) ] * (1/(24*3600)) = 1 mHz IR = 2 mHz 674nm = 4.4e-18  shift / (deltaT/day)
    
    

    2022-05-02

    Measure free-running fiber noise of 674nm fiber between cavity and ion breadboard?
    AOM signal from R&S generator, set to 178.501 MHz gives 2 kHz beat-note
    
    

    analysis of AOM-chirp 30min/30min runs
    See [Kazda2015] and [Tyumenev2016] for similar plots
    Kazda ca -12 mHz shift (0.2 degrees in 125 ms) with 1 W RF power (short off-time!)
    Tyumenev  -45 mHz, -20 mHz, -5 mHz using 1.7 W, 0.537 W, and 0.17 W RF power
    
    Shift scales linearly with RF power, as expected
    Campaign setting of 28 dB attenuation corresponds to +17.9 dBm RF power (62 mW) at AOM
    our 'max' setting is +10 dB from this, +27.9 dBm = 617 mW
    
    With 30 min on-time, residual slope at end of run is <1 mHz, so we are confident that the mixing-frequency is correct.
    
    With Tp=80ms, Td=45ms, we expect duty-cycle scaling (1-Tp/(Td+Tp)) = 0.36
    
    

    2022-04-29

    17:20 start 60 hour AOM-chirp measurement
    1800s on, 1800s off. 10 repeats, so 10 hours per attenuation
    6 different attenuations from 18 to 28 dB: [18., 20., 22., 24., 26., 28.]
    
    60 hour run should finish by 5 am local on Monday morning
    

    2022-04-28

    Leave AOM-chirp run over night
    1800s ON, 1800s OFF. initially 18 dB AOM attenuation, 10 repeats
    SDR HW tune  151664009.99845859408378601074, SW tune +1 kHz
    expect -178 uHz correction needed (see below)
    

    Test SDR frequency measurement accuracy against ARTIQ DDS
    ARTIQ settings Frequency in MU:  21344951171411
    Actual frequency in Hz: 75832500.000000626
    deviation from programmed frequency:  +6.258487701416016e-07 Hz
       2*fAOM = 151665000.00000125
    SDR RX tune 151664004.99910995364189147949
    SDR f-translation in software +1 kHz
    
    expected measured frequency: 4.999108701944 Hz
    
    when unwrapping phase with this frequency, get a positive residual of:
    0.001348 rad/s = 0.000215 Hz  (dataset length ca 100s)
    if we subtract this from the expected measured frequency
    and unwrap with: 4.998893701944351 Hz we get essentially a zero-slope residual
    
    repeat with SDR frequency tuned to be ca -10 Hz
    SDR HW Rx 151664009.99845859408378601074
    f-translation +1 kHz
    
    expected frequency: 9.998457342386 Hz
    phase-residual 0.001121 rad/s = 0.000178 Hz
    corrected frequency 9.998279342386246 Hz -> phase residual below 1 uHz
    
    With knowledge of this correction, try to measure equilibrium beat-note again..
    

    2022-04-27

    Second AOM-chirp run over night
    Attenuations:  [18. 21. 24. 27. 30.] dB
    ON for 300s, OFF for 600s
    10 repetitions for each attenuation
    

    2022-04-26

    Attempt to measure AOM chirp.
    Reference beam before Zeeman-AOM, shifted beam is B-beam after cube/B-shutter
    60MHz one-inch-photodetector + Mini-Circ ZFL amp give ca -55 dBm signal (AOM at 75.8325 MHz, so beat-note at 151.665 MHz)
    Measure with B210 SDR, downconverted signal stored at 200 S/s.
    
    artiq programmed measurement sequence:
    AOM on for 120s 
    AOM off for 600s
    repeat 10 times.
    
    These 10pcs on/off cycles are repeated at different attenuations:
    18 dB down to 28 dB in 5 steps
    as attenuation is increased, we will loose SNR for the beat-note. 28 dB attenuation gives <5% of power at 18 dB
    
    From ARTIQ AD9912 DDS, frequency setting is:
    Frequency in MU:  21344951171411
    Actual frequency in Hz: 75832500.000000626
    deviation from programmed frequency:  +6.258487701416016e-07 Hz
    

    2022-04-25

    RF on/off/on on Sunday:

    2022-04-22

    14:15local switch off cooling laser and ASE.

    2022-04-21

    09:04local ion dark
    all beams, clearout, 405nm on. detuning scan.
    slider pos2 (check if this causes Rb-unlock!)
    still dark after many detuning scans
    relock Rb, detuning scan with 405nm
    reconnect coils to current-driver
    back to old coil-currents: I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    ion still dark or lost?

    2022-04-20

    uMotion: 1.5%, 1.7%, 1.5%

    2022-04-19

    06:11local change aux-servos to measure carrier max peak height (slave-servo FWHM=0, both R and B measure carrier peak)
    
    11:32local stop clock-run
    
    11:30local uMotion: 3.7%, 9.2%, 2%

    15:45 local uMotion minimization, 20 iterations, 50 kphotons
    dV = 0.0256 -0.0861 0.1620
    V = [-6.5000 0.2479 23.7342 58.3074 ]

    check uMotion after optimization: 2.1%, 1.9%, 2.3%

    Coil currents in use:
    I1 = -0.0408, I2 = -11.4939, I3 = -0.7741
    Set these to zero, disconnect coils from driver, install 50R terminators on coils
    Try measuring spectrum at night with grounded coils?
    Zeeman
    AOM att   tau_pi approx    spectrum step Hz AOM
    28        70-80 ms         1 Hz
    23        40 ms            2 Hz
    18        20 ms            4 Hz
    
    
    2022-01-09 coil-characterization B0 = [1.222, 0.1271, -0.2233]
    abs(B0) = 1.249 uT
    
    ID22562 peak search 18dB/20ms, 4Hz step, 400Hz span. abs(B) 1.239 uT estimate.
    1.2uT*5.6 kHz/uT = ca 7 kHz Zeeman shift for C1?

    B2 estimate =  1.2424556640503752
    B4 estimate =  1.2424622174806619
    
    ID22564 28dB/70ms, 1Hz step, 150 Hz span

    2022-04-18

    ion dark after only few hours of clockrun
    usual recovery attempts, many detuning scans
    
    recovers during detuning scan id22532
    continue sb-monitoring clock-run
    
    if sidebands come from 50 Hz AC mains, 80ms probe should average over 4 cycles of length 20ms?
    changing to 70ms should show change, now 3.5 cycles of 50 Hz
    from 06local, change to 70ms probe!?
    
    16:42 local, change to measure +/-100Hz sideband height. keep 70ms probe.
    if AC B-field at 50 Hz causes sidebands, do results support modulation-index ratios 1:2:5 for C1,2,4 Zeeman-components?
    
    

    2022-04-17

    20:30local ion dark
    stop clockrun
    all beams, clearout, 405nm
    detuning scans, ion recovers, camera software stuck - restart
    20:39 restart clockrun with 50Hz sideband-probing
    

    2022-04-16

    ion dark
    slider pos1, all beams, clearout, 405nm on
    voltages to low-EQS
    detuning scans
    still dark
    

    RF off/on to empty trap.
    slider pos1/OD0, all beams, 461nm, 405nm
    low-EQS voltages
    (461nm was left on since yesterday, unintentionally...)
    16:33:00 load with 1.80A
    check Rb-lock
    16:44 after Rb-lock check loaded many ions.. changes between diffuse cloud or coulomb-crystal
    empty trap
    16:51:30 load again with 1.80A
    16:56:55 two ions..
    both disappear when changing V12cm from -6.5 to ca +6 V in one step, with lowered RF-drive 6Vpp
    V12cm>0 gives vertical oritentation of 2-ion crystal, usual -6.5V gives horizontal orientation
    back to 10Vpp RF-drive
    17:00:30 load with 1.75A
    17:11:50 ion, one?
    slider pos1/OD1 
    A counts quite low?
    
    PMT counts with OD1 filter low, so measure uMotion with OD0?
    OD0 uMotion: 3.2%, 1.6%, 2.4% seems ok

    10ms B-counts with slider-pos1 OD0 84.6 (9.4) / 3.2 (1.7) - same as previously with OD1??
    relock to Rb-line, now counts 379.7 (21.2) / 3.9 (1.9) - back to normal levels
    to slider pos2/OD1 -> this causes Rb-unlock!?
    slider pos2/OD1 B-counts after manual Rb-relock now normal: 78.6 (9.3) / 1.8 (1.3) with back lab lights on
    uMotion with OD1: 4.0%, 8.9%, 2.4%

    17:25 start normal clock-run, initialized with comb-linecenter and old abs(B) 4.84331 uT
    C4- counts low, C1/C2 counts ok?

    Try to measure side-bands with extra 'slaved' servos.
    After the 6 usual clock-run servos C1,C4,C2 we run 6 slaved servos with same line-centers, but probing points at +/-25 Hz AOM from the carrier
    servo    R probe freq       B probe freq
    EC1+     C1+ - 25 Hz AOM    C1+ + 25 Hz AOM
    (and similar for EC1-, EC4+, EC4-, EC2+, EC2-)
    Similar to low/high-EQS run these ectra servos have indexes +10, -10, +40, -40, +20, -20
    

    Sideband-monitoring run seems to run OK after ca 18:32 local

    2022-04-15 Good Friday

    Thursday low/hig-EQS run:


    ion dark
    od0 for cooling, clearout, 405nm on
    detuning scans
    still dark
    ion lost?
    RF off, on, to empty trap
    11:35local load with 1.80A
    11:42 two ions!? (counts during loading ID22354)
    lost ions when changing from V12cm=-7V to +7V (RF lowered to 4.9Vpp)
    back to 10Vpp RF drive, OD0 for 422nm
    11:46 load again, 1.80A
    11:49:20 - ion 3min 20s (only one?) counts during loading ID22369, OD1 switched half-way
    slider pos2 = OD1 for 422nm
    uMotion: 1.8%, 2.5%, 4.3%


    Detuning scans:


    start low/high-EQS clockrun
    low-EQS servos ok counts and locked, high-EQS C4 pair not locked..
    12:30local restart with updated abs(B)=4.84331 uT

    2022-04-14

    lasers on, trap RF drive to 10Vpp
    
    10:19:30Z load with 1.80A 
    10:27:05 ion. PMT counts during loading ID22272 (loading-time 7min 35s)
    slot2 = OD1 for 422nm
    A counts 59.5 (7.4) / 2.0 (1.4) in 10ms
    B counts 46.8 (7.5) / 2.6 (1.6)
    C counts 24.4 (5.1) / 1.6 (1.3)   - C showed reduced counts already previously
    
    422nm power measured with slim-sensor after beam-objective
    A 0.4 uW
    B 0.3 uW
    C 0.3 uW
    
    C beam better alignment gives up to 500counts/100ms (up from ca 300/100ms)
    
    Loading: 100ms gate (change from OD0 to OD1 at the end, all beams on)

    uMotion: 2.8%, 5.3%, 13.9%


    Detuning scans:


    uMotion minimization:
    final contrasts C = -0.0309 0.0095 0.0183 RMS = 0.0371
    V = [-6.5000 0.2223 23.8203 58.1454 ]
    dV = -0.0256  -0.1477  0.4813
    

    double check uMotion after minimization:


    Start clock-run to find todays abs(B) value?
    initial abs(B) 4.84835 uT
    C1 counts OK? C4 zero
    15 local re-start with abs(B) 4.84745 uT ?
    15:33 re-start again
    Update AHM3 model, until now we were using ab2022-01 values for all plots:
    MJD0 = 59609
    y0 = 2.78848e-13
    y1 = -2.61147e-15
    
    New ab2022-04 model is
    MJD0 = 59669
    y0 = 1.21876e-13
    y1 = -2.62712e-15
    
    clockrun doesn't lock easily to zeeman pairs, check spectra?
    od4/28dB/80ms spectra:
    comb linecenter 75.827620
    +C1 peak ca 75.841209
    -> abs(B) lower now 4842.8 nT ??
    -C1 75.814029
    

    18:45local start clockrun with updated abs(B) value
    p_exc ok for all peaks, EQS ok, Dsum ok
    low/high EQS run, 2nd try.
    Insert 100ms wait after voltage change time.sleep(0.1)
    
    low-EQS voltage is recent uMotion optimum V0=[-6.5000 0.2223 23.8203 58.1454 ]
    previous high-EQS was V12_cm=0, now try even more? EMM-preserving change dV = [1.0 0 1.3 0]
    high-EQS V1 = V0 + 13*dV = [ 6.5   ,  0.2223, 40.7203, 58.1454]
    
    individual voltages, low-EQS V0
    V1 = -6.7223, V2 = -6.2777, V3 = -34.325100000000006, V4 = 81.9657
    
    high-EQS V1
    cycle idx
    01        V1 = -6.288966666666666, V2 = -5.844366666666667, V3 = -33.761766666666674, V4 = 82.52903333333333
    02        V1 = -5.855633333333333, V2 = -5.411033333333333, V3 = -33.19843333333334, V4 = 83.09236666666666
    03        V1 = -5.4223, V2 = -4.9777000000000005, V3 = -32.6351, V4 = 83.6557
    04        V1 = -4.988966666666666, V2 = -4.544366666666667, V3 = -32.07176666666667, V4 = 84.21903333333333
    05        V1 = -4.555633333333334, V2 = -4.111033333333334, V3 = -31.508433333333336, V4 = 84.78236666666666
    ...
    08        V1 = -3.2556333333333334, V2 = -2.811033333333333, V3 = -29.818433333333335, V4 = 86.47236666666667
    10        V1 = -2.388966666666667, V2 = -1.944366666666667, V3 = -28.69176666666667, V4 = 87.59903333333334
    15        V1 = -0.2223, V2 = 0.2223, V3 = -25.875100000000003, V4 = 90.4157
    21        V1 = 2.3776999999999995, V2 = 2.8223, V3 = -22.4951, V4 = 93.79570000000001
    30        V1 = 6.2777, V2 = 6.7223, V3 = -17.4251, V4 = 98.8657
    
    ca19:08local start low-highEQS 12-servo clockrun. Now with 100ms delay after voltage change, and even higher high-EQS V12_cm=+6.5V
    EQS ramps up during the first 30 servo-cycles
    cycle-time 90.475 s, eqs ramp-up takes ca 45 minutes
    19:54local index 30 and full high-EQS

    2022-04-12/13

    Change 2" mirror on 422 breadboard to 0.5".
    
    422 fiber coupling efficiency before modifications: laser breadboard: 730 uW, trap breadboard: 250 uW
    
    After change Rb- error signal slightly better than before, ~2 Vpp. Similar fiber coupling efficiency.
    
    Piezo slider installed.
    
    

    2022-04-11

    Analysis of selected data from low-EQS - high-EQS comparison run
    (data where laser drift was large not included)
    low-EQS-clock - high-EQS-clock = ca -6(3)e-17 from ca 2000 12-servo-cycles, ca 55 hours of data

    2022-04-09

    ion dark
    all beams, clearout, 405nm, detuning scan, still dark
    no easy recovery, leave all beams + pmt-monitor (405nm off)

    12:16 to OD0 for coolinglaser, check Rb-lock
    405nm on, detuning scan.
    still dark - ion lost?
    turn off coolinglaser, comb, clocklaser

    2022-04-08 Friday


    13:40local terminate clockrun and switch off 422, MNIR, ASEs for height measurements.

    14:55 local restart coolinglaser, comb, ion still stored.
    beat-note was very weak for a while - but then OK (M-NIR warm-up??)

    ASE power measurement (slim sensor): 8.63 mW
    Try improving C-beam alignment, but not much better coutns 200-300/100ms
    also no big gains from ASE-alginment
    To low-EQS voltages V12cm , V12diff, V34cm, V34diff = [ -6.5, 0.2479, 23.968, 57.6641]
    uMotion 4.3%, 7.1%, 7.2%


    15:10local start normal clock-run (not low/highEQS interleaved)
    17:25 RF drive to 6 Vpp

    2022-04-07 Thursday

    high-low EQS clockrun as of Thursday ca 9:40local
    EQS-suppression ca highEQS(5/2) / (low-high sigma) = 12.645 / 0.0151 = 837

    16:35local open tile in floor for height-marker
    coolinglaser unlocked - relock
    cart with electricity-analyzer moved out of lab - may change abs(B) ?

    servos lost lock to transition since clocklaser frequency jumped +80Hz
    17:38 restart clockrun with re-initialized comb-linecenter and B-estimate
    This will start both sets of servos from zero-EQS voltages and ramp up EQS to max during first 30 cycles
    19:35 seems low-EQS locked eventually, but high-EQS had some problems?
    restart clockrun again with lower B-estimate 4.84835uT and comb-linecenter

    2022-04-06 Wednesday

    Interleaved low/high-EQS clock-run.
    12 servos, first set C1,C4,C2 using bias voltages from campaign (V0)
    second set EC1, EC4, EC2 using bias voltages V1 with V12cm=0
    V0 = [ -6.5, 0.2479, 23.968, 57.6641]
    V1 = [ 0.0, 0.2479, 32.418, 57.6641]
    EMM was measured at both voltages and is reasonably low (<10% contrast)
    
    All servos intilized as in campaign at first.
    EQS for the second set is gradually increased from V0 to V1 during the first 30 servo-cycles.
    After cycle 30 second set is kept at V1.
    
    Clockrun parameters otherwise as used in campaign:
    t_probe = 80ms, OD4 filter, AOM_att = 28 dB, g1 = 1.0, g2 = 0.01, n_shots = 30
    12 servos give cycle time of ca 90s.
    
    data stored in clockrun-files for 59674 and 59675 (no other files have data from the 'E' servos)
    E-servos have indexes 10, 40, 20 in the clockrun file.
                low       high
    EQS-slope   28        3793  mHz/mj'-squared  (ratio 135)
    EQS(5/2)    +0.095 +12.643  Hz (ratio 133)
    
    Line-center difference after 549 servo-cycles:
    df = -0.004463 (0.026016) Hz    (stdev scaled with 1/sqrt(N))
    dy = -10(58)e-18
    
    After ca 500 servo-cycles preliminary results:


    2022-04-03 detuning scans analyzed
    Check ASE power and alignment in lab. Also C-beam alignment?


    15:45 local disable comb data collection - work on connecting fibers in comb-box. enable again (need missing parts for fiber, later..)

    EQS-run around 17 local, ca 24h of data, 800 cycles

    2022-04-05 Tuesday

    Carrier peak-search to center scans again
    'a priori' zeeman component centers from comb-linecenter and B-estimate (4.848607 uT): [75.8416583 75.8144477 75.85517744 75.80092856 75.89590717 75.76019883]
    OD0/18dB/200us, 200Hz step, 8 kHz span
    Gaussian fits to spectra
    +C1 Center =  75.841661 Height =  41.3 FWHM =  2054
    -C1 Center =  75.814391 Height =  46.4 FWHM =  2012
    C1 Linecenter =  75.828026 B1 estimate =  4.859030772957006
    +C2 Center =  75.855176 Height =  50.1 FWHM =  1921
    -C2 Center =  75.800871 Height =  42.8 FWHM =  2053
    C2 Linecenter =  75.828023 B2 estimate =  4.8536505160574075
    +C4 Center =  75.895847 Height =  52.5 FWHM =  1915
    -C4 Center =  75.760244 Height =  46.5 FWHM =  1870
    C4 Linecenter =  75.828045 B4 estimate =  4.84483786016688
    

    Check that uMotion reasonable:

    Check secular freqs
    (note text below is artiq-printout, images have separate lorentz-fit with slightly different params!)
    popt index:   0  1  2   3
    lorentzian(f, C, A, f0, g) = C + ( A / ( ( f-f0 )**2 + g ) )
    
    FWHM = np.sqrt(popt[3])*2.0*1.0e6
    center = popt[2]*1.0e6
    
    not clear why artiq-printout and images differ? (try this also in matlab?)
    artiq-fit seems to leave constant C always at zero?
    images may have C>0
    
    Z 24 dBatt 1.955-1.959 MHz: fit params: [0, 0.2866453277079468, 1.9570689655172415, 6.25e-08]
    Y 12 dBatt 0.958-0.962 MHz: fit params: [0, 0.1809272422343231, 0.9597931034482758, 6.25e-08]
    X 15 dBatt 0.909-0.916 MHz: fit params: [0, 0.22061458760248293, 0.9126206896551724, 6.25e-08]
    Z Y X

    +C1 Z BSB at +C1 + Z/2 = 76.82019548275862
    ID22167 is rabiflop on +C1 Z BSB, to 3ms

    ID22168 is same flop to 6ms

    maybe 2ms is OK for looking at RSB/BSB ratio?
    ID22169 is 18dB/2ms Z RSB+BSB spectrum, step 40 Hz to give 5pts/FWHM, span 2 kHz, 51pts/peak
    C1+ Z RSB, BSB
    C1- Z RSB, BSB
    C2+ Z RSB, BSB
    C2- Z RSB, BSB
    C4+ Z RSB, BSB
    C4- Z RSB, BSB

    EQS investigations
    Now at voltages V12cm , V12diff, V34cm, V34diff = [ -6.5, 0.2479, 23.968, 57.6641]
    check uMotions again:

    EMM-preserving voltage vector is Veqs = [1.0, 0, 1.3, 0]
    For V12cm Zero add 6.5*Veqs = [6.5, 0, 8.45, 0] -> [0 , x , 32.418, x]
    uMotion at V12cm , V12diff, V34cm, V34diff = [ 0.0, 0.2479, 32.418, 57.6641]


    To OD4 for 674nm - try to start 80ms clock-run?
    Try interleaved low/high EQS clock-run.
    For each servo-cycle We now run 12 servos, the first 6 with low EQS, and the second 6 with high EQS. Cycle duration ca 90s
    Low EQS same bias-voltages as during campaign: V0 = [ -6.5, 0.2479, 23.968, 57.6641]
    High EQS has V12cm=0: V1 = [ 0.0, 0.2479, 32.418, 57.6641]
    We start with V0 for both sets of servos, and for second set linearly step towards V1 during the first 30 servo-cycles. For index>30 we keep second set of servos at V1 bias.
    18:30local at servo index 16, only half of final EQS seen by second set of servos

    2022-04-04 Monday

    Campaing is over:
    - disable Chronograf/Slack alerts
    - disable daily data-export to ftp-site (comb_data foder)
    
    Change to OD0 for 674nm
    Back to 10Vpp RF drive (ca9:40local)
    Zeeman component centers and line-centers from end of clock-run. If lab-visit doesn't change cavity drift too much, use these for OD0 flops?

    ID22107 is flop on +C1, 25us to 5ms, 50 pts, 18dBatt, 100 shots (default is 1 ms start_delay)
    ID22108 is same flop, but 500ms start_delay - measurement quite slow..
    ID22109 is same flop, 250ms start_delay
    ID22110 is flop with 125ms start_delay
    ID22111 flop with 60ms start delay
    1ms start_delay
    60ms start_delay
    125ms start_delay
    250ms start_delay
    500ms start_delay

    OD0/18dB/200us spectra: (sampling 200Hz AOM is denser than usual, 200us gives 2 kHz AOM FWHM, usual sampling would be 5 pts/FWHM or 400 Hz AOM)

    X  0.9127172
    Y  0.9597931034482758
    Z  1.9572068965517242
    
    X secular
    Y secular
    Z secular

    ID22116 carrier with 200us - aborted

    ID22118 carrier with 200us
    +C1 Center =  75.841600048945   FWHM =  0.0019805568136061387 Height =  47.812942893894956
    -C1 Center =  75.81437703033127 FWHM =  0.0021099454037227794 Height =  46.60398837404041
    C1 Linecenter =  75.82798853963814 C1 B estimate =  4.850820381716847
    +C2 Center =  75.8551206600346  FWHM =  0.0020730820628837047 Height =  44.68321238122099
    -C2 Center =  75.80087380303802 FWHM =  0.001693650472418099  Height =  51.66422095935006
    C2 Linecenter =  75.8279972315363  C2 B estimate =  4.848426676095995
    +C4 Center =  75.89594914157672 FWHM =  0.0018599360616039803 Height =  52.65380800550583
    -C4 Center =  75.76015801466544 FWHM =  0.0020301402620284924 Height =  40.69441032994285
    C4 Linecenter =  75.82805357812109 C4 B estimate =  4.851564603247748
    

    ID22119 red and blue Z-sidebands with 200us - gives low counts 1/100
    +C1 axial red and blue:
    -C1 axial red and blue:
    +C2 axial red and blue:
    -C2 axial red and blue:
    +C4 axial red and blue:
    -C4 axial red and blue:

    axial blue sideband: +C1 + Z/2 = 76.820203
    ID22120 flop on axial blue

    ID22121 axial blue sidebands with 1000us probe (still 400Hz step, corresponding to 200us probe)
    +C1 axial blue
    -C1 axial blue
    ID21123 axial blue, 1000us, better sampling 100Hz, narrower span 8 kHz, 81pts/peak
    peak contained within first 25 points, so could halve span again to 4 kHz
    +C1 axial blue
    -C1 axial blue
    +C1 axial blue center (compare to earlier flop freq) fit fails, visually 75.82020 - agrees ok with earlier estimate
    -C1 axial blue center visually 76.79298
    ID22125 is repeat of +C1 axial blue rabiflop 76.820203 MHz, to 5ms

    ID22126 to 10ms

    ID22127 Compare to rabiflop on carrier with 28dB AOM attenuation?

    ID22128 carrier with 29 dB AOM att

    ID22129 carrier with 30 dB AOM att

    ID22130 with 31dB AOM att (artiq urukul spec is 31.5 dB max attenuation?)


    From the above: use 800us probe, measure carrier with 31 dB AOM attenuation, and axial sideband with 18 dB AOM attenuation.
    ID22133 is axial red and blue sidebands with 18 dB AOM attenuation and 800us probe. 100Hz step, 4 kHz span, 41pts/peak
    +C1 axial red and blue: Data red and blue
    ID22135 is carrier with 31dB att and 800us probe
    C1+ carrier Data C1+ carrier
    C1- carrier Data C1- carrier

    X (lower radial) sidebands with 800us and 18dB. Rabi freq should be double, so we see higher excitation now?
    C1+ X red and blue Data red and blue
    visually +C1 X red 75.38524, X blue 76.29805
    ID22137 is rabiflop on +C1 X blue

    ID22138 same flop, but shorter time to 3ms

    We used 31dB/800us for carrier previously, so now 26dB/400us should work. ID22139 is carrier rabiflop with 26 dB:

    ID22140 is carrier flop with 27.3 dB att:


    X sidebands with 18dB att and 400us probe:
    C1+ X red and blue Data red and blue

    Carrier with 27.3dB att and 400us probe
    Data C1+ carrier

    Y sidebands with 18dB att and 400us probe:
    C1+ Y red and blue Data red and blue

    Carrier with 27.3dB att and 400us (same carrier measurement as previously for X)
    Data C1+ carrier
    C1- carrier Data C1- carrier

    Try to investigate the extra anomalous peaks. Double span to 16 kHz giving 81pts/peak
    Double peaks are due to similar X-Y (47 kHz) and C1 Zeeman-split (54 kHz)
    ID22148 is Y sideband-scan with 18dB/400us and 16 kHz span
    C1+ Y red and blue Data red and blue
    C1- Y red and blue Data red and blue
    C2+ Y red and blue Data red and blue
    C2- Y red and blue Data red and blue
    C4+ Y red and blue Data red and blue
    C4- Y red and blue Data red and blue

    ID22149 is X sideband scan with 18dB/400us and 16 kHz span
    C1+ X red and blue Data red and blue
    C1- X red and blue Data red and blue
    C2+ X red and blue Data red and blue
    C2- X red and blue Data red and blue
    C4+ X red and blue Data red and blue
    C4- X red and blue Data red and blue

    ID22151 is Carrier scan, 27.5dB/400us, 16 kHz span, 200Hz step
    C1+ Data C1+ carrier
    C1- Data C1- carrier
    C2+ Data C2+ carrier
    C2- Data C2- carrier
    C4+ Data C4+ carrier
    C4- Data C4- carrier

    Queued long duration experiments:
    ID22152 X blue sidebands, 500ms start_delay, 18dBatt 400us probe


    ID22153 Carrier, 500ms start_delay, 27.3dBatt 400us probe


    ID22154 Y blue sidebands, 500ms start_delay, 18dBatt 400us probe
    Tue 07:20 local 22154 ongoing. C2- shows no jumps - why? NOTE: center frequencies might be changed, check frequency range scanned for spectra below!


    ID22155 Z blue sidebands, 500ms start_delay, 18dBatt 800us probe, 100Hz AOM step (X and Y have 400us probe and 200Hz step)
    (This run aborted after C2-, see txt-files for data. (to find txt-file from image URL, remove probe_pulses_ and replace .png with .txt)


    ID22156 Z Carrier, 500ms start_delay, 31dBatt 800us probe
    (These run at least 10x slower than with 1ms start_delay - may not complete during the night?)
    Measure only C1+ carrier - then continue with other measurements.

    C1- carrier:
    (abort at pt46/81)

    2022-04-03 Sunday

    100k photons uMotion:

    200k photons uMotion

    Detuning scans: [-15, +12] MHz AOM from nominal 208 MHz detuning
    A 10ms counts 104.8 (10.2) / 0.5 (0.7)
    B 10ms counts 70.8 (8.8) / 0.5 (0.7)
    C 10ms counts 26.7 (5.8) / 0.7 (0.9) repeat 29.4 (5.4) / 0.5 (0.7)
    A B C
    Z secular scan range was 1.951 - 1.956 MHz, 24 dB DDS attenuation

    no resonance, try +5 kHz higher
    Z ca 1.957
    5-fold more photons, 50 kphotons/pt, centered on 1.957 MHz

    X 909-913 kHz, 15 dB ? Y 955-961 kHz, 12 dB?
    X?
    X 50kphotons centered on 913 kHz
    Y? 956-962 kHz 12 dB:
    Y 50kphotons:

    (start ca 16:35 local) ID22096 spectra, 80ms, 28dBatt. 125 Hz AOM span, 1 Hz step, 125pts/peak

    went dark. recovers with all beams (no detuning scan)
    Try again: ID22104



    ROCIT campaign

    Date Uptime Notes
    2022-04-03 Sun / 59672
    clockrun continues
    15:53 stop clockrun to measure uMotion, detuning-scans, seculars, spectra etc.
    20:30local spectra done, continue clockrun
    
    g-ratio ca 1e6(g-1.668) = 77
    noise in g-ratio ca stdev( 1e6(g-1.668) ) = 13
    Prediction from email 2022-01-17
    RF-drive  1e6(g-1.668)
    10 Vpp    77
    8 Vpp     78.8
    6 Vpp     79.6
    to get noise 13/sqrt(N) < 1 need to average ca 13*13 servo-cycles, about 2 hours?
    
    program RF-drive voltage decrease -1 Vpp every 2 hours
    first step 10 -> 9Vpp at 22 local
    
    
    2022-04-02 Sat / 59671
    14:39local: ion dark
    All 422 beams on, no 405, detuning scan (ID22059). Ion reappeared.
    uMotions: A: 1.4%, B: 3.1%, C: 3.5%
    Back to B-beam
    14:53local reinitialize and restart clockrun
    
    
    2022-04-01 Fri / 59670
    clock-run continues
    
    
    2022-03-31 Thu / 59669
    clock-run continues
    
    
    2022-03-30 Wed / 59668
    clock-run continues
    
    07:30local servo index 2485, ca 31 hours
    
    11:40local ion dark
    
    12:30local all 422 beams on (no 405), detuning scan (ID22049). Ion reappeared.
    
    umotions: A: 1.7%, B: 1.0%, C: 1.4%
    
    Back to B-beam
    
    12:38local reinitialize and start clockrun
    
    2022-03-29 Tuesday / 59667
    clock-run continues
    
    18:30local servo index 1446
    
    2022-03-28 Monday / 59666
       
    clock-run continues
    17:30 local servo-index 4730 (ca 59 hours
    
    23:47local ion dark
    all beams, 405nm on,1st detuning scan id22031 -> recovers (blinking because no clearout)
    umotion id22032 10%, id22033 11.1%, id22034 5.2% (see images for contrasts, these numbers 'histogtam constrast'?)
    uMotion minimizatio 20 iterations, 50kphotons id22035
     ca 6+6+6 s=18 s per point
      V = [-6.5000 0.2468 23.9324 57.6128 ] close to last iteration?
      V = [-6.5000 0.2522 23.9147 57.6200 ]
      V = [-6.5000 0.2479 23.9680 57.6641 ] final
      dV = 0.0235  0.0133  0.2325
    uMotion after optimization id22036, 37, 38
     1.8% 1.0% 1.4%
    00:21local continue clockrun
      went dark again??
      recovery id22044
    00:24 clockrun start, old servo positions
    
    
    2022-03-27 Sunday / 59665
       
    clock-run continues
    
    2022-03-26 Saturday / 59664
       
    ion dark, stop clockrun
    try to recover?
    id22021 recovery
    umotion id22022. 3. 4
    05:16local start clockrun from old servo positions
      R-counts high (for C4+ B) - so might lock ok?
    
    
    2022-03-25 Friday / 59663
       
    
    
    2022-03-24 Thursday / 59662
    ion dark
    recovery id22004
    uMotion 6.3%, 8.2%, 2.4%, ID22005,6,7
    06:12local continue clockrun, R-counts for all peaks high so should lock ok?
    
    
    
    
    
    2022-03-23 Wednesday / 59661
    ion dark in early morning 01:51Z
    doesn't seem to recover remotely with OD1 on 422nm
    also no recovery with max 422nm power (around 9 local, after ion dark for 4? hours)
    first loading attempt gave 2 ions
    second loading attempt ok
    uMotions 6.2, 6.7, 1.5% (ID21990,1,2)
    07:27Z start clockrun with B-estimate 4.8479 uT and comb-linecenter
     Note higher abs(B) now ca 4.8485 uT
    
    15:11Z ion went dark for ca 10 servo-cycles
    stop clockrun an recover (ID21999)
    15:16Z restart clockrun with previous servo-positions
     I-terms seem to agree quickly, no significant settling (as seen after 07:27Z start)
     
    
    
    2022-03-22 Tuesday / 59660
    OD4/23dB/80ms spectra look reasonable, but have clear 50Hz sidebands and weaker 100Hz sidebands. 
    C1-pair measured with both 50Hz sidebands visible, C2 only one side visible, C4 not measured
    uMotions <5%
    10:25local restart 80ms clockrun with revised B-estimate 4.484 uT
    10:46local servo-index >25, all excitations >50/300 (similar to previous clock-runs)
     Dsum ok (doesn't diverge, stays <10 Hz)
     B1, B2, B4 consistent ca 4.8478 uT
    11:33 one hour of data. self-comparison looks ok. eqs-slope +0.023 Hz. 
      1e6*(g-1.668) = 76, similar to previous runs
      +0.179 Hz relative to AHM3 model
    
    2022-03-21 Monday / 59659
    no clockrun data :(
    
    linewidth/coherence problems related to either(both?):
    - minus-k not floating
    - pdh-detector tilted
    spectra now look better, drift settles slowly
    
    2022-03-20 Sunday / 59658
    no clockrun data :(
    
    2022-03-19 Saturday / 59657
    various clock-laser tests
    OD0 0.2, 0.4, 0.8us spectra look ok, FWHM close to Fourier limit
    OD2 2ms spectra ok:ish
    
    15local start OD2/18dB/1.5ms clock-run
    Note 18 dB AOM attenuation, so more AOM-chirp than previously?
    beam is B, dark/bright threshold now 10
    cycle time ca 16.8s with n_shots=30
    
    excitation probabilities from 2 hours of 1.5ms clockrun
    ZC     pR    pB
    +1     24.26 24.26
    -1     21.63 21.80
    +2     22.78 22.78
    -2     25.14 25.10
    +4     22.30 22.30
    -4     24.60 24.58
    
    19local stop clockrun to measure spectra
    18dB 1.5ms
    23dB 3.0ms
    28dB 6.0ms
    
    
    2022-03-17 Thursday / 59655
    Work with PDH-lock and possible RAM or electrical feedthrough of 25 MHz EOM signal to PDH-detector
    
    While waiting for cavity to settle:
    15:19local try OD4/18dB/20ms clock-run
    C1 was maybe locked, others not, Dsum diverges
    15:31 restart 
    15:59 restart
    17:00 restart
    17:30 close cavity door
    17:46 restart with OD4/18dB/15ms to see if p_exc changes with 15ms instead of 20ms probe?
    after 30 servocycles p_exc looks similar or slightly worse compared to 20ms run
    20 ion dark
    all beams, 405nm, detuning scan
    ID21738 is revival detuning scan (405nm off, but 405nm during several scans prior to this)
    20:14 continue with OD4/23dB/40ms clockrun - laser drift still 100 mHz/s, B from 20ms clockrun now 4.841 uT?
    p_exc ca 27/300 = ca 10%
    21:52 driftcomp now active, drift < 50 mHz/s
    start OD4/28dB/80ms clockrun
    p_exc low, dsum not ok.. stop clockrun
    spectra tomorrow when drift settled..
    
    
    2022-03-16 Wednesday / 59654
    07:10Z ion dark, no recovery with usual procedures
    to full 422nm power -> ion recovers! (was dark from ca 19 local yesterday!?)
    to OD1 filter for 422nm
    uMotion 1.4%, 3.5%, 3.9% (ID21632,3,4)
    ID21636 is B-beam detuning scan
    ID21637 repeat
    07:14Z start clockrun, 90ms probe, 28dB DDS att, OD4, PMT threshold now 5
    after 20 servo-cycles highes p_exc ca 20/300 = 7%
    
    ID21642 spectrum, OD4/28dB/80ms, 120 Hz span, 1 Hz step
    
    No clocrun data from today :(
    
    
    2022-03-15 Tuesday / 59653
    5:05Z: Stop clock-run.
    
    ID21556: OD4, 28 dB, 80 ms probe peak search (to check if changing PDH-cables had any effect), 1 Hz step, 110 Hz span, 5 Hz Fourier limit (AOM).
    
    ~07:40: Start service break for improved 422nm collimator/Fiber
    
    Measure coupling efficient for 422 before modifications:
    At fiber collimator on laser breadboard (no OD-filter): 750 uW
    At trap breadboard after collimator: 250 uW
    
    Replaced old collimator with f = 7.5 mm (4.6 mm had too small beam).
    
    Coupling efficiency is similar (750 -> 250 uW)
    
    Found issue where first shot of n_shots always was a bright state.
    This is because coolinglaser and repumper were left permanently ON after the n_shots loop.
    Fix by removing cooling+repumper ON-statements at end of loop.
    
    
    13:50Z fix issue with always bright state on first shot.
    Lot's of laser-drift, so difficult to get clock-run going and low excitation-counts
    13:59Z restart with updated comb-linecenter/B - driftcomp now active but lags behind true cavity drift
    14:43 restart again, p_exc seems low?
    set new bright/dark threshold based on 10ms counts: threshold = 5 (was 15)
    
    
    16:15Z possibly higher p_exc with 90ms probe?
    16:20Z try 100ms clockrun? cycle-time ca 52.2s with n_shots=30
    tried intra-cavity power stabilization servo, set-point is 3.5V on cavity-transmission detector
    back to intra-cavity servo off, cavity-transmission detector now ca 4.3 V
    16:32 try 90ms and 26 dB AOM attenuation?
    16:56Z back to original settings OD4/28dB/80 ms -> excitation only 25/300
    17:34Z restart with 90ms
    ion dark
    id21604 recover from dark ion? all beams, 405nm
    still dark, leaave pmt monitor
    
    
    2022-03-14 Monday / 59652
    
    ca03:30 local, ion dark, stop clockrun (no recovery with 405nm on during clockrun)
    
    ion recovers with all beams, clearout, 405nm - no detuning scans required
    restart clockrun
    
    16local stop clock-run to replace PDH-detector, EOM-drive cables.
    uMotion 2.2%, 1.5%, 2.7% (measurements 21546, 21547, 21548)
    PDH-lock on again, -21.2 dBm LO drive from Toptica, LO-angle optimum now ca 154 degrees
    674nm power
    - at corner, before EOM, 26 uW
    - into cavity 19.5 uW
    PDH-detector 1.5 VDC not locked, down to 1.1 VDC when locked (2 mV when dark) - 0.4/1.5 = 27% drop
    If 50% in carrier, 25+25% in sidebands, ideally expect DC-level to drop 50% (?) - we see 27/50, so couple 53% of carrier into cavity
    
    27%*19.5 uW = 5.27uW carrier into (and out) of cavity
    Transmitted power detector 110 kOhm transimpedance, 0.36 A/W responsivity
    Expected transmitted detector signal 5.3 uW * 0.36 A/W * 110 kV/A = 0.2 V? (17.5-fold too little compared to experiment?)
    
    We see 3.5 VDC on detector -> 3.5 V / 100kV/A / 0.36 A/W = 88 uW? (16.6-fold too much, compared to Thorlabs power-meter reading)
    
    16:30local close cavity doors
    
    16:52local start clockrun with new comb-linecenter and previous B4 estimate
    HOX: fiber-noise servo must be restarted off/on, after clock-laser has been unlocked or off
    17:00 reboot fiber-noise servo -> now get quantum jumps as expected
    PMT-counts drop quite low due to temperature swings
    Cavity frequency ca +3 kHz after working with cables etc
    
    
    2022-03-13 Sunday / 59651
    ca 15:30local load new ion
    minimize micromotion
    dV = 0.0007  -0.0333  0.4846
    V = [-6.5000 0.2389 24.0884 57.3503 ]
    uMotion 1.9%, 2.0%, 1.6% (ID 21530,1,2)
    16:08local start clockrun -> Dsum diverges?
    16:16 restart with comb-linecenter and Best 4.84421 uT (remove expanded FWHM probing-points from start of clock-run)
    C1, C2 ok? C4 not locked
    B1 4.8448, B2 4.8449
    16:21 restart with B2-estimate
    Dsum seems better now, all peak-servos give ok R/B counts
    16:29local index 10: C1 average excitation seems normal 50 jumps / 10*30 probes = 17%
    C2 and C4 still low exc-prob. Dsum OK.
    
    16:33Z clockrun restarted, beam used previously today was C - now changed to B as usual..
    
    better idea for initial convergence of clock run?
    - start with only C1, run 10(?) servo-cycles to see that they are locked (C1-linecenter close to comb, and B1 resonable)
    - add C2, run further 10 (?) servo-cycles to see that C2-linecenter and B2 agrees
    - finally add C4 servos, with now accurate B2 estimate?
    
    
    2022-03-12 Saturday / 59650
    Ion dark at 10:58Z
    All beams on, 405 on, detuning scans..
    ion probably lost
    
    2022-03-11 Friday / 59649
    id21436 detuning scan, all beams and 405nm - still dark
    id21437 detuning scan,, 405nm off
    id21439 repeat with 405nm on 
    Check 422 OK, no ion visible still.
    18:33Z: Try to load, 1.75 A. Ion after ~20min. Probably double, since umotions are 50%.
    19:00Z: Try to load again, 1.75A. Ion! after 10 min. Set detuning back to 208 MHz, and OD filter to 0.9.
    uMotions: A: 6%, B: 6%, C: 6%.
    ID21469: Detuning scan with B-beam
    Tried to start clock-run, counts not too good. Leave B-beam on, and leave lab to let things settle.
    Leave lab at 19:17Z
    Start clockrun at 20:06Z.
    
    2022-03-10 Thursday / 59648
    08Z ongoing continuous clockrun since yesterday
    
    21425 detuning scan to wake up ion, all beams and 405nm
    11:19local uMotion 5.2%, 5.6%, 5.3% (21427,8,9)
    21431 B-beam detuning scan
    B-counts 116.6 (9.2) / 0.6 (0.9) in 10 ms
    11:22 local restart clockrun (no changes to parameters)
    11:38 servo-cycle nr20, initial FWHM-shrink done.
    
    2022-03-09 Wednesday / 59647
    ion dark ca 05:00Z
    recovers during detuning scan 21414 (no 405nm) - with all beams, after 405nm was on for a while
    05:37Z uMotion  4.5%, 5.3%, 5.1% (21416, 21417, 21418)
    B-beam detuning scan 21420
    
    clockrun restarted ca 05:40Z
    data looks valid from 06:00Z
    
    ca 14Z sudden jump in B-field of +0.2 nT
    seen also in servo I-terms as jumps in different directions, followed by (slow) convergence
    
    2022-03-08 Tuesday / 59646 clockrun ongoing, looks ok
    14:40Z restart with increased n_shots=30, based on looking at jump-statistics and binomial distribution
    Previous excitation probabilities gave ca 30% zero-jumps with n_shots=10
    New n_shots=30 should give 10-fold less 0-counts, ca 3%
    
    statistics for 59645 (invalid points removed, ca 5730 servo-cycles valid)
        +pR        +pB          -pR       -pB
    C1  0.135618   0.135560     0.151910  0.154135
    C2  0.177879   0.177774     0.153846  0.152224
    C4  0.156620   0.157335     0.139375  0.139421
    
    excitation-probability not expected to change with n_shots=30, but statistics for discriminant improved?
    
    ca 720 servo-cycles from 14:40Z restart to UTC-midnight (720*45s = 9 hours)
    all excitation probabilities are slightly higher
        +pR        +pB        -pR       -pB
    C1  0.145245   0.144645   0.167128  0.165466
    C2  0.188412   0.188735   0.167359  0.167545
    C4  0.173675   0.173859   0.148870  0.151083
    
    2022-03-07 Monday / 59645 09:00Z clockrun ongoing, looks ok
    No visits to lab -> lab temperature stable

    Longer Ion dark moment around 19:45Z (index 5793) - but ion recovers during clock-run (no manual intervention)
    recovers 5 minutes later 19:50Z (index 5813)
    2022-03-06 Sunday / 59644 2022-03-06T19:00 ion went dark, end of servo-cycle 8912. Last valid cycle 8911
    19:31Z uMotion 4.0%, 4.1%, 4.2% (IDs 21401, 21402, 21403)
    19:33Z B-beam detuning scan ID 21405

    ARTIQ clockrun program revised to store jump counts in 2nd detection window also
    store also eqs52=(10/3)*eqs_slope
    19:36Z clockrun restarted
    19:45Z data looks ok
    2022-03-05 Saturday / 59643 Ion fluorescence recovers with all beams, clearout, 405nm, and detuning scan
    B-beam detuning scan ID21386
    05:27Z uMotion 3.5%, 3.0%, 2.6% (IDs 21387, 21388, 21389)
    06:00Z clockrun restarted, Dsum ok
    2022-03-04 Friday / 59642 Coil currents from 2022-01-25: I(-34deg) = [-0.0408 -11.4939 -0.7741] mA
    2022-02-18 has coil current measurements (artiq set-point, set-point as displayed on box display, and DMM measurement)

    80ms spectra RID21340
    OD0 rabiflops (but unsure about 674nm power-stabilization..)
    uMotion minimization ID21341, V = [-6.5000 0.2382 24.1217 56.8657 ]
    12:46Z uMotion: 1.7%, 0.6%, 0.8% (IDs 21342, 21343, 21344)

    From ca 17:45local 80ms clockrun with OD4/28dB Zeeman AOM attenuation
    n_shots=10, g1=1.0, g2=0.01, Zeeman pair order C1, C4, C2
    Zeeman shift coeffiecents used currently in artiq:
    Z1 5.6120442464394
    Z2 11.1885484963681
    Z4 27.9891412391756 
    

    Ion went dark 2022-03-04T22:09 (-C4 peak, index 1568)
    Last valid servo-index 1567

    2022-03-29 Tue

    2022-03-28 Mon

    recovery: (note blinking, 405nm on, clearout off, all beams)

    uMotion:

    uMotion optimization:

    after optimization:

    ion went dark when starting clockrun, recovery: (no image)

    start clockrun

    2022-03-27 Sun

    continuous clockrun, no notes

    2022-03-26 Sat

    ion dark, recovery id22021

    uMotion:

    continue clockrun

    2022-03-24

    ion dark
    recovery id22004

    uMotion 6.3%, 8.2%, 2.4%, ID22005,6,7

    06:12local continue clockrun, R-counts for all peaks high so should lock ok?

    2022-03-23

    ion dark

    09:03 to max 422nm power (ca 800/1000ms background counts with 405nm on)
    detuning scan with 405nm on
    check Rb-lock, seems ok
    detuning scan again
    ion lost?

    RF off, on to empty trap
    07:11:30Z load with 1.80A
    07:15 two ions.. throw away

    07:16 load again, 1.80A
    07:19:50 ion
    counts during loading ID21984

    to OD1 for 422nm
    uMotion: 6.2%, 6.7%, 1.5%

    B-beam counts 69.1 (7) / 1.6 (1.4) (back lab lights on)
    B from clockrun 4.8479, linecenter from comb
    07:27Z start 80ms clockrun

    around 17local ion dark, recovery:

    2022-03-22

    OD4/18dB/30ms:

    OD4/23dB/80ms:
    (attempted clockrun with low p_exc was locked to -50Hz sideband of C1 and C2?)


    uMotion: 4.3%, 4.2%, 4.1%


    Continue with 80ms clock-run, B-estimate now 4.848 uT?
    10:25local restart again, now maybe p_exc ok for all peaks?

    2022-03-21

    08local ion dark
    all beams, detuning scan id21867
    405nm on, detuning scan id21868
    still dark..
    09local still dark
    detuning scan, all beams + clearout
    scan with 405nm on, still dark
    422nm max power, check Rb-lock
    empty trap with RF off, on
    07:18:30 load with 1.80A
    07:24:05 ion
    pmt counts during load id21876
    to OD1 filter for 422nm
    uMotion 4.8%, 4.7%, 2.4%

    Detuning scan:

    OD4/23dB/30ms with new ion?
    FWHM: +C1 257 Hz -C1 264 Hz, +C2 288Hz AOM

    Cavity transmission ca 0.65V now - try to improve alignment?
    PDH detector DC-output 0.72V when laser locked, 0.75V when laser unlocked (very little change??)
    transmission improves to ca 1.77V with alignment
    pdh detector 0.75V unlocked, 0.55V locked
    11:15local cavity transmission ca 1.65V - door closed again, will it drop as components warm up?
    OD4/23dB/30ms after alignment of cavity (laser drift now large..)

    replace fiber-noise VCO/servo with fixed freq?

    Add BLP-30 filter on PDH-detector, removes 50MHz component before mixer? this changes LO angle to 244.3deg
    FWHM 248 Hz - no major change from previous results

    FALC SLI on removes bump in PDH error signal?, 50Hz, 250 Hz, 350 Hz, 450Hz, 550Hz spikes remain

    ion recovers with all beams, 405nm, detuning scan
    FWHM 231 Hz AOM

    mini-circ amplifier on PDH signal?

    remove mini-circ and low-pass filter from PDH signal

    cavity transmission now up to 2.0 V
    PDH-detector DC 0.55-0.61V locked, 0.85V unlocked
    Make sure minus-k is floating! (seems it was not previously??)
    Tilt PDH-detector as much as possible with M6 fixing screws in place
    Laser drift after working with cavity..
    OD4/23dB/30ms spectra anyway while we wait for drift to settle:

    Try OD4/28dB/80ms? still large laser drift - too much drift, need to wait..

    ca21local drift -36mHz/s. OD4/28dB/80ms:
    1 Hz AOM step takes 100*(80ms+45ms) = 12.5s
    scan alternates steps left/right, so in one direction we make progress 1Hz optical/12.5 s = 80 mHz/s
    but driftcomp AOM is at max drift ca +50 mHz/s, so actual drift before Zeeman AOM ca -75 mHz/s -> spectrum scan makes very slow progress in one direction..

    increase step-size to 2Hz AOM, drift now reduced to -22mHz/s, driftcomp still at 50 mHz/s, so Zeeman AOM sees -71 mHz/s

    2022-03-20

    10local ion bright
    uMotion:

    Detuning scans:

    B-beam OD2/18dB/1.5ms:
         B-beam        A-beam        C-beam
    ZC   Height FWHM   Height FWHM   Height FWHM
    +1   36.7    318   8.3     344   20.1    352
    -1   33.1    397   9.4     332   51.8    335
    +2   32.8    372   7.2     341   40.5    341
    -2   33.3    398   5.7     326   17.4    341
    +4   34.3    348   11.5    209   23.8    357
    -4   31.3    403   5.9     365   12.1    298
    
    B:
    A:
    C:
    16:30local uMotion:

    detuning scans:

    OD2/18dB/1.5ms, B-beam
    ZC  Height FWHM freq
    +1  33.6   389  75.84218945598637
    -1  36.2   352  75.81500590050898
     

    ID21850 rabiflop on +C1
    ID21851 flop on -C1

    Try more 422nm power?? From slot10 OD1 (T=10%) to slot8 OD0.8 (T=15.8%)
    B-counts with OD0.8: 100.1 (8.9) / 1.9 (1.4) in 10ms (back lab lights on)
    B-beam detuning scan with OD0.5:

    OD2/18dB/1.5ms
    ZC  Height  FWHM  freq
    +1  40.0     359  75.842173557124
    -1  37.6     309  75.81502638235217
    

    ID21855 flop on +C1
    ID21856 flop on -C1

    To slot6 OD0.6 T=25.1%
    B-counts 151.7 (10.9) / 1.9 (1.5) in 10ms
    Detuning scan:

    OD2/18dB/1.5ms:
    ZC  Height  FWHM  freq
    +1  36.4     383  75.84219016431294
    -1  29.5     388  75.82860247400467
    


    OD4/18dB/15ms: 200 Hz span, 5 Hz step

    OD4/23dB/30ms, 800 Hz span, 5 Hz step

    2022-03-19

    OD4/28dB/60ms spectra, 200Hz span, 2Hz step
    C1:


    OD4/18dB/15ms
    C1: FWHM 222 Hz AOM, 250 Hz

    C2:

    C4:


    FALC input offset seems to adjust 25 MHz central peak height in PDH signal (why?) - try to minimize this?
    peak -58..-59 dBm with 3 kHz RBW
    OD4/18dB/15ms
    +C1

    OD4/23dB/30ms
    +C1

    Switch to ULI setting 500 (V/s)/V - this should keep PDH error-signal better centered?

    Switch to ULI second highest setting 2000 (V/s)/V

    To Toptica -10dBm input level, need to turn down main-gain, and ULI to 500 (V/s)/V

    Remove BPL-10.7 filter at FALC input - this gives slightly lower noise on PDH-detector?

    back to input level 0 dBm

    input level +10dBm also gives similar PDH error signal spectrum

    PDH error signal, 200kHz span, 3 kHz RBW. Green is laser off resonance not locked.
    yellow/red laser locked - note PDH servoloop pushes in-loop noise below off-resonance noise floor for frequencies <20 kHz

    remove transformer/isolators (FTB-1-6 A15)) from EOM-drive and PDH-detector cables. remove BLP-30 from PDH-detector cable


    Sanity-check OD0/18dB/200us spectra?
        18dB/200us    23dB/400us    28dB/800us    OD2/18dB/2ms
    ZC  Height FWHM   Height FWHM   Height FWHM   Height FWHM
    +1  49.8   1986   47.9    972   34.7   541    31.5   346
    -1  40.8   2020   41.6    970   30.9   554    24.8   315
    +2  41.7   2014   38.2    976   25.5   624    
    -2  52.7   1937   50.2    945   42.1   552
    +4  44.1   1989   46.6    936   29.4   577
    -4  53.2   1919   52.3    988   44.1   447
    
    800us +C1 Center =  75.84217488955069
    800us -C1 Center =  75.81502686778107
    2ms +C1 Center =  75.84218164469243
    2ms -C1 Center =  75.8149971315586
    
    C1:

    C2:

    C4:


    OD2/18dB/2ms
    C1:

    ID21814 rabiflop on -C1 with OD2 and 18 dB - tau_pi 1.5ms
    start 1.5ms clockrun

    OD2/18dB/1.5ms:
    Fourier limit 0.5*0.8/1.5ms = 267 Hz AOM
    ZC  height  FWHM
    +1  34.1    360
    -1  32.4    331
    +2  34.7    358
    -2  38.5    322
    +4  34.1    332
    -4  34.7    347
    


    OD2/23dB/3.0ms:
     
    Fourier limit 0.5*0.8/3ms = 133 Hz AOM
    ZC  height  FWHM
    +1  25.0    287
    -1  19.7    318
    +2  20.4    315
    -2  25.1    296
    +4  22.1    300 
    -4  24.3    313
    


    OD2/28dB/6.0ms:
    Fourier limit 0.5*0.8/6ms = 67 Hz AOM
    



    ion bright, uMotion 2.9%, 2.6%, 1.7%

    B-beam detuning scan:
    B counts 69.8 (9.3) / 0.6 (0.7)
    A-beam detuning scan:
    A counts 104.7 (10.2) / 0.7 (0.8)
    C-beam
    C counts 70.4 (8.1) / 0.7 (0.8)
    seculars (909.4, 956.5, 1948.8) kHz on 2022-03-03
    now get (911.2, 958.3, 1953.6) kHz , delta (+1.8, +1.8, +3.8) kHz compared to 2022-03-03
    x-secular search 908-911 kHz, peak far right. 909-913 kHz scan

    y-secular 955-961 kHz

    z-secular 1948-1954 kHz

    2022-03-18

    07local ion bright.

    OD4/18dB/15ms spectra. 5 Hz step, 400 Hz span
    C1: +C1 FWHM 257 Hz AOM, -C1 247 Hz, B1 estimate = 4.839977476480573

    C2: +C2 FWHM 274Hz, -C2 248 Hz, B2 estimate = 4.84063472389163

    C4: +C4 251 Hz, -C4 256 Hz, B4 estimate = 4.840360175543003


    OD4/23dB/30ms spectra, 3Hz step
    C1:

    C2:

    C4:


    slightly changed PDH-lock gains:
    OD4/18dB/15ms
    C1:

    C2:

    C4:


    Low-pass filter BLP-30 on PDH-detector signal to attenuate 50 MHz component
    C1:

    C2:

    C4:


    OD4/23dB/30ms:
    C1:

    C2:

    C4:


    Clock-laser (back table) devices connected to white electricity (was green/UPS)
    OD4/23dB/30ms
    C1: +C1 FWHM 211 Hz AOM, -C1 FWHM 194 Hz AOM, B1 estimate = 4.841657840419675

    C2: +C2 245 Hz, -C2

    C4:


    15:45local Open cavity doors to look for problems..
    0.9uW at cavity output (Thorlabs power-meter) -> 1.7 V on cavity transmission detector
    PDH detector DC-output 0.49V locked -> 0.7V when unlocked (Thorlabs power meter at detector shows 7.1 uW)
    Wideband view of PDH detector signal. Broad bump is photodetector noise floor (blue). Green trace laser locked. Note 50 MHz (twice EOM frequency component)
    25 MHz useful signal should be where photodetector is most sensitive

    Medium span view of PDH signal. Blue is spectrum analyzer floor. Green is laser far off resonance unlocked. Yellow is locked laser.
    Note no 25 MHz component anyomre in this view! This was possibly doe to low-pass filter on EOM PA output ??

    Narrowband view of PDH-signal close to 25 MHz EOM frequency. Blue spectrum analyzer floor, green laser off resonance unlocked, yellow laser locked
    Note very low or no 25 MHz 'feedthru' signal peak - maybe this was a problem prviouusly? 25 MHz signal on detector is mixed to DC in the rror signal!?

    All measurements through 11 dB coupler placed after photodetector. After coupler Aircell7 cable to back of Toptica box
    17local close cavity doors

    OD4/23dB/30ms
    C1:

    C2:

    C4:


    Wider scan with OD4/23dB/30ms ?
    C1:

    C2:

    C4:


    2022-03-17

    Ion was dark at 14:20 local, reappeared with all beams and full 422 power.

    09local ion still stored and bright
    uMotions: 3.1%, 1.5%, 7.4%

    PDH-lock has been working (intermittently?) with 25 MHz component in PDH error signal
    if this is RAM that goes through the mixer into FALC this may cause problems..
    Zoomed view of PDH-signal (through 11 dB couplter) shows 36-37 dB peak (300Hz RBW)


    12local, uMotion minimizer, 30 iterations 100 kphotons, over lunch..
    12:58 still running..
    dV = [ -0.0145 -0.1337 0.0813 ]
    V = [-6.5000 0.2244 23.9547 57.4316 ]
    Set voltages -6.7244 -6.2756 -33.4769 81.3863

    After optimization: 1.1%, 1.3%, 2.1%

    recovery from dark:

    uMotion 21739, 40, 41

    2022-03-16

    uMotions after recovery:

    B-beam detuning scan:
    (fit fails!?)


    repeat, changed detuning range:


    ID21642 OD4, 28dB, 80ms spectrum
    C1:

    ID21643 change PDH-settings so that in-loop error signal looks like yesterday (red trace yesterday, blue trace today)

    C1:


    Similar setting, but only 20 dB attenuation on FALC->ECDL (previously 23 dB or 30 dB)
    C1:

    Toptica PDH-mixer input level to 0 dBm (was -10 dBm)
    Previously (-10dBm setting) PDH-error signal was triangle-shaped, and FALC output integrates this to square-shaped (indicating overload of mixer stage?)
    with 0 dBm setting both signals more rounded shape. PDH error-signal spectrum similar, with slightly better/lower values close to 25 MHz? (green trace)

    C1:

    Try Toptica PDH-mixer input level +10dBm setting
    Now FALC main-gain trimpot can be turned up, and we get slightly lower noise close to 25 MHz (red trace, green was previous setting 0 dBm input level)
    cavity-transmission detector ca 4.5 V during these tests, EOM-drive level kept at -27.2 dBm



    13local ion dark during spectrum scan
    full 422nm power, 405nm on
    ion recovers after re-lock to Rb-line (405nm was on and 422nm was off during re-lock)
    uMotion 0.5%, 4.4%, 2.7% (ID21658,59,60)


    PDH-detector signal measured through 11 dB coupler
    50 Ohm terminated shows 50-53 mV AC RMS on oscilloscope
    11 dB is 12.6-ratio for power, sqrt(12.6)=3.55 ratio for amplitude
    Photodetector output is 3.55*50 = 177 mV AC RMS - should not be close to saturation?
    
    Try increasing driftcomp-AOM attenuation to 22 dB, this lowers cavity transmission from 4.5V to 2.2V
    
    
    Lower power through cavity, detector ca 1.85V
    OD4/28dB/80ms

    PDH LO-angle to 231.4degrees, 0 dBm input level

    try no attenuator on FALC->ECDL, LO angle 263.6deg

    various settings tried..
    OD4/28dB/80ms:

    OD4/23dB/40ms:

    OD4/18dB/20ms:

    OD4/18dB/10ms, FWHM 165 Hz AOM

    OD4/18dB/5ms: -C1 FWHM 167 Hz AOM, +C1 FWHM 214 Hz AOM

    add 10 dB attenuator on EOM PA input (adds isolation between Toptica PDH-mixer electronics and PA?) - increase drive-level correspondingly +10 dB
    OD4/18dB/5ms

    OD4/18dB/10ms


    insanity check with OD0??
    OD0/18dB/200us spectra:
    comb linecenter 75 827 873
    +C1 75 841 522, Height =  37.6, FWHM 1915
    -C1 75 814 273, Height =  55.9, FWHM 1814
    C1 Linecenter =  75.82789708424843, B1 estimate =  4.855455434463197
    


    insanity check with OD1
    OD1/18dB/600us spectra, Fourier limit 667 Hz AOM
     
    comb linecenter 75 827 870
    +C1  75 841 459, Height =  34.4, FWHM 714
    -C1  75 814 306, Height =  50.2, FWHM 679
    C1 Linecenter =  75 827 882, B1 estimate =  4.838311553162914
    


    OD2/18dB/1.8ms - Fourier limit 222 Hz AOM
    comb linecenter 75 827 867
    +C1 75 841 449, Height =  34.9, FWHM =  233  
    -C1 75 814 288, Height =  46.4, FWHM =  257
    C1 Linecenter =  75 827 868, B1 estimate =  4.839806628910559
    


    OD3/18dB/5ms - Fourier limit 80 Hz AOM (OD3 with OD1+OD2 filters, alternative would be only OD3 filter)
    comb linecenter 75 827 864
    +C1 75 841 439, Height =  16.6, FWHM =  197  (not close to Fourier limit anymore...)
    -C1 75 814 267, Height =  27.1, FWHM =  194
    C1 Linecenter =  75 827 853, B1 estimate =  4.8418711643831935
    


    Alternative OD3 config with only OD3 filter?
    +C1 FWHM 217
    -C1 FWHM -
    


    OD4/18dB/15ms (OD3+OD1 filter) - Fourier limit 27 Hz AOM

    2022-03-15

    ion dark, pmt counts:


    Replace long SMB-SMA cable from Toptica EOM-drive output to PA-input with shorter SMB/SMA-adapter + SMA-SMA cable.
    PDH LO-angle optimum around 290 degrees. Servo-bumps (effective PDH-gain?) not very sensitive to drive-level, around -23.2 dBm seems OK
    ca 2.9 VDC on cavity-transmission detector

    ID21556: OD4, 28 dB, 80 ms probe peak search (to check if changing PDH-cables had any effect), 1 Hz step, 110 Hz span, 5 Hz Fourier limit (AOM).
    C1:

    C2:

    C4:


    17:58local sanity check OD0/18dB peak-search
    +C1 Center =  75.8527774548399,  height 49.6, FWHM 1886
    -C1 Center =  75.82563369416457, height 43.5, FWHM 1974
    B1 estimate =  4.836697553222443
    
    C1:


    ID21591 is rabiflop on +C1

    2022-03-14

    Service-break to change PDH-detector and EOM-drive cables to new Aircell 7 cables
    RF-isolators also removed (no clear measurement shows they were helping), both on PDH-detector and EOM-drive cables
    PDH error signal spectrum before (yellow) and after (red) cable change: (surprisingly large change?)


    uMotion:

    2022-03-13

    15:23local cooling-laser power to max, 405nm on, detuning scan - no ion
    461nm on - no ion
    empty trap with rf off/on
    13:27:30Z dispenser on 1.75A, 2.06 W displayed on PSU: 1.182V and 1.749 A (1.70A and 1.888W note on 2022-02-02)
    check 422nm Rb-lock, seems OK
    422nm detuning to 185 MHz
    dispenser to 1.80A
    13:44:23Z ion! shortly after increasing to 1.80A
    PMT counts during loading is ID21520
    To 208 MHz detuning and OD1 filter for 422nm
    uMotuon A 9%, B 11.5%, C 11.1% (IDs 21526,7,8)
    ID21529 is 20 iterations of uMotion-minimizer, 50 kphotons

    uMotion after minimization:


    2022-03-12

    ion dark, pmt-counts:

    2022-03-10

    uMotion

    Recovery from darkness:

    2022-03-09

    07:40local continue clock run
    07:56 servo-index 20, data looks ok

    B-counts 10ms: 106 / 0.5
    B-beam detunign scan 21420

    uMotion 4.5%, 5.3%, 5.1% (21416, 21417, 21418)

    405nm off, detuning scan 21414 -> recovers

    repeat ID21413
    detuning scan ID21412
    all beams?
    405nm + B-beam detuning scan?
    405nm while clockrun ongoing -> no recovery
    07:30local ion dark

    2022-03-08

    changed to n_shots=30
    clockrun ongoing..

    2022-03-06

    from ca 21:45 looks like ok data, peaks locked, dsum ok.
    clockrun program revised so now window2 darK/bright counts also stored (previously raw PMT counts were stored, but not thresholded jump counts)
    eqs52 is new stored field in database, (10/3)*eqs_slope
    21:36 restart clockrun - no re-initialization, start from last peak centers? (but they drift away in random directions when ion goes dark..)

    B-beam detuning scan:

    uMotion:

    21:29 fluorescence recovers some time after scan
    detuning scan with 405nm on?
    all beams, detuning scan?
    21:24 ion dark, stop clockrun.

    2022-03-05

    Data files from clockrun:
    59642 clockrun raw/minimal data from Zeeman peak servos, six rows per servo cycle (add 2nd detection window jumps, valid flag)
    59642 clockrun2 computed quantities, one row per servo cycle
    59642 comb one row per second
    59642 driftcomp approx. one row per minute (only used for cavity frequency)
    59642 fion Ion vs. maser frequency, one row per second (no valid flags.. must use other files to filter valid data from here)

    07:59 data looks ok, Dsum ok, B1,2,4 consistent
    07:50 restart again
    07:40local restart
    Dsum -50Hz, B4 differs from B1 and B2..
    07:30local start clock-run, same parameters as yesterday
    initial peak positions from comb linecenter and B-estimate print:[75.85444097 75.82726103 75.86794488 75.81375712 75.90862872 75.77307328]
    uMotion:

    B-beam detuning scan

    B-beam counts: 8923 / 45 in 1000ms
    detuning scan -> fluorescence recovers
    all beams, clearout, 405nm
    ion dark

    2022-03-04

    PMT count-distribution seems to have long tail towards low counts? related to 422nm polarization problem?

    Start OD4/28dB/80ms clock-run. B-beam. Trap breadboard 674nm power servo set-point 3.68 VDC
    Bright/Dark threshold left at old value 15 counts/10ms
    n_shots=10, g1=1.0, g2=0.01
    Revised Zeeman pair order should be C1, C4, C2
    Note first 30 servo-cycles have gradually reduced FWHM from 4*FWHM_Fourier down to 1*FWHM_Fourier in order to find lock to peaks easier
    servo cycle duration 15.0 s, dead-time 45.11 ms
    from ca 17:45 seems we are locked to carrier peaks, Dsum reasonable

    Back to using fiber-noise servo - no clear difference in spectra with servo ON or OFF

    Before clock-run:
    - Minimize micromotion - Done 1.7%, 0.6%, 0.8%
    - Find best probe length (current OD4/28dB 674nm power seems ok, giving ca 80-100ms probe) - 80ms seems ok
    - Test 50 Hz sideband improvements, if possible. (try fiber-noise servo off, fixed 178.5 MHz source instead) -> no big change without fiber-noise servo
    - OD0 rabiflops -> ion temperature estimate?
    - RF monitor (oscilloscope, or try rectifier?) -> scope IP address configured, CH4 should give monitor-voltage from HelRes side monitor-antenna

    comb linecenter and B-estimate give peak positions [75.85442997 75.82725003 75.86793388 75.81374612 75.90861772 75.77306228]
    These measured with fiber-noise servo OFF - but should not matter? (80ms spectra very similar with servo ON and OFF)
    Possibly measured with trap-breadboard 674nm power servo off - giving higher than usual 674nm power to ion..
    ID21371 is OD0/18dB flop on +C1
    ID21372 is flop on -C1
    ID21373 is flop on +C2
    ID21374 is flop on -C2
    ID21375 is flop on +C4
    ID21376 is flop on -C4
    C1
    C2
    C4

    80ms is optimal?
    OD4/28dB/80ms carrier spectra, to get B-estimate?
    +C1 height 27.8, FWHM 5.33 Hz AOM
    -C1 height 34.4, fwhm 5.48 Hz AOM
    B1 estimate =  4.843143976250383
    +C2 height 37.4, FWHM 5.76 Hz AOM
    -C2 height 26.2, FWHM 5.72 Hz AOM
    B2 estimate =  4.843154264903667
    
    C1:

    C2:

    C4:


    probe   +C1 max count
    ms
    
    60      23
    70      26
    75      30
    80      39 (outlier?)
    85      33
    90      31
    100     29
    
    75ms: max count

    70ms: max count

    80ms: max count 39

    85ms: max count 33

    100ms: max count 29 at ??

    90ms: max count 31 at 75.854405

    60ms: max count 23

    ID21351 flop on 75.854435 - seems hard to stay on resonance while rabiflop running.. visually rabiflop max at 60ms?
    ID21348 rabiflop on +C1 with OD4/28dB - to see tau_pi? 75.854441 not on resonance..

    Replace fiber-noise servo+VCO with fixed frequency 178.5 MHz (R&S locked to 10 MHz from maser)
    OD4/28dB/80ms spectra: 110 Hz span, 1 Hz step, Fourier limit 5 Hz AOM
    Any change in 50 Hz sideband?
    C1:


    uMotion after minimization: 1.7%, 0.6%, 0.8%

    ID21341 is uMotion-minimizer, 20 iterations, 100 kphotons
    change from previous voltages dV = -0.0109 -0.1065 0.2202
    New voltages V = [-6.5000 0.2382 24.1217 56.8657 ]


    Fits to carrier:




    OD4/28dB/80ms spectra: 110 Hz span, 1 Hz step, Fourier limit 5 Hz AOM
    B estimate 4.8432 uT
    C1:

    C2:

    C4:


    OD4/23dB/40ms spectra: 200 Hz span, 2 Hz step, Fourier limit 10 Hz AOM
    50 Hz optical sidebands at 50/4 = 12.5 steps from carrier?
    C1:


    ID21338 rabiflop on +C1 with OD4/18dB
    Modified FALC settings (blue trace) - servo-bumps now closer in +/-506 kHz - linear slope of PDH-spectrum towards zero?

    ID21335 rabiflop on C1+ with OD4/18dB


    Servo-bumps on beat-note seem clearly smaller now? effective PDH gain reduced?
    1.67V on cavity-transmission photodiode - this was 3V yesterday before closing cavity box??
    increasing FALC main-gain to '4' gives servo bumps as usual, still 1.67V on cavity transmission
    PDH LO angle 276.4 degrees from yesterday. still seems roughly optimial
    PDH EOM drive -24.2 dBm from yesterday. optimum is maybe -21.2 dBm
    PDH input level was 0 dBm from yesterday. -10 dBm gives higher gain - but need to turn down main-gain to avoid excessive servo-bumps?

    Oscilloscope is on close to T110 sensor - shows as high T110 temperature?

    OD4/18dB/25ms spectra: 300 Hz span, 3 Hz step, Fourier limit 16 Hz AOM
    (stop after C4+)
    C1:

    C2:

    C4:

    2022-03-03

    OD2/28dB/8ms, with settled laser drift
    5Hz AOM step, optical 50 Hz sidebands at 5 steps from carrier?
    
    +C1 FWHM 99 Hz AOM (sidebands?)
    -C1 90 Hz
    B1 estimate =  4.842979170129967
    
    C1:

    C2:

    C4:


    OD2/28dB/8ms spectra: 2 kHz span, 5 Hz step, Fourier limit 50 Hz AOM
    B4 estimate =  4.842658064145234
    
    C1:

    C2:

    C4:


    OD2/23dB/4ms spectra: 2 kHz span, 10 Hz step, Fourier limit 100 Hz AOM
    C1:

    C2:

    C4:


    laser drift high after work with cavity..
    OD2/18dB/2ms spectra: 2 kHz span, 30 Hz step, 67pts/peak, Fourier limit 200 Hz AOM
    C1: +C1 FWHM 198 Hz AOM

    C2:

    C4:


    18:30 uMotion: 1.4%, 3.6%, 7.2%


    Work on increasing gain of PDH detector, previously 8.2 kOhm transimpedance replaced with 39 kOhm
    Detector DC-output was not working properly, fix with 100 kOhm and 100 nF RC-filter, 16 Hz low-pass filter
    Detector DC-output now 1.19 VDC
    PDH-detector DC-level 1.23 V laser unlocked, 0.90 V locked - 27% coupled into cavity (optimal sideband power is 50% of total, 50% in carrier)
    Now ca 3 VDC on cavity transmission detector - more than before

    Zeeman AOM drive PA checked on 2022-01-27 power levels seem ok?
    PA output DDS off  -43 dBm
    PA output DDS on   +27.9 dBm     (2022-01-27 measurement has -3 dB less: +24.8 dBm)
    ARTIQ attenuation 18 dB -> output measures -6.3 dBm
    -> PA gain 27.9 - (-6.3) = 34.2 dB
    AMPA-B-30 spec >33 dB gain
    
    ARTIQ attenuation 0 dB -> output measures +10.5 dBm
    back to 18 dB attenuation -> measures -6.3 dBm
    
    based on 2022-01-27 measurements:
    ARTIQ att.    ARTIQ output   PA-output
    18            -8.2           +24.9 dBm    (measured values)
    
    
    ARTIQ att     B-beam thru chamber P_674
    28            90 nW, background 22 nW
    26            179 nW bg 21 nw
    24            364 nW bg 21 nW
    22            715 nW bg 21 nW
    20            1.25 uW bg 13 nW
    18            1.83 uW bg 13 nW
    
    NOTE: when AOM kept on for several seconds, some time-dependence of P674 !? AOM chirp/heating effect?
    
    B beam 674nm intensity vs. time, AOM manually switched on. Thorlabs power-meter voltage output measured with DMM, NPLC=10 (0.2s per point)


    B-beam realtive intensity vs. DDS attenuation:

    tau_pi vs. DDS attenuation:


    ID21300 rabiflop on +C1 with 18 dB DDS attenuation (bright counts now 110-140 counts/10ms)
    ID21301 with 20 dB attenuation
    ID21302 with 22 dB attenuation
    ID21303 with 24 dB attenuation
    ID21304 with 26 dB attenuation (PMT counts now 80-120 / 10ms)
    ID21305 with 28 dB attenuation
    18dB
    20dB
    22dB
    24dB
    26dB
    28dB

    OD0/18dB/200us spectrum:
    +C1 Center = 75.85339931200139, FWHM = 2086


    slot9 OD0.9 filter, B-counts 8576 / 177 in 1000ms
    Fluorescence returns after re-locking to Rb-line. Rb-signal seems weak - maybe laser jumped out of lock?
    OD0 for coolinglaser
    detuning scan? still dark
    13:06 ion dark? all beams, clearout, 405nm


    OD3/18dB/4ms spectra? 1 kHz span, 20 Hz step, 51pts/peak, Fourier limit 100 Hz AOM
    C1: +C1 FWHM 261 Hz AOM, -C1 FWHM 262 Hz, B1 estimate = 4.84187703599264

    C2: +C2 266 Hz,

    C4:


    B-beam detuning scan (rabiflop showed fluorescence of only 3000-4000 counts/1000ms ??)


    ID21290 OD3, 18dB, rabiflop on +C1, to 32ms
    rise to maybe 4ms - then flat.. Bright PMT counts now only at level of 30...40 counts/10ms - previously was ca 100 counts/10ms


    ID21289 OD2, 18dB, rabiflop on +C1, to 16ms - tau_pi ca 1.5ms? (laser drift off resonance height only 30? comb data shows 35Hz AOM laser drift during past hour)


    ID21288 OD1, 18dB, rabiflop on +C1, to 8ms - tau_pi ca 600us


    ID21286, OD0, 18dB, rabiflop on +C1, to 4ms


    OD0/18dB/200us spectra: Fourier limit 2 kHz AOM
    +C1 Center =  75.85334737929657, Height =  43.1, FWHM =  1920
    -C1 Center =  75.82608352850822, Height =  44.3, FWHM =  2075
    B1 estimate =  4.85809619295892
    
    C1:


    Zeeman AOM adjustments
    
    before adjustments, 674nm power at B-beam output through chamber
    18 dB DDS attenuation
    AOM off 0.02 uW - AOM on 1.95 uW
    
    Before AOM-cube, with trap-breadboard intensity servo set-point 3.68V - ca 60 uW
    before B-beam cube ca 23 uW (double-pass AOM efficiency 23/60 = 38%)
      power split into 3 for A, B, C beams
    before A-fiber 5.2 uW
    before B-fiber 6.7 uW
    before C-fiber 8.5 uW
     (sum of these 20.4 uW    loss-beam at B-cube 0.4 uW   some loss in dichro that joins 422nm and 674nm also)
    
    after adjustments still 1.9 uW at B-beam output through chamber
    
    

    uMotion: A: 2.8%, B 3.8%, C 6%

    B-beam detuning scan:
    repeat 1000ms: repeat 100ms: repeat 10ms: repeat 1000ms:
    uMotion: A: (paper blocks shutter..), B 5.2%, C (paper blocks shutter..)

    09:30 ion stored and bright

    2022-03-02

    OD3/28dB/12ms spectra: 5 Hz step
    C1:

    C2:

    C4:


    OD3/23dB/6ms spectra: 10 Hz step
    C1:

    C2:

    C4:


    OD3/18dB/3ms spectra:
    C1: +C1 FWHM 282 Hz AOM

    C2:

    C4:


    ID21254 is rabiflop on -C1 with OD3
    rises to ca 3ms - then noisy?


    OD2/18dB/1.5ms spectra, Fourier limit 267 Hz AOM
    +C1 Center =  75.85328806307895, Height =  29.6, FWHM =  366
    -C1 print:Center =  75.82610870579471, print:Height =  48.120204741294884, print:FWHM =  0.0003301814701812016
    print:C1 Linecenter =  75.83969838443683, print:B1 estimate =  4.843040448493246
    
    C1:


    ID21251 is rabiflop on -C1 with OD2 filter - expect ca 2.4 ms tau_pi?
    tau_pi ca 1.5 ms


    Try 240 us spectra? Fourier limit 1667 Hz AOM
    +C1 Center =  75.85331037205954, Height =  40.7, FWHM =  1488
    -C1 Center =  75.8261581589811,  Height =  53.1, FWHM =  1579
    C1 Linecenter =  75.83973426552032, B1 estimate =  4.838203671623676
    
    C1:


    OD0/18dB/300us spectra - need to do rabiflops to see new OD0 tau_pi?
    +C1 Center = 75.8532862343632
    ID21248 is rabiflop on +C1 - tau_pi now slightly over 200us


    Adjust Zeeman AOM retroreflecting mirror to get more 674nm power in B-beam - seems this alignment was not optimal
    Adjust Zeeman AOM waveplate, Zeeman AOM diffraction angle, and waveplate before B-beam cube for max power. After adjustments OD0/18dB gives ca 2.15 uW 674nm on output side of B-beam (through chamber)

    OD4/18dB/25ms - 300 Hz span, 5 Hz step
    Intra-cavity power servo on - trap-breadboard power servo modified for faster response 150 nF integrator cap changed to 22 nF
    C1: +C1 FWHM 186 Hz, -C1 FWHM 272 Hz, +C1 slightly better than before??

    C2: +C2 FWHM 258 Hz

    C4:


    OD4/18dB/25ms - Fourier limit 16 Hz AOM
    200Hz span, 5 Hz step. PK304 still off
    PK304 back to 35Hz setting during C4
    C1:

    C2:

    C4:


    OD2/18dB/2.5ms - Fourier limit 160 Hz AOM
    1500 Hz span, 50 Hz step
    Turned OFF vibrating motor PK304A (now running at 35Hz) during -C2 scan, C4 peaks scanned with motor off
    C1: +C1 FWHM 396 Hz, -C1 FWHM 300 Hz

    C2: +C2 314 Hz, -C2 307 Hz

    C4: +C4 502 Hz, -C4 413 Hz


    OD0/18dB/400us
    C1: +C1 FWHM 1186 Hz, -C1 FWHM 1028 Hz

    C2: FWHM 908 Hz, FWHM 843 Hz

    C4: FWHM 926 Hz, FWHM 861 Hz


    OD0/28dB/1600us - Fourier limit 250 Hz AOM
    +C1 FWHM 340 Hz AOM
    -C1 FWHM 352 Hz
    +C2 FWHM 329 Hz
    -C2 
    
    C1:

    C2:

    C4:


    OD0/18dB/400us - 8 kHz AOM span, 100 Hz step - 1 kHz AOM Fourier limit
    adjusted Zeeman AOM and 674nm alignment into beam-fibers - 400us might not be exactly tau_pi..
    comb linecenter 75 849 657 (start of run)
    +C1 75 853 275,  height 41.1, FWHM 928 (Gaussian fit doesn't fit rabi-sidebands..)
    -C1 height 57.8, FWHM 966
    C1 Linecenter =  75.83968368838535 B1 estimate =  4.843499657984703
    +C2 height 51.1, FWHM 948
    -C2 height 36.3, FWHM 890
    C2 Linecenter =  75.83967475362617 B2 estimate =  4.846489510685067
    +C4 height 48.8, FWHM 959
    -C4 height 35.1, FWHM 952
    C4 Linecenter =  75.83971115161137 B4 estimate =  4.844432784122737
    
    C1:

    C2:

    C4:


    Back to Toptica as EOM source and mixer
    PDH signal, via 11 dB coupler, before Toptica mixer stage
    green: noise floor, laser unlocked
    yellow/red: laser locked, noise PSD reaches noise-floor close to DC, servo-bumps at +/- 700 kHz
    Now using 20 dB attenuator on FALC->ECDL, -24.2 dBm EOM drive, 0V lock level, 289.3 degrees LO phase
    NOTE RBW here 1 kHz (older figures with 30 kHz)


    Comb data collection zmq-receiver had stopped - restart. Clock-laser still locked

    Continue B-beam count monitor
    B-counts vs. temperature:


    X: 15 dB
    XA: XB: XC:

    Y: 10 dB
    YA: YB: YC:
    Z: now 20kphotons/pt (yesterday 5kphotons/pt)
    ZA: ZB: ZC:

    uMotion:
    A: B: C:

    detuning scans:
    A: B: C:

    2022-03-01

    PMT count monitor, B-beam:
    2022-03-01-counts.txt


    Check secular frequencies with tickler
    x search 0.90-0.91 MHz, 15 dB attenuation, B-beam -> peak ca 0.909 ?
    y search 0.95-0.96 MHz, 15 dB attenuation -> weak resonance 0.956 MHz? try 12 dB attenuation
    z search 1.94-1.95 MHz, 26 dB, B-beam -> 1.9485 MHz
    
        resonance (FWHM) Hz
        A-beam               B-beam             C-beam
    X   909262.9 (500.2)     909205.0 (419.7)   909259.9 (520.3)    (measurement order B, A, C)
    Y   956179.5 (1288.2)    956299.7 (800.5)   956308.5 (777.4)    (order A, B, C)
    Z   1948544.0 (310.0)    1948500.6 (854.2)  1948525.4 (447.7)   (order A, B, C)
    
    cooling-rate proportional to scattering rate?
    A counts 9305.8 bright / 62.6 dark in 1000ms
    B counts 8058.8 bright / 48.0 dark
    C counts 6620.4 bright / 76.0 dark
    
    
    detuning scans:
    A: B: C:
    X:
    XA: XB: XC:
    Y: 12 dB att
    YA: YB: YC:
    Y: 10 dB att
    YA: YB: YC:
    Y: 8 dB att
    YA: YB: YC:
    Z: 24 dB att
    ZA: ZB: ZC:

    17:11 start OD4/18dB/25ms - probably lots of laser drift now..

    C2:

    C4:


    17:05 uMotion:

    17 close doors again - probably need to wait for drift to settle now..
    new more rigid mount for QWP
    QWP at driftcomp-AOM was mounted very loosely..
    open cavity box to look at driftcomp-AOM etc

    Restart again - intra-cavity servo was limiting power for ion-breadboard servo..
    Raise set-point on intra-cavity power, so that breadboard-servo has enough input power to operate..
    OD4/18dB/25ms


    Restart scan - now with intra-cavity power servo enabled. 1.23V set-point on transmitted power detector.
    Tomberg notes on detector have ca 40 mV/uW sensitivity measurements at 633nm - slightly higher at 674nm for Si detector
    OD4/18dB/25ms


    Try various FALC settings - not so clear what signal to optimize... Transmitted power still shows rising spectrum towards lower frequencies which seems wrong.
    OD4/18dB/25ms

    C2:

    C4:


    Cavity transmission photodiode signal, average around 1.8 V with current settings (blue trace)
    short term variations ca 30 mV, 0.03/1.8 = ca 0.02 relative intensity change
    long-term slow variations larger - due to changing input intensity to cavity?


    Removing 10 dB of attenuation on FALC output to ECDL gives the usual servo bumps at ca +/-800 kHz from carrier (both in beat-note and PDH detector spectrum)
    OD4/18dB/25ms
    +C1 FWHM 219 Hz AOM

    C2:

    C4:


    OD4/18dB/25ms: 3.2 Hz step
    NOTE FALC gains not very optimized for new EOM-drive + mixer settings.. beat note has servo-bumps at only +/-200 kHz, but PDH detector spectrum seems OK close to 25 MHz
    C1:



    Try PDH modulation from siglent signal generator? Demodulation using external mixer -> IF to FALC input
    Monitoring 25 MHz carrier on photodiode shows much less variation compared to notes on this see e.g. 2022-02-09. CH1 to EOM drive 40 mVpp, CH2 as LO for mixer 2Vpp
    OD2/18dB/2.5ms - 32Hz step, Fourier limit 160 Hz AOM
    +C1 FWHM 312 Hz AOM (central feature narrower, sidebands on each side?)
    -C1 318 Hz
    
    C1:



    25 MHz modulation frequency from Toptica PDH lock seems to (slowly) vary around 160-170 Hz - but does this matter? Same source used for demodulation internally also
    If EOM resonator has steep phase vs. frequency slope then changing modulation frequency could cause problems??
    Toptica PDH settings: channel2, 25 MHz, -24.2 dBm (40 mVpp), input level -10 dBm, lock level 0V, phase 315 degrees
    

    OD2/18dB/2.5ms - 32Hz step, Fourier limit 160 Hz AOM
    +C1 FWHM 312 Hz AOM
    -C1 292 Hz
    B1 estimate =  4.847911109194807
    +C2 336 Hz
    -C2 380 Hz
    B2 estimate =  4.844192617677165
    +C4 399 Hz
    -C4 329 Hz
    B4 estimate =  4.845883965614396
    

    C2:

    C4:



    OD0/18dB/0.4ms

    C2:

    C4:

    Fourier limit: 1000 Hz AOM.
    C1+ FWHM: 1040 Hz.

    uMotions: A: 0.5%, B: 1.7%, C: 4.6%

    07:58Z: 1.75 A. Ion after 10 min.

    Try to load

    Recheck Zeeman-AOM aperture, correct order now passes through (same as before...).

    Ion seems to have been lost.

    All beams on, 422 to full power, 405 on, detuning scans.
    Ion has gone dark during the night.

    2022-02-28


    1 point takes 9 seconds, 251 points per scan, 6 peaks = 3.8 h. 4 full scans take 15 h

    Scheduled as ID21139 - ID 21142, DB: 2022-02-28-search##: 13 - 16


    OD4/18dB/40ms, search13

    C2:

    C4:


    OD4/18dB/40ms, search14

    C2:

    C4:


    OD4/18dB/40ms, search15

    C2:

    C4:


    OD4/18dB/40ms, search16

    C2:

    C4:


    uMotions: A: 2.5%, B: 2.1%, C: 2.4%

    OD2/18dB/2.5ms
    C1:

    C1+ FWHM: 313 Hz.
    C1- FWHM: 324 hz.

    Remove foil around camera objective.

    No detectable residual 674 nm light after B-beam cube with Thorlabs amplified Si-detector (70 dB gain).

    Replug USB3 camera.

    Unplug USB3 camera
    OD2/18dB/2.5ms
    C1:

    C1+ FWHM: 327 Hz
    C1- FWHM: 320 Hz
    Fiber servo back on.

    OD2/18dB/2.5ms
    C1:

    C2:


    C1+ FWHM: 328 Hz.
    C1- FWHM: 343 Hz.
    C2+ FWHM: 324 Hz.
    C2- FWHM: 291 Hz.

    Replace fiber noise servo VCO with open loop 178.5 MHz from function generator.

    From now on, PK102 on again.

    PK102 off
    OD2/18dB/2.5ms
    C1:


    Fourier limit: 160 Hz.
    C1+ FWHM: 308 Hz
    C1- FWHM: 326 Hz

    PK102 off for C2- restart at ca 14:05
    OD4/18dB/40ms - wide scan with 9 Hz step
    C1:

    C2:

    C4:


    OD4/18dB/40ms - wide scan with 9 Hz step

    C2:

    C4:


    uMotions: A: 0.3%, B: 1.6%, C: 6.5 %

    422 Detuning back to 208 MHz. OD1 ND filter for 422. B-beam counts now 667 in 100 ms.

    9:47Z: Try to load. 1.75 A. Ion! after 5 min.

    8:50Z: Try to load. 1.7 A. No ion after 7 min.

    8:43Z: Try to load. 1.8 A. 2 ions after 5 min.

    detuning-scan w. full power and 405nm? still dark
    422 to full power. Check Rb-lock.

    PMT counts with all beams:

    detuning scan with 405nm on, still dark
    detuning scan with all beams, still dark.
    ion dark.

    2022-02-27

    OD4/18dB/40ms - wider 900Hz AOM span, 9 Hz step
    Fourier limit 20 Hz optical, 10 Hz AOM
        FWHM AOM
    +C1 220 Hz
    -C1 214 Hz
    +C2 238 Hz
    -C2 250 Hz
    +C4 248 Hz
    -C4 263 Hz
    B4 estimate =  4.844944644042546
    

    C2:

    C4:


    OD4/18dB/40ms

    C2:

    C4:


    B-counts 10304 / 85 in 1000ms
    detuning scans with beams A, B, C:

    uMotion:

    08:55 ion stored and bright

    2022-02-26

    uMotion after optimization:

    30 iterations of uMotion minimizer, 100kphotons

    18:30 uMotion:


    Finfoam box around ECDL
    OD4/18dB/40ms

    C2:

    C4:


    rubber feet under ECDL (as used inside finfoam box previously..)
    OD4/18dB/40ms
    clock-laser unlocked..


    FFT-analyzer off, EOM PA (with 120mm computer fan) moved away from close proximity of ECDL
    OD4/18dB/40ms


    Try spectrum without comb-fiber connected (optical feedback from beat-detection splitter, photodiode, comb??) (disconnected in comb-box, before splitter that joins 1348nm laser and comb)
    PM fiber from ECDL to comb was installed 2022-01-17. last reasobable OD5/90ms spectra from 2022-02-09
    reconnect comb after C1+. Turn off FFT-analyze (acoustic noise close to ECDL?)
    OD4/18dB/40ms



    check cavity ring-down (but photodetector on cavity output might be too slow?) (store also CSV data files with same name)
    Photodetector bandwidth ca 100 kHz, so rise-time < 10 us. Measured results ca 8-fold compared to photodetecot rise-time.
    F=Finesse, (1-rho)=loss per round-trip.



    OD4/18dB/40ms, 2 Hz step, 300Hz span, 151pts/peak - 422nm shutter mounted rigidly to trap breadboard
    C1 peaks:
    +C1 FWHM 221 Hz AOM


    Maybe 422nm shutter mounting on rubber-feet (to eliminate vibration noise on comb beat-note) was causing more vibrations to 674nm light seen by ion??
    Try mounting SRS shutter rigidly to table again

    Clock laser had jumped to wrong TEM00 mode!? now 17.5 MHz beat-note!?

    B-beam detuning scan:
    B-counts: 13877 / 92 in 1000ms

    uMotion:

    11local ion still stored and bright

    2022-02-25

    In-loop PDH error signal when laser locked. FFT analyzer results scaled to 1 Hz RBW
    Conversion from dBm to Hz*Hz/Hz may not be completely correct..


    OD4/18dB/40ms, 2 Hz step
    C1 peaks:

    C2:

    C4:


    remove 10 dB attenuator from PDH-detector signal
    Should give 10 dB or sqrt(10)= 3.2-fold more PDH signal?
    modulate with 100mVpp 10mHz, comb data from ca 17:40 -> comb sees modulation amplitude 27.93 Hz -> PDH_coeff = 100 mVpp / (2*2*27.93 Hzpp) = 0.895 mV / Hz 674nm (worse?)
    PDH error signal through 11 dB coupler, after photodetector and mini-circ amplifier (before Toptica PDH mixer input)
    red (lowest) directly from photodetector
    yellow: photodetector -> 10 dB attenuator -> mini-circ +28 dB amplifier
    green (highest): photodetector -> mini-circ +28 dB amplifier


    OD4/18dB/25ms, 3.2 Hz step
    C1 peaks:

    C2:

    C4:


    OD2/18dB/2.5ms (FALC main-gain trimpot at max, 30 dB attenuator). 32 Hz step
    +C1 FWHM 296 Hz
    -C1 295 Hz
    B1 estimate =  4.844254243956416
    +C2 342 Hz
    -C2 291 Hz  (more input gain, slightly less main-gain for this peak)
    B2 estimate =  4.8471796451916385
    +C4 329 Hz
    -C4 308 Hz
    B4 estimate =  4.846315478829265
    
    C1 peaks:

    C2:

    C4:

    With 30dB attenuation on ECDL current control, can use max gain on main-gain trimpot

    Repeat PDH-signal sensitivity test. 10 mHz 100 mVpp modulation to FALC input
    comb data from ca 15:50
    fit to comb data gives amplitude 18.12 Hz (earlier today 34.22 Hz with 200mV -> 17.1 Hz/100mVpp)
    PDH_coefficient = 100 mVpp / (2*2*18.12 ) Hzpp = 1.38 mV / Hz at 674nm (earlier result 1.46 mV/Hz)

    Increase FALC input gain and try again? increased ca 5 half-turns(?) - gain should be adjustable 1-10x (10 turns?)
    comb data from ca 16:10 -> 14.74 Hzp at 1348nm -> 1.7 mV / Hz
    Turn up input gain more, ca 5 turns - now servo-bumps become a lot higher - turn down main-gain trimpot for reasonable servo-bumps (from 16:22)
    comb data from ca 16:22 -> fit to comb data 17.35 Hzp -> 1.44 mV / Hz ?? worse?


    Unplug noisy Thorlabs shutters for A and C.
    OD2/18dB/2.5ms spectra
    Gaussian fit FWHMs: 273, 322, 296, 250, 297, 279
    C1 peaks:

    C2:

    C4:


    C1+ FWHM: 272 Hz AOM (160 Hz Fourier limit)
    OD2/18dB/2.5ms spectra
    C1 peaks:

    C2:

    C4:


    C1+ FWHM: 260 Hz AOM (160 Hz Fourier limit)

    OD0/18dB/400us spectra
    C1 peaks:

    ID21017: OD0 Rabi flop.


    OD2/18dB/4ms spectra
    C1 peaks:

    C2:

    C4:


    OD0/18dB/400us spectra - after adjustments to Zeeman AOM - 2 nW 674nm was passing through with AOM off, 6.5 uW with AOM on - after adjustments much less than 2 nW
    C1 peaks:

    C2:

    C4:


    OD4/23dB/50ms spectra with 0.5Hz step
    C1 peaks:

    C2:

    C4:


    OD2/23dB/5ms spectra with 5Hz step - after PDH-lock modifications..
    C1 peaks:


    Try to measure PDH-signal sensitivity. Siglent siggen modulation on FALC input
    50mVpp (50r output mode), 10 mHz, comb data from ca 10:09:30-> stop modulation 10:17
    Modulation seems to give changing 25MHz peak-height on PDH photodiode signal? (why?)
    NOTE: we wants steepest possible PDH error-signal slope, so fixed voltage variation should give smallest possible frequency variation as seen by comb

    Matlab fit to data gives 56.19 (52.81, 59.57) Hz sine-wave amplitude for comb data
    f_peak-peak_IR = 2*56.19 = 112.38
    PDH_coefficient = 50 mVpp / (2*112.38 Hzpp) = 0.22 mV / Hz at 674nm

    Add mini-circuits amplfier on PDH photodetector (with 10dB attenuator before amplifier)
    100mVpp modulation 10 mHz, from ca 10:58. Sine-fit gives 18.92 Hz (1348nm) amplitude
    PDH_coefficient = 100 mVpp / (2*2*18.92) Hzpp = 1.32 mV / Hz at 674 nm
    improvement is 1.32/0.22 = 6-fold
    mini-circ nominal gain +28 dB, attenuation 10 dB -> +18 dB corresponding to 7.9-fold amplitude gain
    Adjust PDH mixing angle, set-point 300deg, Toptica software rounds to 295.7 degrees - do we now get close to 8-fold more sensitivity?
    200mVpp 10 mHz modulation from ca 11:19 - sine fit to comb data gives 34.22 Hz amplitude
    PDH_coefficient = 200 mVpp / (2*2*34.22) Hzpp = 1.46 mV / Hz at 674nm (11% improvement over previous LO-angle result 1.32 mV/Hz)
    compared to original 1.46/0.22 = 6.6 fold improvement

    PDH settings with no mini-circ amplifier on PDH-detector
    ULI: fastest, SLI: 3, FLD: 7, FLI: 8
    main-gain ca 2.5
    no attenuator on ECDL current control cable
    
    PDH settings with mini-circ amplifier (10dB attenuator after detector, before mini-circ amplifier)
    16dB attenuation on ECDL current control
    ULI: 100V/s/V, SLI off, FLD 6, FLI: 8
    main-gain 4
    
    

    B-counts 5815 / 22 in 1000ms
    B-beam detuning scan: repeat, repeat 100ms gate

    uMotion after optimization:

    30 iterations uMotion optimizer:
    V = [-6.5000 0.2874 24.3318 56.5690 ]


    uMotion: 8%, 7.7%, 4.9%

    ion stored and bright

    2022-02-24

    OD2/23dB/5ms spectra with 5Hz step
    C1 peaks:

    C2:

    C4:


    locking with pos1 seems unstable..
    OD2/23dB/5ms: FALC SLI pos2 (2.4 kHz - 140 kHz):
    +C1 (ion went dark.. recovers with all beams and detuning scan)


    OD2/23dB/5ms: FALC SLI pos3 (1.1 kHz - 65 kHz):
    +C1


    OD2/23dB/5ms: FALC SLI pos4 (500 Hz - 30 kHz):
    +C1


    OD2/23dB/5ms: FALC SLI from pos6 (110 Hz - 6.5 kHz) to pos5 (240 Hz - 14 kHz):
    +C1


    OD2/23dB/5ms: increase FALC ULI speed from 500 to 2000 (V/s)/V
    +C1


    OD2/23dB/5ms: increase FALC ULI speed from 100 to 500 (V/s)/V
    +C1


    OD2/23dB/5ms: increase FALC ULI speed from 20 to 100 (V/s)/V
    +C1


    OD2/23dB/5ms: new GND-wire from helres alu-mount (drive side) to table
    +C1


    OD2/23dB/5ms: reconnect yellow/green GND-wire again.
    +C1


    OD2/23dB/5ms: disconnect yellow/green GND-wire coming from trap?? (bolted to optical table, corner)
    +C1


    OD2/23dB/5ms: back to -24.2dBm EOM drive. disconnect USB3 camera at computer end
    +C1

    -C1 with camera connected again:


    OD2/23dB/5ms: +3 dB more EOM drive -> higher PDH-lock overall gain (2nd order servo-bump in beat-note visible)
    +C1


    OD2/23dB/5ms: Tickler cable connected to artiq DDS ch0, with 20dB fixed attenuator. (was 50R terminated)
    +C1


    OD2/23dB/5ms: trap dispenser wires floating(shorted together), outer shield GND-wire removed?
    C1 peaks:


    OD2/23dB/5ms: Fourier limit 80 Hz AOM
    +C1 FWHM 222 Hz
    -C1 FWHM 201 Hz
    C1 peaks:

    C2:

    C4:



    OD2/18dB/2.5ms:
    +C1 Center =  75.85317005690537, Height =  32.4, FWHM =  249
    -C1 Center =  75.82592917944645, Height =  38.4, FWHM =  230
    C1 Linecenter =  75.83954961817591, B1 estimate =  4.854002616997453
    
    C1 peaks:

    C2:

    C4:


    17 local uMotion: 1.3%, 3.3%, 3.2%


    B-beam detuning scan:


    Add Waveplate on trap breadboard, to mazimize beam through cube and to B-beam? (plate removed again..)

    Detuning-scans with beams A (fit fails?), B, C, repeat A:


    ca15:50local uMotion: 0.4%, 0.8%, 2.4%

    B-beam Detuning-scan with OD1:
    Counts: 6764 / 73

    B-beam detuning scan, OD0:
    Counts at 208MHz: 27432 / 183

    Changes to 422nm waveplate and fiber..

    RID20909 B-beam PMT cuonts, 100ms gate, when fiber-temperature modulated with 10 mHz ramp
    Full power/OD0, variation ca 2400-3000, or 20 %


    RID20906 Repeat OD2/18dB/2.5ms spectrum (now with B-beam, not sure if previous result was B or C...)
    800 Hz span, 20 Hz step (usual 5 pts/FWHM stepsize would be 32 Hz)
    Now with OD0 for cooling laser, 10-fold more power!? PMT threshold still 15 counts/10ms - seems to work?
    2.5ms probe gives 160 Hz AOM Fourier limit
    
    comb linecenter 75 839 525 (start of run)
    +C1 Center =  75.85317034677945, Height =  32.1, FWHM =  225
    -C1 Center =  75.82592562351964, Height =  22.0, FWHM =  273
    C1 Linecenter =  75 839 548, B1 estimate =  4.854688
    +C2 Center =  75.86670286208648, Height =  25.2, FWHM =  244
    -C2 Center =  75.81237696790912, Height =  28.7, FWHM =  268
    C2 Linecenter =  75 839 540, B2 estimate =  4.855491
    +C4 Center =  75.90749916601803, Height =  26.9, FWHM =  233
    -C4 print:Center =  75.77159753708638, print:Height =  26.73811334304485, print:FWHM =  0.0002828593188409724
    C4 Linecenter =  75.8395483515522, B4 estimate =  4.855513
    comb linecenter 75 839 533 (end of run)
    
    C1 peaks:

    C2:

    C4:


    B-beam counts with full 422nm power - no temperature modulation on fiber, waited over lunch (ca 1 h) without temperature modulation on fiber
    Still quite large count variations. Time-series ca 250s long, 500 pts of 100ms gate (but ca 400ms dead time, 0.5s/pt*500pts = 250s)


    RID20891 repeat with B-beam (still no changes on laser breadboard)
    PMT count varies from 900 to 500

    Try heating/cooling fiber between 422-laser and trap breadboard. Large changes in ion fluorescence from 700 counts down to 2-300 counts
    temperature modulation with peltier, maybe +15..+35C
    run1: no changes to waveplate on laser breadboard. Note this was beam C (intention was beam B...)


    PDH detector signal after adjustments yesterday. Measured through 11 dB coupler on photodetector output. No Mini-Circuits amplifier in use. Laser locked.
    Distance to servo-bump ca 750 kHz - same as seen on comb beat note


    OD2/18dB/2.5 ms: 800 Hz span, 20 Hz step, 41pts/peak
    comb linecenter 75 839 513 (start of run)
    Fourier limit 160 Hz AOM
    +C1 Center      75 853 134, Height =  28.1, FWHM =  244
    -C1 Center      75 825 882, Height =  23.2, FWHM =  344  (more activity in lab - disturbance or not??)
    C1 Linecenter   75 839 508, B1 = 4.856092
    +C2 Center      75 866 646, Height =  41.4, FWHM =  254
    -C2 Center      75 812 347, Height =  9.4,  FWHM =  277
    C2 Linecenter   75 839 497, B2 = 4.853101
    +C4 Center      75 907 385, Height = 25.6,  FWHM =  237
    -C4 print:Center =  75.77155583401284, print:Height =  8.175377539725075, print:FWHM =  0.00021086636619431357
    C4 Linecenter   75 839 471, B4 = 4.852934
    comb linecenter 75 839 476 (end of run)
    
    C1 peaks:

    C2:

    C4:



    ID20887 rabiflop on +C1, tau_pi maybe 2.5 ms?

    OD2/18dB/4ms
    comb linecenter 75 849 516 (start of run)
    +C1  =  75.853119 
    
    C1 peaks: probe too long??


    OD0/18dB/400us sanity check..
    comb linecenter 75 839 511
    +C1 Center      75 853 169, Height =  31.5, FWHM =  896
    -C1 Center      75 825 886, Height =  55.7, FWHM =  921
    C1 Linecenter   75 839 527, B1 = 4.861601
    +C2 Center      75 866 635, Height =  49.5, FWHM =  1033
    -C2 Center      75 812 396, Height =  36.1, FWHM =  994
    C2 Linecenter   75 839 515  B2 = 4.847667
    +C4 Center      75 907 418, Height =  41.8, FWHM =  891
    -C4 Center      75 771 562, Height =  30.0, FWHM =  988
    C4 Linecenter   75 839 490, B4 = 4.853862
    comb linecenter 75 839 517 (end of run)
    
    C1 peaks:

    C2:

    C4:



    10:12local uMotion: 1.6%, 0.8%, 3%

    B-beam detuning scan:

    B-counts: 5624 / 54 in 1000ms

    Menlo FiberComb software had crashed again. Restart.

    OD4/23dB/80ms
    +C1 fit to noisy data gives FWHM 178 Hz (similar to 40ms spectrum earlier)
    -C1 FWHM 165 Hz
    B1 estimate =  4.851566097194596
    
    C1 peaks:


    OD4/18dB/40ms spectrum
    Initialized with last B-estimate 4.871905 uT - but now B seems lower?
    peaks offset by maybe 1/4th of span, 150 Hz optical, so B should be adjusted ca -30 nT (assuming ca 5 Hz/nT for C1) -> 4.84 uT
    
    +C1 fit to broad noise peak, Center =  75 853 123, FWHM 170 Hz AOM
    -C1  Center =  75.82588931631606, FWHM 163 Hz
    B1 estimate =  4.852793995323336
    B2 estimate =  4.852588307569552
    
    C1 peaks:

    C2:

    C4:


    uMotion:

    laser drift settled ca 04 local (12 hours after doors closed)
    ca 06:30 local uncompensated drift 1.7 mHz/s at 674nm

    2022-02-23

    OD4/23dB/80ms spectrum
    C1 peaks:

    C2:

    C4:


    OD4/18dB/40ms spectrum
    C1 peaks:

    C2:

    C4:


    18:00 OD4/18dB/40ms spectrum - restart
    C1 peaks:

    17:30 OD4/18dB/40ms spectrum - still big laser drift..
    C1 peaks:

    ca 16 local close cavity door
    OD4/18dB/40ms spectrum
    C1 peaks:

    Time-stamps for later analysis, 500k each (now GND-wire removed): 852, 853, 854

    B-beam detuning scan:

    remove GND-wire from camera objective (caused uMotion problems previously):
    now uMotion:


    collect photon time-stamps to analyze ion motion?
    RID20835 B-beam 100kphotons in 13s
    20836, 500k photons
    20837 repeat 500k photons
    20838 repeat 500k photons
    all B-beam!


    remove amplifiers from PDH-detector chain - seems mixed error-signal was saturated?
    Tune FALC settings..
    OD4/18dB/40ms spectrum
    C1 peaks:

    C2:

    C4:



    Fiber-noise signals
    Yellow: VCO output to AOM, at 178.5 MHz, when fiber-noise servo is locked.Shows spurs at 43Hz (-45 dB down from carrier) and 71 Hz (-52 dB down from carrier)
    Red: reference signal from ARTIQ DDS at 357 MHz. Spur at 88Hz (-75 dB down from carrier)


    OD4/18dB/40ms spectrum
    C1 peaks:

    C2:

    C4:


    PDH signal as measured by photodiode, around 25 MHz EOM modulation frequency. Laser locked.
    Measured through 11 dB coupler
    
    Yellow: signal from photodetector
    Red: original configuration with one mini-circ +28 dB amplifier (some compression maybe?)
    Blue: signal after various tests and improvements - both high-pass and low-pass filters to keep 25 MHz signal but attenuate 50MHz and close to DC
    SHP-20 high-pass filter
    BPL-30 low-pass filter (seems strong 50 MHz component was causing amplifier saturation before this filter was added, after photodetector, before mini-circ amplifier)
    

    PDH error signal
    servo-bumps at ca -50 dBm (measured at photodetector output, with 11 dB coupler) -> signal at -39 dBm
    
    mini-circ ZFL-500LN gain at +15V is +28 dB
    
    with two minicircs PDH photodetector spectrum doesn't seem any higher??
    
    SHP-20 high-pass filter tried on photodetector output, removes 0-1 MHz frequency content of photodetector.
    SHP-20 has large group delay at 25 MHz (not ideal?), Toptica PDH mixing angle ca 190 degrees now?
    27.87ns group delay, period is 1/25MHz = 40 ns -> 251 degrees !?
    
    Try to increase PDH error signal, filters to block DC-2MHz and also 50 MHz!
    
    

    OD4/18dB/40ms spectrum
    C1 peaks:

    C2:

    C4:

    B-beam detuning scan
    B-counts 6236 / 22.1
    (fit fails?)

    07:30 uMotion 1.3%, 0.8%, 2.5%

    ion dark. recovers during detuning scan, with clearout, all beams, and 405n on.

    2022-02-22

    OD4/28dB/160ms spectrum
    C1 peaks:


    OD4/23dB/80ms spectrum
    C1 peaks:

    C2:

    C4:


    Change to -10dBm input level on Toptica (win program) - no visible changes in PDH error signal 50Hz peak or in beat-note spectrum?

    vibration spectrum on stone shows 87.5Hz peak
    OD4/18dB/40ms spectrum
    C1 peaks:


    Try mini-circ isolation transformers: at EOM close to cavity, on PDH photodiode close to Toptica DLC-box mixer input
    PDH error-signal shows clear (>20dB) 50Hz peak

    OD4/23dB/80ms spectrum
    C1 peaks:


    OD4/18dB/40ms spectrum
    C1 peaks:


    OD0/28dB/1600us spectrum
    +C1 FWHM 292 Hz
    -C1 FWHM 249
    B1 estimate =  4.867662494842605
    +C2 268 Hz
    -C2 268 Hz
    B2 estimate =  4.870678084197625
    +C4 286 Hz
    -C4 289 Hz
    B4 estimate =  4.871904886216983
    
    C1 peaks:

    C2:

    C4:

    FALC settings 2022-02-22 ca 14 local:
    ULI 20 (V/s)/V
    SLI pos 7: 50 Hz - 3 kHz (lock first with FLI, then turn on SLI)
    FLD pos 7: 190k - 1MHz
    FLI pos 8: 6k5 - 80kHz
    output 20 dB attenuator
    main-gain trimpot ca "4"
    
    Toptica
    EOM drive level -24.2 dBm
    Input level 0 dBm
    Lock level +0.073 V
    Phase 12.9 degrees
    

    FALC settings currently in use:
    XSLI off
    SLI: off
    FLI: DIP 8: 6.5 kHz - 80 kHz
    FLD: DIP 7: 190 kHz - 1 MHz
    ULI: DIP 1+6: 5 (V/s)/V
    10 dB attenuator at main output

    B-beam detuning scan
    B-counts 7968.2 (684.4) / 75.9 (11)



    afternoon 15:45 repeat uMotion: 0.6%, 0.7%, 1.04%

    (morning) uMotion 0.2%, 0.7%, 0.5%

    08:44 ion still stored and bright

    2022-02-21

    With -36 dBm EOM drive
    With -10 dBm ?
    With 0 dBm setting get ca 0.91 kHz/V
    With +10 dBm setting get ca 0.5 kHz/V

    -29 dBm EOM drive Data from 17:11
    -27 dBm EOM drive from 17:13 (lower modulation seen by comb)
    -30 dBm from 17:16
    -32 dBm from 17:17 (much higher modulation seen by comb!?)
    -34 dBm - need to change FALC settings.. 20mVpp modulation (34 gives larger modulation than 32)
    -36 dBm? similar to -34 (now 10mVpp modulation)
    -38 dBm? (better?)
    -36 dBm data with 10mVpp modulation 10mHz frequency from 17:58->18:06 (todo: analyze data)

    switch to 50mVpp modulation
    -24.2 dBm might give slightly higher servo bumps - but comb shows higher modulation with -27 dBm
    Optimal EOM modulation depth? (previous setting -27.2 dBm from Toptica)

    valid data ca 16:58 - 17:05
    Input level setting -10dBm from ca 16:47local, now with 20dB attenuator on FALC output

    modulation ca 45.6 Hz seen by comb
    Modulate FALC with 100 mVpp and 10 mHz frequency. valid data from ca 16:11 local (Note siglent 50ohm output mode, FALC input impedance is 50 Ohm)
    Need to turn down gain-trimpot to ca 2/10 value to avoid exessive servo-bumps
    changed Toptica PDH-mixer input sensitivity from +10dBm to 0 dBm -> increased gain, should see as (3-fold?) increased PDH error signal sensitivity?

    OD0/28dB/1600us spectrum:
    Fourier limit 250 Hz AOM: 
    +C1 267 Hz
    -C1 297 Hz
    +C2 269 Hz 
    -C2 246 Hz
    +C4 292 Hz
    -C4 298 Hz
    B4 4.87268 uT (was typo "2.87.."?)
    
    C1 peaks:

    C2:

    C4:


    OD0/18dB/400us spectrum:
    Fourier limit 1 kHz AOM: experiment +C1 1119Hz, -C1 957 Hz AOM
    C1 peaks:


    get ca 2 V/kHz (at 674nm) for PDH error signal sensitivity with current settings
    from 13:47local -> reduced amplitude to 50mVpp
    valid data starts ca 13:34local
    100 mVpp and 10 mHz? should give 50s at lower frequency and 50s at higher?
    raise amplitude to 100 mVpp - now see something in comb data?
    Try 10mVpp square wave at 50 mHz into FALC noninverting input - to get PDH error signal units V/Hz?
    PDH settings: Toptica EOM drive -27.2 dBm, PDH input level +10dBm, lock level 0V, LO phase 347.1degrees.

    ID20788 B-beam detuning scan
    B-counts: 7111.3 (1100.8) / 81.7 (10.6) in 1000ms

    uMotion after optimization: 1.7%, 1.9%, 1.6%

    Leave 30 iterations, 50 kphotons, of uMotion optimizer over lunch:
    C = 0.0114 0.0165 0.0239 RMS = 0.0312
    V = [-6.5000 0.2458 24.2582 56.5455 ]

    Restart scan:
    C1 peaks:


    OD4/18dB/40ms spectra:
    modify FALC-gain (lower servo bumps) during +C2 scan
    C1 peaks:

    C2 peaks:


    ID20777 B-beam detuning scan:
    B-counts 7657.3 (179.1) / 74.7 (8.6) in 1000ms


    optimized uMotions with more photons: 2.4%, 2.9%, 2.0%

    15 iterations of Nelder-Mead:
    C = 0.0400 0.0317 0.0192 RMS = 0.0546 at voltages V = [-6.5000 0.2491 24.2236 56.4999 ]


    now uMotion: 19%, 1.7%, 25%

    rough B-minumum at V12cm , V12diff, V34cm, V34diff = [ -6.5, 0.18, 23.7505, 56.786]
    ID20767 repeat with GND-wire disconnected again, OK V-shape with 2% min contrast

    ID20766 B-contrast vs. V12diff, camera objective foil GND-wire connected to camera-breadboard. Optimum moves a lot - out of scan range!

    ID20765 B-contrast vs. V12diff scan, disconnect GND-wire from camera objective foil (does it help or not?)

    NOTE: with camera-objective-foil GND-wire disconnected, 14.424 MHz almost disappears from beat-note! -60 dBm (close to noise-floor of beat-note)
    14.424 RF signal in beat-note now ca -50 dBm (RBW 30kHz) - compare to -36.5 dBm yesterday??
    B-counts quite low? 6998 / 64 in 1000ms
    09:45 uMotion: 16%, 22%, 20%


    Back to OD1 filter for 422nm
    09:43 ion recovers with 422nm full power (405nm and clearout still kept on)
    all beams, 405nm and detuning scan -> still dark
    Connect USB3 camera.
    09:30 ion dark

    2022-02-20

    peaks with various new FALC settings tried..
    C1 peaks:

    C2 peaks:

    C4 peaks:


    Trap RF peak in beat-note spectrum ca -36.5 dBm (RBW 30 kHz)
    
    FALC settings from 2022-02-19
    SLI pos 6: 110 Hz - 6.5 kHz
    FLD pos 7: 190 kHz - 1 MHz
    FLI pos 8: 6.5 kHz - 80 kHz
    ULI pos 1on, 4on: 100 (V/s)/V   Fine-1 ARC-factor -1 V/V
    

    B-beam detuning scan
    B-counts at 208MHz: 5568 (92) / 17.5 (3.6)

    14:30 uMotion: 8.8%, 8.5%, 7.7%


    B-contrast with 10min intervals:

    after optimization:

    15 iterations of uMotion-optimizer, 20kphotons:
    V = [-6.5000 0.0698 23.7505 56.7860 ]


    15k measurement with 10min interval to see time-dependence?
    id2722, 23, 24, 25, 26, 27, 28, 29 - move scan range up, 30kphotons: 30, 31, 32


    Sharper V-shape, much better phase-fit before and after uMotion minimum - disturbing RF field mostly removed now?

    id20721 15 kphotons
    id20720 repeat with 120kphotons
    id20719 repeat with 60kphotons
    id20718 repeat with 30kphotons, -0.2..0.2 V12diff scan
    id20717 B-contrast vs. scan of V12diff, 30kphotons

    uMotion now:

    Ion still stored and bright

    2022-02-19

    This improves significantly when disconnecting USB3 camera at computer end. Camera still installed as normal and cable runs from trap breadboard to computer-cabinet as usual.
    B-contrast vs V12diff scan shows blunt minimum and lowest contrast ca 10% - indicates out-of-phase 14.424MHz field seen by ion?

    Schedule B-beam uMotion once per hour - is uMotion constant?
    ID20701 - ID20711 - scheduling failed..


    uMotion after optimization:

    ID20695 uMotion scan 4x4x4 grid, 100kphotons

    Clock-laser unlocked..
    Add FALC SLI gain also, with 20 dB attenuator on output and gain-trimpot ca half-way.
    40ms spectra
    C1 peaks:

    C2 peaks:

    C4 peaks:


    after 60 iteration optimization:

    60 iterations:


    200k uMotion after optimization:

    Nelder-Mead again?
    V = [-6.5000 0.0670 23.8056 56.7755 ]
    C = 0.1303 0.1222 0.1502 RMS = 0.2334
    30 iterations


    200k uMotion after optimization: 10.4%, 7.7%, 8.9%

    Optimal voltages (Common mode, diff): [ 0.06701685 23.80559423 56.77547893]
    ID20680 is old uMotion optimization, 4x4x4 grid, 50kphotons. duration 64pts*(9s + 8s + 7s)= 26 min

    200k uMotion after optimization:

    20 iterations with Adaptive option on?


    200k uMotion after optimization:

    30 iterations Nelder-Mead with initial simplex size=1.0, n_photons=50k
    finds reasonable optimum - but further iterations give worse results? (or true uMotion nulling voltages change over time?)


    repeat same spectrum as yesterday OD4/18dB should give 40ms ?
    Fourier limit 10Hz AOM. step is 2Hz, span 300Hz AOM
    C1 peaks:

    C2 peaks:

    (stop at -C2)

    ID20660 B-beam detuning scan with this micromotion:
    B-counts: 9401.3 (112.8) / 77.1 (7.6) in 1000ms


    uMotion 100k: 5.0% 2.5% 2.4% (compare to 20k optimizer C-vector below)

    Disconnecting camera USB3 from computer helps to get lower contrasts??
    run Nelder-Mead optimizer, 10 iterations, 20kphotons
    C = -0.0467 -0.0207 -0.0507 RMS = 0.0720 at voltages V = [-6.5000 0.0808 23.9465 56.7529 ]

    RF-drive   min B-contrast when scanning V12diff
    9.9        17 %   id20643
    9.8        11 %   id20644
    9.7        10 %   id20645
    9.6        10 %   id20646
    9.5        11 %   id20647
    9.4        13 %   id20648
    9.3        16 %   id20649
    9.2        17 %   id20650
    

    2022-02-18

    OD4/18dB should give 40ms ?
    (run restarted once, so +C1 has some overlapping old datapoints also..)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    OD0/28dB/1600us 50 Hz step
    +C1 75 853 327, height 38.6, FWHM 302, Fourier limit 250 Hz AOM
    -C1 75 826 047, height 40.8, FWHM 281
    C1 75 839 687, B1 4.861081
    +C2 75 866 894, height 43.5, FWHM 274
    -C2 75 812 486, height 33.4, FWHM 345
    C2 75 839 690, B2 4.862868
    +C4 75 907 746, height 32.2, FWHM 291, (far right in scan)
    -C4 75 771 625, height 36.2, FWHM 290
    C4 75 839 685, B4 4.863354
    comb linecenter 75 839 682
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    OD0/23dB/800us 100 Hz step
    +C1 75 853 363, height 40.9, FWHM 472, Fourier limit 500 Hz AOM
    -C1 75 826 030, height 48.2, FWHM 444
    C1 75 839 697, B1 4.870498
    +C2 75 866 697, height 40.0, FWHM 498
    -C2 75 812 487, height 36.2, FWHM 461
    C2 75 839 691, B2 4.862817
    +C4 75 907 728, height 38.7, FWHM 499
    -C4 75 771 642, height 41.7, FWHM 419
    C4 75 839 690, B4 4.862455
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    5dB AOM attenuation should double tau_pi?
    Sanity check: OD0/18dB/400us peak search:
    comb linecenter 75 839 687
    +C1 75 853 365, height 45.9, FWHM 921
    -C1 75 826 068, height 41.1, FWHM 921 
    C1 75 839 717, B1 4.86392
    
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    uMotion at optimal voltages:

    Optimal voltages (Common mode, diff): [ 0.06218931 23.76733537 56.82001601]
    ID20393 Try old uMotion optimization in 4x4x4 grid, 10kphotons

    1st run Nelder-Mead 5 iterations 50kphotons: V = [-6.5000 0.0669 23.7630 56.8506 ] C = -0.0879 -0.1047 -0.0673 RMS = 0.1524

    around 15-local uMotion results now seem more reasonable again (but nothing major changed in lab?)
    Try repeated B-beam contrasts - previous data showed high contrast peaks with 2-3 min period?? can't repeat this result maybe old data had voltage-scans ongoing also

    B-contrast while scanning V12diff between -0.2 and +0.2 V in 21 steps:
    no good minimum, always at least 10% contrast. optimum around +0.02 V12diff
    10 iterations of Nelder-Mead optimizer, 50 kphotons/pt
    C = -0.1473 -0.1153 -0.1483 RMS = 0.2388 V = [-6.5000 0.0273 23.7568 56.5703 ]

    Test B-beam repeatability of uMotion results, between 14% and 21% in 25 measurements


    uMotion after secular measurements: 24%, 22%, 27%

    Seculars 2022-02-18, 10Vpp RF-drive, Vbias = [-6.5 0.0273 23.7568 56.5703]
    
    Z 1.939298  (repeat 1.939265 )
    Y 0.951322  (repeat 0.951240 )  camera image blurry but no clear side-side movement, height <0.2
    X 0.903902  (repeat 0.903876 )
    
    Try tickler measurement, prediction [0.9292 0.9751 1.9371]
    RF drive back to 10.0 Vpp (max)
    Z: search 1.93 to 1.94 MHz - resonance close to 1.939
    ID19926 B-beam, 1.938 - 1.941, 26 dB attenuation (+fixed 20dB attenuator on tickler output). 5k photons per point.
    height only 0.2, can use more tickler?
    ID19927 repeat with 23 dB attenuation


    Y search 0.97-0.98 with 15 dB attenuation - no clear resonance?
    search 0.98-0.99 - 0.997 , 0.96 - 0.97
    coarse tickler with 20dB artiq and no fixed attenuation finds X? around 0.904
    with 15 dB attenuation Y possibly around 0.951
    Y-scan 0.945-0.955 with 20+15 dB


    X-scan 0.9-0.91 with 20+15 dB



    Try Nelder-Mead uMotion optimizer:
    1st run contrasts: C = -0.1875 -0.2036 -0.1363 RMS = 0.3085  
    2nd run V = [-6.0000 -0.0728 24.0370 56.7160 C = -0.1515 -0.0956 -0.1246 RMS = 0.2182
    
    Lower RF-drive to 9.7V and run optimization again
    3rd run, 100k photons, n_iterations=20
    V = [-6.0000 0.0375 24.3179 56.8077 ]   C = -0.1603 -0.1043 0.1746 RMS = 0.2590
    
    Try V12cm = -6.5 ?
    4th run, 50k photons, 10 iterations
    V = [-6.5000 0.0625 24.3000 57.3080 ]   C = -0.3672 -0.1260 -0.1409 RMS = 0.4130
    
    Set RF to 9.4V
    5th run, 50 kphotons, 10 iterations
    V = [-6.5000 0.0273 23.7568 56.5703 ]  C = -0.1452 -0.1378 -0.1881 RMS = 0.2747
    
    B-contrast repeatability now slightly better?

    10:44 uMotion now: 37%, 68%, 60%

    Vbias = V12cm , V12diff, V34cm, V34diff = [ -6.0, 0.0, 27.0, 55.0]

    Bias voltage measurements:
    adjust V1 and V2 to zero - while keeping ion in trap.
    
    Setpoint    measured
    V1=0        V1=1.13(1) mV
    V12cm=-2V   V1 = -1.99872(1) V  V2 = 1.99874 V
    
    ramp down V34cm in steps of 2V, and V34diff in steps of 5V -> this keeps ion trapped but camera image is blurry
    Setpoint                  measured
    V34cm=0, V34diff=0        V3 = +0.0026(16) V   V4 = +0.0024(17) V
    V34cm=0, V34diff=55       V3 = -54.9980(16) V  V4 = +55.0038(17) V
    V34cm=27, V34diff=55      V3 = -27.9986 V      V4 = +82.0030(17) V
    
    V12cm = -6                V1 = -5.9989 V       V2 = -5.99937(1) V 
    
    Coil current measurements (mA):
    I1: Reading:  -0.0413, set-point: -0.04080, DMM-measurement:  -0.0384
    I2: Reading: -11.4941, set-point: -11.4939, DMM-measurement: -11.490
    I3: Reading:  -0.7749, set-point: -0.77410, DMM-measurement:  -0.7678
    

    10:10 uMotion: 17%, 17%, 17%


    change to slot10 OD1.0, transmission 10% - normal operating point(?)
    ID19812 is B-beam detuning scan with OD1.0
    B-counts: 5956.4 (583.7) / 21.2 (6) - B-beam power ca 0.12 uW (large variations, hard to measure?)
    (lineshape fit fails - no image?)


    ID19810 B-beam detuning scan with OD0.8
    B-counts 8109.4 (1006) / 24.2 (6.6) - B-beam power ca 0.4uW

    change to slot8, OD0.8, transmission 15.8% - ion dark again..

    ID19806 detuning scan with slot-4 on filter wheel OD=0.4, transmission 39.8%
    B-counts: 22208.1 (368.1) / 50.2 (7.1) in 1000ms (lab lights off) - B-beam power ca 1.4uW

    recovers in a few seconds
    ion went dark again when changing OD-filter on 422nm laser (which caused loss of Rb-lock..)

    ID19804 B-beam full power detuning scan
    B-counts: 29711.4 (625.1) / 131.4 (14.5) in 1000ms - B-beam power ca 2.9 uW (with variations)


    09:47 ion recovers with all beams and full 422nm power (405nm not continuously on, just 1-2 mins after 09:33)
    09:33 try 405nm again.
    continue count monitor
    07:38 all beams and detuning scan tried - ion still dark
    07:30 ion dark, clearout and 405nm on

    2022-02-17

    Leave B-count monitor:


    repeated B-contrast with 100k photons, ca 10s per measurement, then 200k, ca 20s per measurement


    uMotion contrast seem to vary a lot over time, when voltages kept constant.
    Voltage supplies at back of bias-supplies seem OK, +/- 5V, +/-15V, and +/-173 V
    Repeated B-contrast measurements with 10k photons, ca 3s per measurement:


    Try new uMotion optimizaiton with scipy.optimize.minimize( method='Nelder-Mead' )
    
    First run
    C = -0.3262 0.0710 -0.0561 RMS = 0.3386    at voltages [-0.0022, 23.6312, 56,9558]
    
    Second run
    print:C = 0.0606 -0.0654 -0.0422 RMS = 0.0987
    
    

    repeated contrasts
    C = -0.6464 -0.5158 -0.3178
    C = -0.6413 -0.5400 -0.3960
    C = -0.6197 -0.5084 -0.3581
    C = -0.6199 -0.5190 -0.3378
    C = -0.6397 -0.4739 -0.3023
    C = -0.6200 -0.4581 -0.2982
    C = -0.6168 -0.4594 -0.2775
    
    uMotion now: 65%, 57%, 47%

    Optimal voltages (Common mode, diff): [-2.20992619e-03 2.36312417e+01 5.69558080e+01]
    ID19472 is uMotion optimization, 4x4x4 grid, 30kphotons

    10 repeats with 100k photons:
    -6.51, -6.38, -6.81, -5.76, 
    
    10 repeats of B-beam uMotion measurement with 50 kphotons:
    +2.23
    +6.75
    +3.79
    +4.39
    +1.35
    +4.67
    +1.15
    +2.5
    +2.05
    +1.83
    

    Repeated measurement of the same uMotion contrasts - without changing trap voltages:
    
    C = 0.1271 -0.1746 0.0968
    C = 0.1333 0.1077 0.1054
    C = 0.1112 -0.1625 0.0930
    C = 0.1122 -0.1583 0.0853
    C = 0.1252 0.1278 0.1337
    C = 0.0949 0.1459 0.1042
    C = 0.0873 0.1324 0.0782
    C = 0.1021 -0.1509 0.0865
    
    after optimization: 10%, 10%, 11%

    Optimal voltages (Common mode, diff): [ 0.13633055 24.01770899 56.62724697]
    ID19370 new uMotion optimization, 80 kphotons, narrower scan interval around previous optimum

    uMotion after optimization: 5.9%, 6.2%, 4.0%


    Optimal voltages (Common mode, diff): [ 0.1654255 24.15374045 56.59760042]
    ID19362 is new uMotion optimization, 40k photons, V12cm range [0, 0.3] V (previous optimum was close to scan-limit [-0.15, +0.15] V12cm)
    scan intervals: [0, 0.3], [23, 36], [55, 58]

    ID19364 B-beam detuning scan
    B-counts at 208 MHz: 8557.4 (170.6) / 59.8 (8.7) in 1000ms


    uMotion after optimization A 10%, B 12%, C 9%


    Optimal voltages (Common mode, diff): [ 0.14627313 24.06084981 56.63862896]
    ID19359 uMotion optimization - only 20kphotons
    ID19352 is B-beam detuning scan:

    blink-test with 405nm doesn't show clearly 2 levels... looks like one ion
    Lowering RF voltage to 6V doesn't show two ions clearly
    Problem could be multiple ions in trap?

    2022-02-16

    Leave B-beam count monitor:


    Try FALC lock without slow ULI branch?
    OD2/28dB/17.5 ms:
    C1 peaks:

    C2 peaks:

    C4 peaks:


    uMotion after optimization:

    Optimal voltages (Common mode, diff): [5.54709299e-02 2.38043781e+01 5.66646017e+01]
    ID19322 is new uMotion optimiziation try. 4x4x4 grid, 100kphotons
    ion went dark during optimization? recovers with 405nm and detuning-scan (all beams)
    ID19315 uMotion optimization 4x4x4 grid

    OD2/28dB/17.5ms spectrum:

    OD0/28dB/1.75ms spectrum:
    C1 peaks: +C1 FWHM 345 Hz (Fourier limit 229Hz), -C1 FWHM 333 Hz


    uMotion with 208 MHz detuning: 14%, 12%, 8%


    B-beam detuning-scan again. B-beam 422nm ca 0.4 uW
    B-beam counts in 1000ms 9198.4 (224.4) / 21.0 (2.7) (lab lights off)

    Back to 208MHz detuning
    B-beam detuning-scan after optimization

    uMotion after optimization: 13%, 15%, 18% - not so great!?

    Optimal voltages (Common mode, diff): [4.89253644e-02 2.37689955e+01 5.67042011e+01]
    ID19297 is uMotion minimization 4x4x4 grid, 100kphotons

    uMotion (now detuning -15 MHz!) A 10.5%, B 11.7%, C 9.4%


    remove attenuator on FALC-output -> now servo-bumps as before, 0.35 V on transmitted detector
    C1 peaks:

    C2 peaks:

    C4 peaks:

    Repeat OD0/28dB/1.75ms spectra (PDH-lock not optimized for lower signal levels, no servo-bumps, but does it matter?)
    +C1 FWHM 329 - same as before?

    Try ND 1.0 filter after EOM, before cavity?
    Thorlabs NE510B-A (A-coat ok for 674nm?)
    Now ca 24 uW before filter, and 3 uW after filter. Transmitted photodiode shows 0.5 V (down 5-fold from 2.5 V previously)

    ID19289 OD0/28dB/1.75ms spectra?
    comb linecenter 75 839 357
    +C1 75 852 918, height 30.5, FWHM 311 (Fourier limit 133 Hz AOM!?)
    -C1 75 825 805, height 32.3, FWHM 367
    C1  75 839 362, B1 4.8313 uT
    +C2 75 866 379, height 39.9, FWHM 271
    -C2 75 812 331, height 29.3, FWHM 354
    C2  75 839 355, B2 4.8306 uT
    -C4 75 906 925, height 32.5, FWHM 328
    +C4 75 711 790, height 31.1, FWHM 287
    C4 75 839 357, B4 4.82813 uT
    comb linecenter 75 839 357 (end of run)
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID19287 OD0 rabiflop on +C1 (AOM att 28 dB), tau_pi maybe 1.75 ms

    ID19285 OD0 rabiflop on +C1 (AOM att 23 dB), tau_pi maybe 900us

    ID19285 OD0 rabiflop on +C1 (AOM att 18 dB)


    ID19284 OD0/18dB/400us spectra, sanity check?
    past B-estimate 4.827 uT
    comb linecenter 75 839 356
    +C1 75 852 912, height 44.3, FWHM 974
    -C1 75 825 828, height 48.9, FWHM 1030
    C1 75 839 370,  B1 4.826 uT
    
    C1 peaks:

    B-fiber 674nm: 7.2 uW before fiber, 0.7 uW towards ion from B-objective
    B-fiber 422nm: 8.7uW before fiber, 0.5 uW towards ion from B-objective

    Try larger 422nm detuning -15 MHz, AOM = 203 MHz?

    ID19268 is B-beam detuning scan
    B-counts at 208 MHz: 8690.2 (507.4) / 54.8 (5.1) in 1000ms


    OD5/90ma spectra (no changes since yesterday?)
    C1 peaks:


    uMotion after optimization: 8.7%, 6.7%, 9.5%


    Optimal voltages (Common mode, diff): [9.97083474e-03 2.36976634e+01 5.67785507e+01]
    ID19260 uMotion optimization 4x4x4 grid, 100kphotons

    uMotion: !4.7%, 14.9%, 14.5%

    B-counts 80 (9.2) / 0.5 (0.7) in 10ms
    0757 recovers with detuning scan
    all beams, clearout, 405nm
    0754 ion dark

    2022-02-15

    C1 peaks:

    OD5/90ms spectra:
    C1 peaks:



    Testing various FALC settings..not sure what signal to optimize?
    Moglabs application note FIG13 has log-log plot of transmitted power PSD - but can't get similar slope with lower noise at lower frequencies?

    OD1+OD3=OD4/30ms spectra:
    Strongest peak in FALC-monitor now at 155.6 Hz
    C1 peaks:


    ID19239 OD2/2ms spectra
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID19237 OD2/4ms spectra, updated B estimate to center peaks better?
    +C1 peak height worse with 4ms compared to 3ms??
    C1 peaks:

    C2 peaks:

    C4 peaks:



    ID19235 OD2/3ms spectra:
    +C1 75 852 910, height 23.8, FWHM 241 (Fourier limit 133 Hz AOM)
    -C1 75 825 810. height 25.6, FWHM 247
    C1 75 839 360, B1 4.8288398
    +C2 75 866 369, height 31.9, FWHM 203
    -C2 75 812 362. height 26.5, FWHM 239
    C2 75 839 365,  B2 4.8269495
    +C4 fit fails, peak to right of scan range 
    -C4
    C4 
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    OD2 filter, try rabiflop on +C1 to find Tau_pi?
    ID19234 is rabiflop on +C1, tau_pi maybe 3ms

    ID19230 OD0/400us peaks:
    comb linecenter 75 839 353 (start of run)
    
    +C1 75 852 937, height 41.1, FWHM  934
    -C1 75 825 778, height 40.2, FWHM 1060
    C1  75 839 274  B1 4.83927 uT   (old estimate was 4.80587)
    +C2 75 866 377, height 46.4, FWHM  972
    -C2 75 812 294, height 43.6, FWHM 1042
    C2  75 839 336, B2 4.833757
    +C4 75 906 947, height 48.9, FWHM 1018
    -C4 75 771 707, height 42.5, FWHM  992
    C4 75 839 372   B4 4.8286595
    
    comb linecenter 75 839 356 (end of run)
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID19225 C1 peaks:
    abort scan - Zeeman AOM aperture still wrong?

    Adjust Zeeman AOM aperture - some 674nm was previously reaching ion when AOM off also!?

    ID19218 OD5/90ms spectra with SLI+FLI on FALC
    C1 peaks:

    Adjust FALC gains and offsets to use both SLI and FLI integrators - this should raise gain at low frequencies.
    Using SLI requires input offset to be trimmed with SLI on.
    Using SLI also requires ULI-offset to be trimmed correctly.
    

    PSD of f1 beat-note from 14.2 evening and night - -1floor motor now at lower 35Hz? VFD-setting
    peak at 0.188782 Hz has disappeared
    peaks at 0.0913 and 0.454 Hz remain

    ID19215 is OD5/90ms peak search, initialized with comb linecenter and old B estimate
    sudden +40Hz jump in cavity frequency when work started in lab? (or just coincidence??)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    uMotion: A 8.7% in 59s, B 9.3% in 34s, C 6.3% in 18s

    9:40 recovers during detuning scan! was dark since ca 21 local yesterday..
    ion doesn't recover with full 422nm power and 405nm - lost?

    relock fiber-noise servo, Siglent SA measures 178.507 MHz (not locked to maser-reference) - seems ok
    90ms spectra yesterday did not work - probably because fiber-noise servo was not locked?

    2022-02-14

    all beams, clearout on, 405nm tried for ca 5 mins

    ca21 ion dark

    C1 maybe 50 Hz wrong B-field, so increase by 10 nT? Wider scan to see everything?
    C1 peaks:


    ID19198 is OD5/90ms peak-search
    around 145/201 points into +C1 scan, increase FALC gain to get servo-bumps as before. Why background 'continuum' of counts, no clear carrier? Can we have clearly higher B-field than before?
    Search initialized with comb-linecenter and old B-estimate 4.805872 uT
    Coil currents not changed: -0.0413, -11.4941, -0.7749 mA
    C1 peaks:


    Try to start OD5/90ms clock-run - doesn't seem to find peaks?

    Comb data from weekend
    Clear (aliased?) peaks in 1 S/s comb data at 0.0913 Hz, 0.188782 Hz, 0.454 Hz - compare to new data when -1 floor motor at lower frequency?
    beat-note sum shows ca +3 uHz bias - probably related to AD9912 DDS tuning word error - new data with corrected actual tuning word.

    DDS mixer frequency for comb-counter channel2
    Nominal 215 MHz
    ARTIQ reports tuning word 60517119992791 corresponding to 214999999.999999850988
    if this is true, difference between nominal and actual -0.000000149012 Hz
    
    DDS_output = 1 GHz * FTW / 2^48
    
    However past AD9912 measurements show tuning-word is always even!
    With 60517119992791-1 we get  actual-nominal= -0.000003695488
    with 60517119992791+1 we get  actual-nominal= +0.000003397465
    
    Comb data f1+f2 sum from weekend shows a sum of:
    Saturday +0.00333  mHz 
    Sunday   +0.003892 mHz
    
    beatnote f1 ca -15.5 MHz
    beatnote f2 ca +234.5 MHz
    Mixing frequency for f2 215 MHz -> SDR measures 234.5-215 = 19.5 MHz
    
    SDR computes f1+f2 sum as:
    f1 + f2 + 215 MHz
    we observe +3 uHz error here, so actual mixing frequency is lower than nominal: FTW-1 row above!?
    FTW = 60517119992791-1 -> mixing-frequency is 214999999.999996304512
    Try using this in SDR data processing from now on! 
    

    FALC monitor still shows ca 20 dB peak at 50 Hz
    Blue trace is vibration on lab floor - -1 motor set to 35 Hz today, no clear peak?
    Green trace is FALC monitor


    New AHM3 model from Atomic Bulletin 2022-02
    # ab 2022-01 AHM3
    MJD0 = 59579
    y0 = 3.57713e-13
    y1 = -2.59867e-15
    
    # ab 2022-01 AHM3
    MJD0 = 59609 
    y0 = 2.78848e-13
    y1 = -2.61147e-15
    
    Change in predicted frequency for 59624 (2022-02-14)
    y_AHM3(ab2022-01, 59624) = 2.4077285e-13
    y_AHM3(ab2022-02, 59624) = 2.3967595e-13
    difference  -1.096899999999993e-15   (less than half a day of drift)
    

    With waveplate before B-beam cube-polarizer?


    B-counts 56.9 (6.5) / 0.5 (0.7) in 10ms
    uMotion: A 6.3%, B 2.6%, C 3.7%

    Tune camera/PMT focusing and try again:


    B-beam detuning scan with front lab lights off:


    front lab-lights off: 71.6 (9.6) / 0.5 (0.7) - dark counts improve 16-fold!
    B-counts in 10ms: 80.5 (10.1) / 8.1 (2.9)
    B-beam detuning scan
    B-counts at 208MHz in 1000ms: 8334.9 (163.7) / 816.3 (29.6) - dark counts seem high, lab lights on, foil-covers of camera/PMT bad?


    A 4.7%, B 4.1%, C 5.6%

    Optimal voltages (Common mode, diff): [ 0.08938722 23.88913466 56.64129159]

    A 5.9%, B 5.7%, C 19%
    Try manually? V12cm , V12diff, V34cm, V34diff = [ -6.5123, 0.0245, 23.721, 57.1]

    uMotion after optimization:
    A 12%, B 12%, C 12%

    ID19132 is uMotion optimization 4x4x4 grid, 100kphotons
    Optimal voltages (Common mode, diff): [2.45432158e-02 2.37215216e+01 5.66976610e+01]

    slightly improved C-counts with alignment
    initial uMotion 200kphotons: A 12% (35s), B 11% (28s), C 21% (36s)

    To OD1 filter for 422nm
    B-beam was poorly aligned, counts rise from 500 to 3000 with alignment
    Camera image not focused correctly, not aligned to PMT
    09:23 ion! (422nm shutter was in wring position partly during loading..)
    09:13Z load with 1.8A. 422nm full power, all beams, detuning left at 208 MHz
    empty trap RF OFF/ON

    Move trap-breadboard towards middle of big optical table, to allow for a wall that helps with T-uniformity

    2022-02-11

    Leave trap RF on (but no ion!?), Coolinglaser on, clocklaser on and locked (doors again open today), ASE-sources off, PMT off.

    Notes on EOM PA drive level:
    PA is mini-circ ZHL-5W-1, nominal gain 40 dB (acual ca 45.5 dB from datasheet, +24VDC supply), +37dBm output @ 1 dB compression
    
    Toptica source     Measured -50dB dummyload-monitor   Actual PA output(+50 dB from measurement)   measured gain
    
    -6.2 dBm           -11.5                              +38.5                                       44.7
    -9.2 dBm           -14.2                              +35.8                                       45
    -12.2 dBm          -17                                +33                                         45.2
    -20.2              -24.8                              +25.2                                       45.4
    
    We could use lower-Q resonator (insensitive to temperature?) and higher drive-level?  
    
    Use of slow integrator seems to depend on input-offset trimming - if offset is wrong servo goes to limit very quickly with SLI enabled
    PDH-lock was probably locked to EOM sideband - changing LO-mixer angle or invert/non-invert input on FALC locks to carrier again and get 3 V on cavity photodetector
    if resonator Q is too high, EOM drive phase could be sensitive to temperature changes? Try lowering Q of resonator circuit?
    EOM resonator 480 kHz BW, 25 MHz f0 so Q ca 25/0.48 = ca 50
    PDH-lock works with replaced mini-circ amplifier, but changed PDH-mixer LO angle 109.3 degrees? Now only 0.3V on through-transmission detector - previosly 2-3 V
    PDH-lock again strange, doesn't lock? replace mini-circ amplifier on PDH-photodetector?

    10:07 ion seems lost?
    detuning scan
    09:50 full 422nm power, check lock to Rb-line, 405nm on
    ion dark?

    2022-02-10

    Try to start 180ms clock-run, 23 dB AOM attenuation, g1=1.0, g2=0.01
    From Zeeman AOM attenuation vs. rabi-freq result, if 18 dB attenuation gives 90ms probe, ca 23 dB attenuation gives 180 ms?
    idea for intial convergence of peak-servos: start with expanded probing points at -4*fwhm/2, +4*fwhm/2, and gradually reduce towards -1*fwhm/2, +1*fwhm/2 during first 30 (?) servo-cycles?

    ca17 close cavity door
    better logging from clock-run
    Work on re-ordering probing of zeeman pairs: C1, C4, C2
    While testing PDH-lock, switched to Toptica Channel2 for EOM-source and PDH-mixer
    EOM 25MHz -23.2 dBm, LO Phase 347.1 degrees, Input level 0 dBm
    

    ca 12 uMotion:


    50 Hz peak in FALC ULI / Fine-in1 used for slow Piezo control is ca -121 dBm
    This is ca 300 nV peak, if piezo sensitivity is 500 MHz/V this gives 150 Hz modulation depth which is at least 10-fold too large compared to observed spectrum?
    connect yellow/green PE wire of FALC+Toptica to main PE-wire -> peak rises to 117 dBm??
    Disconnect PE-wire -> back to -122..123 dBm
    -123 dBm also with FALC turned off.

    10:19 stop clock run
    continue clock-run?
    returns with all beams and 405nm
    07:47 ion dark

    2022-02-09

    restart again with B1 estimate.
    restart with comb-linecenter and B4 estimate from C4-
    C- peaks found and locked
    start 90ms clock-run
    g1=1.0, g2=0.01, n_shots=10. servo-cycle length ca 16.2 s

    stop scan during C4+
    OD5/90ms peak-search (still laser drift ongoing but maybe it will settle..)
    C1 peaks:

    C2 peaks:

    C4 peaks:

    Micromotion-scans from past days, voltage-vectors that change only one beam contrast, aka. columns of inv(Cfit)
    (sign is arbitrary and depends on RF-phase vs. artiq-phase - should be constant as long as RF-siggen or artiq not booted?)
    
    Voltage vector components (V12diff, V34cm, V34diff)
    
                VA                         VB                        VC
    ID18981     ( 0.0091, 0.5237, 0.3895)  (2.5971, -1.6778, 1.2260) (2.0105, 0.5066, -3.2377)
    ID18874     (-0.0041, 0.5938, 0.4672)  (3.3642, -1.9986, 1.5413) (2.5628, 0.5140, -3.9767)   (6x6x6 scan)
    ID18865     (-0.0108, 0.5956, 0.4660)  (3.3781, -2.0160, 1.5509) (2.5882, 0.5313, -3.9668)   (6x6x6 scan)
    ID18853     (-0.0121, 0.4117, 0.3734)  (2.4743, -1.5128, 1.0989) (1.9394, 0.3955, -3.1357)
    
    sensitivity to uMotion seems to change slightly (scales all values?) so normalize to e.g. V34diff?
    
                VAn                  VBn                   VCn
    ID18981     ( 0.0234, 1.3445, 1)  (2.1184, -1.3675, 1)  (-0.6210, -0.1565, 1)
    ID18874     (-0.0088, 1.2710, 1)  (2.1827, -1.2967, 1)  (-0.6445, -0.1293, 1)
    ID18865     (-0.0232, 1.2781, 1)  (2.1782, -1.2999, 1)  (-0.6525, -0.1339, 1)
    ID18853     (-0.0324, 1.1026, 1)  (2.2516, -1.3766, 1)  (-0.6185, -0.1261, 1)
    these agree slightly better between measurements?
    
    Use average of these for alternative uMotion optimization search?
                VAn                  VBn                  VCn
    mean        (-0.0102, 1.2491, 1) (2.1827, -1.3354, 1) (-0.6341, -0.1364, 1)
    origininal un-normalized voltage vectors change one contrast by 100%, so use only <10% of these vectors at most for search?
    
    
    16:30 uMotion:

    ca 15:40 close cavity doors

    Siglent settings
    CH1 EOM-drive to Power Amplifier: 25 MHz, 45 mVpp, 0 deg phase
    CH2 PDH-mixer: 25 MHz, 1.7 Vpp, 250 deg phase
    PDH-mixer: no filter on LO/IF port, SHP-20 on photodiode signal to RF-port, IF to FALC inverting input
    
    Beatnote with Siglent driving EOM and PDH-mixer


    ca 0.4 uW in B-beam before chamber
    B-beam counts with OD1 filter: 99.9 (10.4) / 0.5 (0.7) in 10ms
    ion returns after 16s with full 422nm power
    all beams, clearout, 405nm
    15:04 ion went dark
    repeat +C1 peak:
    Now PDH lock with Siglent driving EOM and PDH-mixer - try also Toptica


    OD5/90ms spectra:
    C1 peaks:


    Very loud motor in -1: P.K.304A (poistoilmakone?) - turned it off - now FALC-monitor does not show 45.8 Hz peak anymore!?

    Sanity check of OD0/400us spectra:
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID19003 rabiflop, OD0, on +C1


    FALC-monitor 45.7 Hz peak
    
    minus-k floating      -92 dBm
    minus-k not float     -92 looks about the same...
    
    Toptical FIne-in-1 ARC Factor to -1.0V/V - to keep laser better locked?
    ULI settings 1:on, 4:on -> 100 (V/s)/V

    ID18994 OD5/90ms again
    (laser unlocked..)

    re-start drift-compensation

    ID18990 try again OD5/90ms:
    Fourier limit 9 Hz optical, 4.4 Hz AOM
    C1 peaks:
    ion went dark during C1+ for ca 3 frequencies


    ID18989 OD5/90ms peak-search
    C1 peaks: (user error, this was OD5, but only 9ms probe...)


    uMotion after optimization:

    ID18981 4x4x4 uMotion optimization.
    Optimal voltages (Common mode, diff): [ 0.15641643 24.11483179 56.30722076]

    repeat with modified histogram bins:

    uMotion:

    0753 fluorescence returns (all beams on)
    0736 ion dark.

    2022-02-08

    uMotion after optimization:

    ID18962 is 4x4x4 uMotion optimization
    17:35 uMotion:


    Leave OD5 filter - try 90ms spectra when drift has settled
    cavity doors open - work on softer cable routing tomorrow? to couple less 45Hz to minus-k/cavity

    Try to adjust FALC-settings so two integrators work together:
    SLI   pos 5  240 Hz - 14 kHz  (higher gain at lower freq: from 1.6 dB to 37 dB)
    FLI   pos 7  14 kHz - 170 kHz (higher gain at lower freq: from -6 dB to 15 dB)
    
    Differentiator has biggest effect on servo-bumps
    FLD   pos 7  190 kHz - 1 MHz (gain rises from 0 dB to 15 dB)
    

    ID18954 is rabiflop on +C1, modified PDH-lock settings


    remove 10dB attenuator on FALC output (to ECDL current)
    remove 15 MHz low-pass filter on PDH-mixer IF port (to FALC input) -> slightly higher locking BW, beat-note servo-bumps at +/-740 kHz from carrier

    ID18951 OD0 rabiflop on +C1
    looks reasonable but has some noise in the counts also..


    ID18946 is rabiflop on +C1, OD2, now with FALC-gains that eliminate 45.8Hz peak on cavity transmitted photodiode
    (clock laser unloced at end of run)


    fluorescence returns with OD0/full power of 422nm in a few seconds - no 405nm needed?
    detuning-scan?
    all beams, clearout, 405nm on
    15:27 ion went dark?

    Higher FALC gain giving clear servo-bumps at ca +/-600 kHz in beatnote -> transmitted photodiode 45.8 Hz peak readuced significantly or gone completely.
    Cavity transmitted power photodetecto also shows ca 10 dB peak at 45.8 Hz

    ID18940 is rabiflop on +C1, OD3
    ca 0-1.5 ms looks OK - after that noisy, and settles at incoherent 25% level


    OD3/9ms ??
    C1 peaks:

    C2 peaks:

    C4 peaks:

    menlo comb software crashed again - uptime only 1-2 days?
    Try OD5/90ms spectra with this setup? ID18932
    low and noisy counts with this setup also?
    C1 peaks:

    Switch to Siglent 25 MHz source for EOM and PDH-mixer

    Toptica settings 25 MHz, -27.2 dBm, Phase 0 degrees
    Last ion spectra with Toptica as 25 MHz EOM source and PDH-mixer - try also with Siglent 25 MHz source and DIY mixer?
    SDR tracker with 1ms gate time, fft of frequency shows 50/100Hz peaks


    after optimization:

    Optimal voltages (Common mode, diff): [ 0.13857255 24.07741037 56.29296353]
    ID18931 4x4x4 uMotion optimization
    08 local uMotion:

    2022-02-07

    ID18927 is OD5/90ms peak search, 1Hz step, 300Hz span
    Need to wait for laser drift to settle to see real spectrum
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18923 is OD2/4ms peak search, 20 Hz step, 500 Hz span, 100 Hz AOM Fourier limit
    +C1  75 856 704, h 24.5, FWHM 246 Hz AOM
    -C1  75 829 734, h 19.6, FWHM 303
    C1   75 843 219   B1 4.805624
    
    C1 peaks:


    ID18922 is OD0/400us peak search. 200 Hz step, 5 kHz span, 1 kHz AOM Fourier limit
    comb 75 843 213
    +C1  75 856 721, height 40.7, FWHM 1086 Hz AOM
    -C1  75 829 765, height 40.2, FWHM 1037 Hz AOM
    C1   75 843 243,  B1 4.803319
    +C2  75 870 089, height 48.7, FWHM 976
    -C2  75 816 346, height 45.6, FWHM 999
    C2   75 843 217,  B2 4.803374
    +C4  75 910 455, h 49.1, FWHM 1019
    -C4  75 775 973, h 42.6, FWHM 991
    C4   75 843 214   B4 4.804774
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    16:58 ion again dark? recovers 17:05 with full 422nm power and detuning scan (405nm also tried.)
    16:40 ion went dark, but recovers with all beams, 405nm, and detuning scan..

    Concrete-pillar, with MIKES Z-transducer (green trace). FALC monitor (blue trace)

    MIKES piezo-transducer measurements of 45.7 Hz peak, with FFT-analyzer
    Today fft-analyzed 1 MOhm termination, AC-coupled, 1 Hz RBW
    
                                                          (different input-settings on friday?)
                                      rel.floor           rel.floor(2022-02-05)
    lab floor               -86 dB            0           0
    minus-k top            -102 dB          -16           -20   (minus-k floating but weight on the side pulls top to side)
    stone                  -110 dB          -24           -35
    big opt.table          -110 dB          -24           -37
    concrete below stone   -119 dB          -33                (87.5 Hz, 131 Hz, 175.5 Hz narrow peaks visible also, 400 Hz broad peak)
    
    top of cavity          -125 dB minus-k floating, zensor on top of pumping-T (small breadboard)
    not-floating           -110    (load-adjustment lowered so floating platform at bottom)
    peak rises +10 dB when 'hard' alu-block couples lab-floor to stone!
    
    
    

    Transmitted photodiode shows small 45.7 Hz peak, maybe only 2-3 dB high from noise-floor (50 Hz electrical peak higher)
    with 22 dB attenuation photodiode 3.8 V OK?
    cavity transmitted photodiode 4.8V (saturation) - reduce AOM drive level (was at 20 dB attenuation)

    ID18907 peak-search OD5/90ms - but work in lab will create laser drift..
    FALC-monitor 45-46Hz peaks similar as before, expect supressed carrier and strong sidebands
    C1 peaks: no clear peaks?


    09:30local uMotion:

    2022-02-06

    FALC fast (blue) and slow (green) branch both show 10-15 dB peak at 45.7 Hz


    EOM amplifier also off floor and onto anti-vib?

    Now uMotion:

    Optimal voltages (Common mode, diff): [ 0.16296459 24.16227442 56.19753975]
    ID18892 is once again uMotion optimization 6x6x6 grid, now 100k photons
    ca 15s/point/beam, 6x6x6 x 3beams x 15s = 2.7 hours

    now get: A 12%, B 12.4%, C 12%
    try matlab solution V12cm , V12diff, V34cm, V34diff = [ -6.5123, 0.0869, 23.9509, 56.2829]

    now get: A 10%, B 12%, C 9.3%
    back to latest optimization result V12cm , V12diff, V34cm, V34diff = [ -6.5123, 0.1032, 23.9923, 56.2773]

    get: A 8%, B 11%, C 20%
    Try voltages V12cm , V12diff, V34cm, V34diff = [ -6.5123, 0.0595, 23.9905, 56.8112]

    Try drift-compensation from 1h data.

    wavelength of 45.7 Hz sound ca 7.5 m (+20 C air)
    length of lab is 280+240+160 cm = 680 cm
    width 245+123 = 368 cm
    diagonal sqrt( 680**2 + 268**2 ) = 731 cm

    back to Toptica as source for EOM-signal + mixer - but no big change in 45 Hz component of fast or slow branch signal of FALC
    Signal-level from toptica mixer is probably better than quick DIY-system, with toptica FALC signals don't show 50 Hz (electrical?) components (but with DIY these show depending on settings)

    various tests with PDH-lock, but no clear results.
    fft-analyzer input dc-coupled, 1 MOhm termination
    45.7 Hz peak height     piezo-signal      falc-monitor
    overnight                -107 dBm          -122 dBm         -118 dBm @ 50 Hz
    anti-vib. OFF            -105 dBm          -120
    boa/dbl on table         -105              -117
    

    12:50 fibercomb-control program had crashed again - restart. (comb without autocontrol will loose rep-rate lock if program crashed for too long?)

    Now uMotion:

    Optimal voltages (Common mode, diff): [ 0.10319468 23.99231199 56.27726532]
    ID18874 is 6x6x6 scan again

    Now uMotion:

    alternate ID18865 solution where contrasts <0.1 discarded:
    V12cm , V12diff, V34cm, V34diff = [ -6.5123, 0.0764, 24.0361, 56.86]

    uMotion after optmimization:

    Optimal voltages (Common mode, diff): [ 0.09954851 23.97216324 56.26529901]
    ID18865 uMotion optimization 6x6x6 grid. 40kphotons

    2022-02-05

    Finfoam box or cavity door (partly) closed makes things worse??
    ID18884 all peaks again with settled laser drift. Cavity doors closed, finfoam cover on ECDL
    C1 peaks:

    C2 peaks:

    C4 peaks:


    1905 install finfoam top cover for ECDL
    19local install camera-side cavity door, input-side door not fully closed (seismometer cables etc)

    ID188862 "search3" is OD5/90ms peak search with 300 Hz AOM span - to see all sidebands?
    C1 peaks: (+C1 with ECDL on antivibr-platform, cavity doors open)

    C2 peaks: (closed doors, lots of laser drift)

    C4 peaks: (laser drift settled now?)



    ID188861 "search2" is OD5/90ms peak search with 60 Hz AOM span - to possibly see carrier quickly first?
    C1 peaks:

    C2 peaks:

    C4 peaks:


    16:30 leave lab
    ca 16 local Try anti-vibration platform under ECDL? platform is TS-140 Low Power
    Things to try next week: platform on/off, rubber feet yes/no, finfoam-floor/box yes/no?
    NOTE: previous ECDL was mounted on back table with Finfoam + 4 pcs black rubber (sorbothane?) feet under each corner - (are rubber feet and finfoam here good or bad?)

    uMotion after optmimization:

    ID18853 is new uMotion optimization, 4x4x4 grid, 100kphotons.
    Optimal voltages (Common mode, diff): [ 0.08765912 23.94991012 56.30508654] V12diff quite low?
    ion went dark during scan - recovers after few seconds of 405nm
    ID18845 is uMotion optimization, 4x4x4 grid, 100kphotons.

    14local ion still stored. uMotion ids18842, 43, 44

    2022-02-04

    45.781 Hz peak as measured with MIKES006787 vertical sensor, using FFT-analyzer.
    Lower trace is FALC monitor signal showing clear 20 dB peak at 45.781 Hz (but no other frequencies/peaks?)
    
    sensor position                   peak-height
    top of Cavity, next to pumping-T 
    fitting, minus-k locked           -60.68 dBm     15 dB better than floor
    
    Lab-floor, at cavity output       -42.99 dBm
    
    back table, close to ECDL         -45.7 dBm      3 dB better than floor?
    
    main optical table,
    close to T110 sensor              -80.11 dBm     35 dB better than floor (lowest so far!)
    
    Cavity rack
    above Rpi clock-display           -42.98 dBm     (same as lab floor)
    
    Cavity stone                      -78.0 dBm      close to main optical table (but note new high peak 87.5 Hz -66 dBm!?)
    camera end                        
    
    Minus-K top plate
    camera end, next to tilt-sensor   -63.1 dBm        (why is this 15 dB higher than on stone?)
    
    Repeat top of cavity              -63.3 dBm        (86.5 Hz ca -75 dBm)
    next to pumping-T
    minus-k locked                    
    
    top of cavity-box,                -38.3 dBm      5dB higher than floor  (highest peak is 70.93 Hz, -24 dBm)
    plywood roof
    
    Floating minus-k doesn't seem to have effect on FALC-monitor 45.781 Hz component.
    Z-spring is tuned rougly right, so that minus-k indicator in middle of range.
    
    A/C setting in -3 floor has no great effect on PDH-lock oscillation-peak.
    tried settings 10, 15, 20 Hz, normal 30 Hz - increase to 35 Hz gives maybe a smaller new peak (ca 34 Hz) in PDH-signal?
    

    ID18839 is B-beam detuning scan, 1000ms gate
    Waveplate before B-beam fiber now removed
    PMT counts at 208 MHz detuning ID18840: 8638.2(230.2) / 60.6(11.8) in 1000ms
    smooth curve, no linear bits?


    FALC monitor-output with FFT-analzer
    20dB peak, very sharp, at 45.781 Hz


    MIKES z-axis piezo-transducer on top of cavity, next to pumping-T
    Minus-K not floating with one-sided load (drifts to one sides)
    two or more peaks between 40-50 Hz? 50 Hz probably electrical artefact?


    461nm dispenser seems to still work
    Install timer on AC-socket to avoid prolonged use?


    Slower ULI-setting transfers the 46 Hz noise onto the fast branch of FALC, visible at monitor output
    ULI 1on, 4on os 100 (V/s)/V gives monitor spectrum:


    uMotion after optimization ID18834, 835, 836

    Optimal voltages (Common mode, diff): [ 0.14940904 24.13016108 56.21492309]
    ID18833 is uMotion optimization 4x4x4 grid, 100kphotons
    10:17 uMotion A 3.9%, B 3.4%, C 8.2%

    fluorescence returned at some point.
    08:03 405nm off, no ion
    07:58 405nm on
    all beams on, clearout on
    07:57 ion dark?

    2022-02-03

    peak-search once more:
    C1 peaks:

    C2 peaks:

    C4 peaks:


    sudden jump of clock-laser frequency around 22 local? 674nm frequency jumps down by ca -40 Hz.


    ID18822 peak-search with OD5/90ms - no visible 46 Hz peak in FALC monitor now (slow branch may still show this peak??)
    drift-correction is wrong now - need to wait 2h for good comb data to get best results
    C1 peaks:

    C2 peaks:

    C4 peaks:


    FALC monitor 46 Hz peak was possibly due to overlap between fast and slow branch of servo?
    Change ULI settings to 2-6 OFF: Integrator speed 10 000 (V/s)/V (fastest setting) -> peak disappears
    ULI 2-6 off


    Seems 461nm laser helper-dispenser was probably on (at 5 A) for a long time?? ion was loaded tuesday afternoon, so 48 hours?

    Remove wave plate from trap bread board and re-run OD0 peak-search: C1+ height: 38.1, C1- height: 48.5
    try OD0/400us peak-search: C1+ height 35.8, C1- height 60.4
    further adjust waveplate 283 ?. Extinction ratio now 28 dB???
    674nm B-beam through chamber: 12 dB extinction ratio
    422nm B-beam after cube-polarizer on breadboard (before B-fiber): 25 dB extinction-ratio
    repeat with full power: 17 dB (close to fiber spec?) fiber is marked PM-S405-XP
    422nm extinction-ratio 12...13 dB for B-beam through trap chamber
    OD0/400us peak-search gives C1+ height 36.3, C1- height 60.8
    Waveplate 268 seems to give better PMT-counts with B-beam, independent of heating fiber with hands
    16:29 fluorescence returns
    16:20 ion went dark (while trying full power for polarization measurement?)

    downstairs A/C is still on 30 Hz setting
    FALC monitor-output shows strong peak 45..46 Hz!!

    2nd order sidebands at -93, +95 Hz -> modulation at 46.5...47.5 Hz ??
    ID18801 peak-search again with improved PDH-locking
    low carrier in middle?? 2*(75855805 - 75842318)/5.6120442464394 = 4.80645 ?
    2*(75855828 - 75842318)/5.6120442464394 = 4.8146 ?

    carrier 75 855 820 ? comb linecenter 75 842 331 - compute new B1 from this? -> B1
    2*(75855820 - 75842331)/5.6120442464394 = 4.80716 ?

    FALC PSU is TDK Lambda model DTM65PW150C, +15 VDC 4A, 60W max. Try replacing this also?
    Output connector Kycon KPPx-4P: pins 1+2 -Vout, pins 3+4 +Vout

    ID18798 is OD5/90ms peak-search with modified PDH locking, 25 MHz EOM and PDH-mixer signals from Siglent signal generator. ZAD-6 mixer IF -> BLP-15 low-pass -> FALC-input
    C1 peaks:


    Beat-note doesn't look optimal - requires FALC-tuning..
    Try driving EOM from Siglent signal-generator, and own external mixer for generating PDH error signal?
    -20 dBm drive from Toptica should be ca 50 mVpp into 50 R
    Mixer is ZAD-6, should work with LO from +4 dBm to +10 dBm, i.e. 1Vpp to 2 Vpp.
    15MHz low-pass filter BLP-15 on IF-port, IF-port to FALC input

    PDH-locking singals measured with NI DAQ:
    Note NI DAQ with terminated long cable shows significant 50 Hz and harmonics (improve grounding on DAQ breakout-board)


    ID18792, 793, 794 is uMotion after optimization:
    Optimal voltages (Common mode, diff): [ 0.19370117 24.23808645 56.06816221]

    ID18791 is uMotion optimization, 4x4x4 grid, 100 kphot/pt
    11:17local uMotion: A 9.5%, B 7.1%, C 4.1%

    stop scan of C4 - B-estimate needs improvement..
    07:30local ID18786 is OD5/90ms peak-search
    C1 peaks:

    C2 peaks:

    C4 peaks:

    2022-02-02

    ID18783 is OD5/90ms peak-search
    C1 peaks:

    C2 peaks:

    C4 peaks:

    Increase max drift accepted by driftcomp-program to 50 mHz/s - to make cavity drift settle faster?

    run PMT count monitor while waiting for cavity to settle. B-beam, OD1 filter for 422nm


    predicted x 0.9012 y 0.959
    Z-secular ca 1.954 MHz
    fixed 20 dB attenuator on tickler. B-beam, max attenuation 30 dB, blurry ion at resonance
    all beams ABC, max attenuation 30 dB, still blurry ion at resonance
    x, B-beam, 23 dB
    y, B-beam, 23 dB

    Leave OD5 filter, for 90ms spectra when laser drift settled.

    ID18759 is OD0/400us peak search - check that we find clock-transition as usual after work with cavity and PDH-lock
    +C1  75 855 758, height 35.4, FWHM 849 Hz AOM
    -C1  75 828 619, height 62.5, FWHM 858
    B1   4.835885
    +C2  75 869 230, height 61.1, FWHM 924
    -C2  75 815 250, height 33.0, FWHM 986
    B2   4.824504
    +C4  75 909 719, height 59.9, FWHM 945
    -C4  75 774 671, height 42.4, FWHM 864
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    16:55local close cavity doors
    
    Toptica EOM drive level -21.2 dBm
    max input level 0 dBm
    Lock Level 0 V
    Phase 57.9 degrees
    Piezo ARC signal input Fine-In-1
    Piezo ARC Factor -0.1 V/V (this was lowered from -0.5 V/V previously)
    
    FALC settings
    -Input: mixed down PDH error signal
    +Input: not conencted
    Main output: 20 dB attenuator -> ECDL
    Main-gain trimpot: ca "4"
    
    SLI slow-limited-integrator: pos 6, 110 kHz-6.5 kHz
    Fast-integrator: pos 8, 6.5kHz - 18 kHz
    fast differentiator: pos 7, 190 kHz - 1MHz
    BOA amplifier 500 mA
    1348/674nm doubler temperature +48.8 C
    through-transmitted power-detector 4.5 V
    artiq driftcomp-dds attenuation 20 dB ( amp-a-b-30 power amplifier )
    
    674nm power levels:
    splitter->ion     600 uW
    splitter->cavity  133 uW
    before EOM (corner mirror) 30 uW
    before cavity 21 uW
    PDH detector 17 uW
    
    With these settings SDR-tracked stdev lower than before ca 24-25 Hz for 1348nm = 1.1e-13 (1s gate time)
    
    FALC settings (green trace) move servo-bumps a bit further out to 800 kHz and lower flat noise-floor around carrier?

    Low-pass filter on PDH mixer output seems to be wrong - without low-pass filter servo bumps move to much higher frequencies and we suppress noise close to carrier better!?
    Blue trace has filter removed. 10 dB better suppression of noise close to carrier?

    With more power into the cavity, PDH signals now better?

    Changed back to 70:30 splitter
    Adjusted drift comp AOM.
    Drift comp Urukul DDS attenuation optimization
    
    P_c: optical power after corner mirror towards polarizer/EOM.
    P_cav: optical input power to cavity.
    
    RF attenuation	P_c	P_cav	Ratio
    12 dB:		63 uW	41 uW	0.65
    15 dB:		69 uW	45 uW	0.652
    18 dB:		54 uW	35 uW	0.648
    19 dB:		45 uW	29 uW	0.644
    20 dB:		34 uW	21 uW	0.618
    
    Setting left to 19 dB attenuation.
    
    Initial setting of 12 dB attenuation was too little: amplifier was in compression.
    

    Beamsplitters for split to ion and cavity
    Current is BSX04 90:10 (R:T)
    after isolator  630 uW
    - towards cavity  19 uW  3.2%
    - towards ion    580 uW  (same result as yesterday)
    
    previous was BST04 70:30 (R:T)
    - towards ion      535 uW
    -- ion-breadboard fiber output ca 130 uW
    -- after beam-sampler with servo on 55 uW
    
    - towards cavity   115 uW  17.7% (5.5..6-fold compared to 90:10 splitter)
    

    10:52 Now get uMotion A: 5.1%, B 5.2%, C 6.6%
    Optimal voltages (Common mode, diff): [ 0.18177753 24.19879588 56.13538437]
    ID18748 is uMotion scan 4x4x4 grid, 50 kphotons/histogram
    10:13local uMotion A 9%, B 9.2%, C 10.5%

    Stop scan to work on clock-laser
    ID18743 is peak-search again with OD5/90ms - however lab-temp oscillates and laser-drift remains
    1Hz step 200Hz span, search initialized with comb-linecenter and B4 from OD0 peak-search yesterday
    C1 peaks:

    C2 peaks:

    C4 peaks:

    2022-02-01

    ID18741 is peak-search OD5/90ms. 1Hz step 300Hz span
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18737 uMotion A

    ID18738 uMotion B

    ID18739 uMotion C

    ID18736 uMotion optimization, 4x4x4 grid, 40 kphotons/pt
    Optimal voltages (Common mode, diff): [ 0.16283734 24.18368609 56.14395544]

    uMotion A 8.4%, B 10.9%, C 11.1%

    ID18731 is B-beam detuning scan


    ID18726 is rabiflop on C1+
    ID18727 flop on C1-
    ID18728 flop on C2+
    ID18729 flop on C2-
    ID18730 flop on C4+
    ID18731 flop on C4-
    Compare to earlier OD0 flops? at least EOM-driving power amplifier now fixed.
    C1+ tau_pi 400us, C1- tau_pi 400us
    C2+ , C2-
    C4+ , C4-

    Try new currents: I(-34deg) = [-0.0408 -11.4939 -0.7741] mA
    ID18725 is peak-search with I(-34deg) and 400us probe (OD0)
    C1+  75 861 613  height 38.9  FWHM  987 Hz AOM
    C1-  75 834 530  height 57.9  FWHM 1001 Hz AOM
    B1  4.825780
    C2+  75 875 051  height 59.4  FWHM 923
    C2-  75 821 095  height 35.3  FWHM 913
    B2  4.822419
    C4+  75 915 580  height 57.9  FWHM 978
    C4-  75 780 629  height 37.5  FWHM 903  
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    g
    ID18721 peak-search with 400us probe
    C1+  75 861 588
    C1-  75 834 569
    B1  4.814573
    C2+  75 875 061
    C2-  75 821 105
    B2  4.822407
    C4+  75 915 539
    C4-  75 780 552
    B4  4.822843
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    some shutter problem with past values.
    id18706 is uMotion optimization again
    Optimal voltages (Common mode, diff): [ 0.09022032 24.0129805 56.26878233] - out of scanning range
    these voltages give: A 4.8%, B 5.9%, C 4.7%

    id18702 uMotion optimization in 3x3x3 grid
    Optimal voltages (Common mode, diff): [ 0.13472994 24.16311805 56.08945038] - out of scanning range...

    uMotion A 15.7%, B 17%, C 16.6%

    all beams with OD1 filter 1600/100ms counts
    counts during loading, detuning
    counts to 3600/100ms with detuning at 208 MHz
    13:51:30 ion! looks blurry with all beams, counts 2400/100ms with 190 MHz detuning
    13:49 to 1.80A?
    13:35 try to load again with 1.70A

    13:08 no ion seen in 20mins?
    12:48Z load with 1.70A - power 1.888W initially
    422nm on 'full power' OD0 (no filter)
    check 461nm against dispenser: 196.3 mA

    Cavity power levels 674nm
    fiber from doubler: 810 uW
    after isolator:     610 uW (next is 3/97 splitter to cavity/ion)
    
    split->cavity:          19 uW  (3 %)
    after driftcomp AOM:     7 uW
    
    split->ion:            580 uW  (97 %)
    after fibernoise AOM   320 uW (1st order AOM beam into ion-breadboard fiber)
    ion-breadboard fiber   130 uW (after fiber output)
    power-servo on          58 uW (after power-servo beam-sampler)
    

    Unattenuated 422 powers: A: 5.5 µW, B: 4.2 µW, C: 4.3 µW

    PDH detector measured with Siglent SA
    reasonable agreement with 8k2 transimpedance + 10 V/V postgain model
    test lightsource without modulation corresponds to 160 uW shot-noise
    (20 dB subtracted from measurement results to account for 10x postgain)

    PDH signal with 10 kOhm detector
    ca 10 mVDC -> 1uA photocurrent / 0.4 A/W = 2.5 uW roughly right? Thorlabs power meter shows ca 4 uW at detector
    25MHz signal -52.7 dBm but this is mostly RF pickup and doesn't go away when beam is blocked!!

    Carrier and sidebands on cavity transmitted photodetector, slow piezo-scan of ECDL with external signal generator 0.2 Hz, 1Vpp applied to Toptica fine-1 -0.5 V/V
    peak heights quite variable depending on how much light happens to couple into cavity(?) and how scope samples waveform(?)
    -21.2 dBm EOM drive from Toptica to power-amplifier

    8:10L: Check EOM resonator: Still ok, with minimum at 25.000 MHz (1.18 SWR).
    Toptica 25 MHz phase noise looks as bad as a day before, even after warming-up for a night.

    2022-01-31

    Current 674nm power levels in cavity system
    
    before EOM-polarizer   6 uW
    after polarizer 5.6 uW
    at detector 3.3 uW
    towards cavity 4 uW
    
    only -72 dBm signal at 25 MHz from PDH photodetector? (no mini-circ attached, seems weak?)
    photodetector should have 10 kOhm transimpedance, and 10x postgain.
    DC-output of photodetector ca 120-130 mVDC  compare to power meter 3.3uW * 0.4 A/W * 10 kOhm * 10 V/V = 132 mV - OK?
    
    
    Testing with a similar RF = 10 kOhm photodetector gives 8 mV DC-level (adjustment slightly off, no aftergain). and -46 dBM of 25 MHz!!!

    25MHz AC-signal from photodetector vs. Toptica EOM drive level
    
    EOM     PDH-det
    dBm     dBm
    -30.2   -50.5
    -27.2   -47.5
    -24.2   -44.7
    -21.2   -41.7   (this drive-level seems OK for Ps/Pc = 0.5 roughly from transmitted photodiode signal)
    -18.2   -38.8
    -15.2   -35.8
    -12.2   -32.8
    -9.2    -29.8
    
    Open resonator to check things out. Resistors have heated up, but are still ok. Readjust coil to get 1.18 SWR at 25.0 MHz (min 1.16 at 25.02 MHz).
    6 Ohm resistors a bit hot..
    EOM resonator hot. Cut power to the amplifier, and check impedance matching. SWR minimum now 1.17 at 25.3 MHz. SWR at 25.0 MHz is now 2.21!!

    RF-chain for EOM now: Toptica -12.2 dBm, PA +45 dB -> EOM resonance ciruit. (no attenuators or filters)
    Ps/Pc 0.54 sideband-to-carrier power ratio - seems good!
    carrier 140-35 mV= 105 mV
    sideband 92-35 mV = 57 mV (locked to sideband with inverted gain on FALC - not the best way to measure sideband strength maybe)
    PDH detector DC-output ca 120 mVDC
    PDH AC signal now ca -25 dBm at 25 MHz (was -43..-47 dBm with faulty PA) (both measurements with ZFL-500LN used, 24 dB gain)
    AC signal rises to -20 dBm with -6 dBm drive from Toptica (nominally +38.5 dBm output from PA, close to compression point +39.7 dBm)

    Change to new Mini-Circ amplifier (found in storage room)
    Now get RF-levels from PA at 25 MHz as expected. Maximum output from Toptica should be ca -3 dBm
    RF chain for measurement: Toptica (-12.2 dBm), PA (+45 dB), dummyload (-50 dB) -> predicted -8.2 dB, measured -9 dBm (difference -0.8 dBm)
    With old faulty PA difference between measured and predicted was ca -15 dB

    EOM amplifier (Minicircuits ZHL-5W-1+) seems to output Topt15 dB too little power (tested with dummy load).
    check resonance-circuit 1.14 @ 25.06 MHz. 1.25 @ 25.0 MHz

    2022-01-29

    Phase noise of Toptical 25 MHz driving EOM
    Not that great? 50Hz/100Hz spurs at least +25 dB compared to ARTIQ DDS. 100Hz harmonics at 2, 3, 4, 6, 7, 800 Hz at least.


    EOM resonator tuning not optimal but maybe good enough? SWR 1.14 @ 25.060 MHz
    Compare to 2020-12-11 note: SWR 1.13 @ 25.060 MHz

    2022-01-28

    turn off coolinglaser, clock-laser. trap RF remains on but ion most probably not trapped

    cavity doors open.
    DC-output of detector shows 50 Hz signal - independent of light on/off - but does this affect PDH-lock signal around 25 MHz?
    AC-output of PDH-detector, 3dB splitter to spectrum analyzer and to Toptica/FALC.
    signal seems quite weak -43..-47 dBm. Small sidebands 50/100Hz maybe visible?



    removed laptop and phase-meter from optical table - it was heating reference sensor T110.. DAQ-page showed curoius results.

    0900 still no ion (loading was 2022-01-14 - lost 2022-01-27, stored 13 days)

    second test of Urukul DDS through AA power amplifier. does not show any significant change in AM or PN noise spectrum
    20 dB attenuator used on input of Microsemi phase-meter

    2022-01-27

    sensor for fast magnetic fields: Magcheck-95e

    16:02 still no ion..
    14:11 still no ion..
    13:47 to max power of 422nm
    405nm on 13:42 (kept on for 1-2min several times while trying to find ion)
    13:40 ion dark again? all beams on, clearout on

    Test phase-noise of ARTIQ Urukul DDS with Microsemi 3120A phase-meter. Max frequency for phase-meter is 30 MHz, so here ca 25 MHz output from DDS
    Spurs as expected, 100Hz component is -100 dB down from carrier - should not be a problem!?

    when this signal goes through a power-amplifier - could there be added spurs created??
    we usually drive Zeeman AOM with 18 dB DDS attenuation, corresponding to ca -8.2 dBm output from DDS
    power amplifier is AMPA-B-30, gain >33 dB, power at 1dB compression >30 dBm, 20 MHz to 450 MHz, gain flatness <+/-1dB
    -8.2+33 = +24.8 dBm should be below compression of power amplifier
    measured with siglent SA this is 24.9 dBm at typical Zeeman AOM frequency 75.8 MHz

    11:07 start work in lab - temperature will rise and laser will drift?


    10:55local uMotion A 4.7%, B 5.1%, C 3.0% (IDs 18651...3)

    id18638 try rabiflop on +C1? hope there is little laser-drift in past 3 hours
    average values for peak-centers from end of clock run
    +C1 75 864 198, ID18638
    -C1 75 837 075, ID18647
    +C2 75 877 673, ID18639
    -C2 75 823 599, ID18648
    +C4 75 918 272, ID18640 - ion went dark during run... ID18646 is new run after fluorescence came back
    -C4 75 783 001, ID18649
    +C1 tau_pi maybe 65ms?
    -C1
    +C2 max at maybe 70ms?
    +C4 (fail)
    +C4 (again)
    -C4

    id18637 peak-search, 90ms, 1Hz step, 120Hz span, 120 pts/peak
    measurement time is 10457 s = 2.9 hours!
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ca07:15 local stop clock run, measure spectra

    Images/results form 8h of 90ms clock-run
    C+/C- balance now worse than before?

    2022-01-26

    21:37 local, re-start clock-run with doubled gain g1=0.6 (still g2=0)
    Shuold reach QPN-line sooner?
    ca 21:30local notes from 4 hours of clock-run with g1=0.3
    
    ion vs. AHM3-model: +1.153 (56.541) Hz, N=14372
    EQS = [-0.01 (0.43), -0.02 (0.42), +0.03 (0.45) ] N=888   (add more digits to figure)
    EQS-slpe 0.004 (0.131) Hz/mj-prime-squared, N=888
    1e6*(g-1.6680) = 78.057 (9.196), N=888
    Dsum 0.048 (2.509) Hz, N=888
    S-balance -0.086, +0.137, +0.230
    B = [4.832919 (97.254pT) , 4.832903 (63.812pT) , 4.832907 (44.924pT) ]
    B1 slope +30.015 pT/h
    with g1=0.3 we reach 1/sqrt(tau) QPN-line at tau= 7..10 * T_servocycle
    T_servocycle = 16.2s
    T_deadtime = 45.111 ms
    
    Try changing to g1=1.0 at some point - to see what happens. We should then reach QPN 1/sqrt(tau) line already at 1 servo-cycle?

    17:10 C4-servo lost lock? re-start clock-run

    NOTE A/C adjustments by Repo et al. afternoon 2022-01-25. VFD is now on 30 Hz setting, but tuning improved to avoid oscillation?
    visits and working in lab seems to cause temperature rise (+0.6C with two persons maybe ca 1 hour work before lunch? ) but seems to settle back down without oscillation
    (plot a few ambient sensors in one figure on ion/comb page also?)


    With updated B1 value - try to start 90ms clock-run again.
    g1=0.3, g2=0, n_shots=10. Fourier limit 0.8/90ms = 8.9 Hz
    comb linecenter 75 850 642
    +C1 estimate 75 864 158 (this is -50 Hz or -100 Hz side-band?, not carrier!?) 
    -50Hz SB maybe 75 864 180 -> carrier at 75 864 205 ?
    -C1 estimate 75 837 127
    
    B1 4.833 uT guess?
    gives peak-positions: [ 75.864204  75.837081  75.877679 75.823605  75.918278  75.783006 ]
    
    comb-linecenter - found+C1 = 850624 - 864204 = 13580 Hz AOM
    2*13.580 kHz / 5.6120442464394 kHz/uT = 4.83959 uT
    
    from this we predict -C1 at 75 837 062 ?
    measurement shows ca -C1 at 75 837 078    (meanwhile comb-linecenter drifted to 75 850 639, so -3 Hz AOM)
    this gives 4.83282 uT (?)
    
    New peak-search with B1 guess 4.833 uT??

    C1 peaks: (initialize search with comb-linencenter and B1 4.81624 uT)

    Difficult to get clock-run going? meausre C1 spectra to get better B1?
    C1-pair maybe locked? re-start with B1 estimate 4.81644 uT
    bad C2 counts and no C4 counts
    Try to start 90ms clock-run?
    servo-cycle ca 16.2s with 10 shots. dead-time 45.11ms as usual

    grounding of siglent PSUs?
    (large laser-drift as drift-data from last 2 hours ok now?)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    with cavity-system +/-24V middle wire grounded to cavity rack
    grounding middle wire raises current on one of the 24VDC supplies slightly
    +C1
    Try grounding middle black wire to ground?
    Check cavity-system +/-24VDC Kikusui supplies. Here it seems 0V (black middle wire) is +0.2 V above ground?

    10:47local check uMotion: A 2.9%, B 3.2%, C 3 %


    Update the AHM3 model from atomic bulletin 2022-01
    linear frequency drift maser model
    y_maser = y0 + y1 * (MJD - MJD0)
    
    # ab2021-12 AHM3-model
    MJD0 = 59544
    y0 = 4.47416e-13
    y1 = -2.64742e-15
    
    # ab 2022-01 AHM3
    MJD0 = 59579    
    y0 = 3.57713e-13
    y1 = -2.59867e-15
    
    2022-01-26 is MJD 59605
    
    difference between old and new model)
    AHM3( 59605, ab2021-12 ) = +2.8592338 e-13
    AHM3( 59605, ab2022-01 ) = +2.9014758 e-13
    difference +4.22e-15  (new model predicts higher frequency for AHM3 - corresponding to ca 2 days of daily maser drift)
    
    add ferrite choke to FALC input?
    re-locking laser causes problems with drift-computation and drift-compensation? this causes carrier peak to be lower and noisier?
    +C1

    stop scan at start of -C2
    cavity drift settles slowly after box open and working with pdh-detector..
    start ca 06local id18588 peak-search od5/90ms, 1 Hz AOM step, 300 Hz span, 301pts/peaks
    C1 peaks:

    C2 peaks:

    C4 peaks:

    2022-01-25

    expect 90ms tau_pi?
    Leave OD5 filter - check spectra later when laser-drift low?

    OD0 C1+ Rabiflop:


    Remove attenuators from FALC output. Set integrator gains according to FALC manual.
    FALC monitor was oscillating significantly (over full range +-2V) during PDH lock.

    Notes on coil-currents and B-direction:
    
    current values had a typo, intention was I = [0.1390 -11.4598 -0.8224]
    actually used values were I = [0.1319 -11.4598 -0.8224]
    
    based on last clock-run, new suggested I(-34deg) = [-0.0408 -11.4939 -0.7741] mA
    

    17:15 close cavity door
    13:36local open cavity door to look at PDH signals more..

    ID18567 is DDS attenuation 22 dB

    ID18565 change DDS attenuation to 20 dB

    ID18564 is rabiflop to 9ms

    ID18563 is OD0 rabiflop on +C1, after uMotion optimization


    ID18555 is NPL umotion minimization 4x4x4, 50kPh
    Optimal voltages (Common mode, diff): [ 0.24307518 24.43305244 55.99184995], V12cm = -6.5123 V
    Now get A 1.3 (0.4) %, B 0.4 (0.4) %, C 0.8 (0.4) %


    uMotions: A: 10.6%, B: 11.8%, C: 7.3% at [-6.5123, 0.1795, 24.2356, 55.8717]

    ID18545 is OD0 rabiflop after PDH-adjustments


    CH1 (yellow): Input signal to FALC-servo (this is the downmixed PDH error signal)
    CH2 (blue): piezo scanning-voltage from signal-generator, to Fine-1 input of DLCpro
    CH3 (magenta): Cavity through-transmission photodiode
    Toptical Windows program PDH Channel1: EOM signal level +0.8 dBm, 25 MHz. Input-level -10dB, Lock Level 0.0V, phase 83.6 degrees (was 57.9 degrees)


    ID18543 is flop on +C1, with OD0 and 18 dB - after better mounting for SRS-shutter

    ID18532 is flop on +C1, with 20 dB DDS attenuation (usual setting is 18 dB)
    jump in counts at 6ms?

    ID18530 is flop on +C1, to 30ms (similar to flops measured yesterday)

    ID18529 is flop on +C1 with OD2 - again bump after 6ms and then noisy counts?


    10local turn down A/C to 25 Hz

    ID18528 is flop on +C1 with OD2. tau_pi maybe 2.5ms - bump after 6ms?


    counts 85.1(9.5) / 0.4(0.6) - with lab back lights on
    PMT counts OK when rubber-mounted shutter realigned (needs better mounting - but soft enough so vibrations not visible on comb beat-note)

    06:57local to voltages for zero EQS? Change along vector [1.0, 0, 1.3, 0]
    new V = [-6.5123 , 0.1795037814604 , 24.2356292033976 , 55.8716815575049]
    Notes from 8 hours of 30ms clock run (before change to zero-EQS voltages)
    
    - ion vs. maser-model -0.763 (66.361/sqrt(28788)) Hz
    - quantum-jump counts reasonable [ C1- 18.8/18.7, C1+ 20.3/20.1, C2- 18.5/18.6, C2+ 17.7/17.8, C4- 19.3/19.3, C4+ 18.4/18.4  ] n_servocycles ca 3196
    - EQS = [-0.69 (0.69), -0.18 (0.70), +0.87 (0.70) ] Hz, n_servocycles=3195, EQS-slope +0.262 (0.199) Hz/mj-prime-squared
    - Dsum -0.856 (4.253) Hz, N=3196  (compare to Fourier limit 0.8/30ms = 26.7 Hz) 
    - B1 4.815300(165) uT, slope during 8h -2.5 pT/h
    - B2 4.815299(87) uT
    - B4 4.815284(45) uT
    - g-ratio: 1e6*(g-1.6680) = 76.994 (17.151), N=3195
    - S-balance = [+0.038 (0.109), -0.023 (0.109), -0.024 (0.105) ], N=3195
    - servocycle-length 9.013 s (10 shots per side, 6 peaks)
    - dead-time 45.111 ms
    - self-comparison ADEV reaches 1/sqrt(tau) quite slowly, around tau=300s (due to low gain g1=0.1 ?)
    

    2022-01-24

    (date corrected from 2022-01-23)
    21:35 local switch to g1=0.1, g2=0
    try to start 30ms clock-run? g1=0.2, g2=0.00025
    8.994s servo cycle, n_shots=10, t_dead=45.11 ms
    ID18519 is flop on +C1
    30ms seems reasonable
    now get only 25-30 counts in 10ms (temperature swings cause polarization changes?)
    stop peak-search during +C4
    (18:40local leave lab)
    ID18515 is OD4 / 30 ms peak-search
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18514 again OD2 flop on +C1


    Turn off RF-monitoring Rigol oscilloscope

    ID18512 again OD2 flop on +C1
    better?


    Turn off dispenser Siglent PSU

    ID18510 again OD2 flop on +C1


    Lower FALC-gain from trimpot - this makes the TEM00 on camera slightly larger/brighter, and slight rise in through-cavity photodiode voltage

    ID18509 again OD2 flop on +C1


    Disconnect Thorlabs shutters for A+C beams (they emit acoustic noise?)

    ID18507 repeat OD2 flop on +C1


    disconnect also 405nm psu (AC/DC wallplug-type..)

    ID18505 is OD2 flop on +C1
    visually tau_pi ca 3ms


    ID18502 is OD2/4ms peak-search
    Fourier limit 0.8/4ms = 200 Hz optical, 100Hz AOM
           peak-center     FWHM      height
    +C1    75 864 033      123 Hz    33.6  (Gaussian doesn't fit sidebands)
    -C1    75 837 004      115 Hz    30.5    
    C1 linecenter 75 850 518 
    B1  4.816312 
    comb lincecenter ca 75 850 538
    +C2    75 877 451      124 Hz    34.4   (left of scan center)
    -C2    75 823 580      120 Hz    37.9   (right of scan center, initial B-estimate too high?)
    C2 Linecenter =  75 850 516
    B2 estimate =  4.814839
    +C4    75 917 906      155 Hz    26.4   (left of scan center)
    -C4    75 783 123      111 Hz    42.6(!)
    C4 Linecenter =  75 850 515
    B4 estimate =  4.815529
    comb linecenter 75 850 536 at end of run
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18501 is OD0/400us peak-search
    C1 peaks:

    C2 peaks:

    C4 peaks:

    ID18500 is flop on +C1 after PDH-lock tests and adjustments


    PDH
    lo
    angle   Transmitted photodiode (V)
    90.0	0.62
    83.6	0.65
    77.1	0.66
    70.7	0.66
    64.3	0.71
    57.9	0.7
    51.4	0.74
    45.0	0.80
    38.6	0.89
    25.7	1.00
    
    current PDH settings seem roughly right?
    FALC input-signal seems to be best way to look at PDH LO angle-adjustment, maybe also EOM drive level
    PDH LO level -0.2 dBm from Toptica program initially

    Toptica PDH lo angle was 57.9 degrees

    ID18498 is rabiflop on +C1 - looks similar to ID18494?


    now max 674nm power at 49.4C set-point - ca 90 uW
    tune temperature, get ca 87 uW 674nm before ZeemanAOM cube
    PDH lock now locks less well to different TEM-modes? temperature change of doubler gives less 674nm?
    replaced doubler temp-controller with +5VDC from Siglent PSU
    1348nm->674nm doubler temperature controller also has low-quality AC/DC supply +5VDC/3A - replace?
    temperature setting was 48.2C

    ID18494 rabiflop on C1+ (better oscillations with replaced PSU?)


    674nm power detector on ion-breadboard, replace Thorlabs +/-12V with Siglent +/-12V supply

    ID18491 rabiflop on C1+


    from peak-height of gaussian fits to OD0/400us spectra
    Waveplate  S1       S2       S4       C1_mu_h  C2_mu_h  C4_mu_h
    270        0.104    0.023   -0.049    50.31    49.04    51.62
    271        0.060   -0.081    0.002    46.57    47.11    45.86
    272        0.008   -0.044   -0.032    51.81    46.75    47.74
    
    ID18490 is peak-search with waveplate at 272
        +height   -height
    C1  52.21     51.40     0.008
    C2  44.71     48.78    -0.044
    C4  46.23     49.24    -0.032
    centers [75.86404701 75.8369897  75.87744034 75.82361354 75.91790339 75.78314154] 
    FWHMs   [ 0.00100361  0.00081457  0.00094873  0.000995    0.00099193  0.0009882 ]
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    ID18489 is peak-search with waveplate at 270
        +height   -height   S
    C1  55.54     45.07      0.104
    C2  50.17     47.90      0.023
    C4  49.08     54.16     -0.049
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18488 is +C1 flop to 6 ms
    (sawtooth at end?)

    ID18487 is rabiflop on +C1, to 3ms

    ID18486 is od0/400us peak search
    Waveplate position from last tests 271
        +height  -height  S
    C1  49.37    43.77    +0.060
    C2  43.30    50.91    -0.081
    C4  45.77    45.94    +0.002 (?)
    (results from clock-run were ca S = [+0.15  -0.15 -0.15] form probing at +/-HWHM points)
    
    centers and FWHMs from Gaussian fits
    [75.8639922  75.83695904 75.87743174 75.82352362 75.91789901 75.78312343] [0.00100536 0.0008917  0.00090837 0.00098247 0.00095103 0.00105696]
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18477 is uMotion optimization, 4x4x4 grid, 50 kphotons/point
    Optimal voltages (Common mode, diff): [ 0.17950378 24.8366192 55.87168156]
    now we get uMotion 200kphot: A 2.5(0.4) %, B 2.4(0.3) %, C 2.2(0.4) %


    RID18473 is detuning scan with B-beam, 1000ms gate
    counts with B-beam at 208 MHz: 8855.9(100.8) / 71.8(7.6)

    10 ms gate

    uMotion 100kphot: A 7.5%, B 7.5%, C 6.9%


    with 422nm/OD1: 90.7(9.4) / 0.8(0.9) in 10ms
    with full power B-counts 368(20) / 9.7(2.8) in 10 ms
    ca 10 local ion fluorescence recovers after 1-2min of all beams full power and 405nm laser on
    optimized Rb-lock, 422nm power as before

    2022-01-22

    Rb-lock was unlocked - tried to fix but lost Rb-signal altogether - needs fixing... turn off 422nm laser. leave clock-laser/comb and trap RF on
    1s counts bright 25.8(6.2), dark 20.5(4.5) - maybe ion still trapped? repeat 21.8(4.3) and 22.6(4.5) so maybe not more photons with repumper on?
    16:49 still dark. Rb-lock??
    clearout on, 405nm on
    16:26 ion dark, stop peak search.
    id18431 "search7" with 20 dB
    C1 peaks:

    C2 peaks:

    C4 peaks:

    id18430 "search6" is 40ms 18dB peak search - but we now have laser-drift again..
    C1 peaks:

    C2 peaks:

    C4 peaks:

    id18429 is +C1 rabiflop, 18dB AOM att


    ca12:55 local turn down A/C to 30Hz - it was on 45 Hz
    Vaisala temp-sernsore just after change of A/C:


    id18417 - id18427 are photon time-stamps, 600k each, for later analysis.

    id18416 is with 24 dB att.

    id18415 is +C1 flop, 20dB attenuation

    id18413 600kphot saved stamps
    id18412 uMotion 300kphot, saved timestamps for later analysis. uMotion 5.9%
    id18411 24 dB attenuation

    id18410 is same peak with 22dB attenuation, 40ms probe
    C1 peaks:

    id18409 is 40ms peak-search with 20dB attenuation
    C1 peaks:

    id18408 +C1 flop with 20dB AOM attenuation, should halve power, 1/sqrt(2) * tau_pi?


    ID18406 is peak-search with 40ms, OD4/18dB att.
    220 Hz AOM span to see all sidebands?
    (daatabase tag wrong, overlaps with tag used yesterday, so these plots show two searches..)
    C1 peaks: 'best sidebands ever', carrier lower than 1st sideband, 3rd sideband visible..


    ID18405 is same flop, but all points between 0 and 50ms
    less noise between 20ms and 30ms?

    ID18404 is rabiflop on +C1
    OD0 gives ca 500us tau_pi, so expect 100-fold 50ms with OD4
    low counts at 1/50Hz=20ms? noisy counts at peak-excitation around 50ms


    after activities end in lab ca 18local it takes until ca 22local before laser drift-compensation settles

    ca09_15local B-beam uMotion 6.6% (id18402)
    B-beam PMT counts 100(9.9) / 0.3(0.5) in 10ms (all three shields and lab lights off give low dark counts)

    2022-01-21

    ID18400 peak-search od4/50ms, 18 dB DDS att.
    C1 peaks:

    C2 peaks:

    C4 peaks:



    Note Thorlabs shutters now off - can't use A or C beam!
    switch to OD4 and try this once lab-temp/laser-drift settled?
    OD4 flop 18396


    ID18393 is longer rabiflop with OD1

    stop clock-run, try later/tomorrow with OD5 instead?
    Start 1300us clock-run
    g1=0.2, g2=0.00025, t_probe=1.3 ms, t_dead=45.11 ms, t_servocycle=5.57s, n_shots=10
    Initial coil-currents: I1 = 0.1319, I2 = -11.4598, I3 = -0.8224
    Initial Vbias:  [ -6.05  0.24589062 25.0666472 55.88396938 ]
    EQS-changing Vbias vector [1.0 0.0 1.3037 0.0]
    zero-EQS estimate V12cm -6.5123
    
    Things to try during clock-run:
    - zero EQS voltages?
    - change coil-currents slightly to get better C1/C2/C4 balance?
    -- With 1.3ms FWHM is 615 Hz
    -- C4 sensitivity is 28 Hz / nT -> so as not to loose lock, change B by 615/28 = 22 nT at most?
    -- Coil sensitivities [0.2668 0.4124 0.2731] uT/mA -> max delta-I for any coil = [0.08  0.05  0.08] mA
    - rotate B-beam waveplate to get better S-balance
    - sanity-checks for data, computed after each servocycle
    

    Reasonable rabi-oscillations only with OD0 and OD1

    OD0 500 us
    OD1 1.5 ms
    OD2 3 ms
    OD3 ca 9 ms? incoherent 25% exc after peak
    OD4 
    

    shutter on rubber-post on trap-breadboard - improves comb-disturbance maybe slightly?
    more rigid mounting maybe needed - now risk of slight shutter movement that blocks beam completely?

    Shutter on 422nm-breadboard now connected to ARTIQ CH8. Shutter on trap breadboard in manual control
    Try SRS shutter on 422nm-breadboard - to get rid of vibrations near ion/comb?
    Shutter causes glitches in Rb-error-signal..

    ID18370 OD0 Rabi flop with longer time-axis. Operated remotely.


    ID18369 flop, with 3 kHz audio tone output:


    ID18367 flop with at rising part of first oscillation


    ID18366 flop with more points at shorter taus?


    ID18365 is same flop, 100shots per probe-length
    low count at 200us, 1500us,


    ID18364 is same flop, but Thorlabs shutters for A/C beams not powered

    ID18383 is flop on +C1, no changes from usual settings (Thorlabs shutters powered, SRS-shutter for 422nm on breadboard)
    noisy count at 200us and 1300us, 1700us, 2900us

    Try to look at decoherence-problems with OD0 rabiflops?

    ID 18359 is rabiflop on +C1, OD0. Tau_pi maybe 0.4 ms.


    Move aperture closer to mirror. Spot size should be smaller, and selectivity better?

    ID 18349 is rabiflop on +C1, OD0. Tau_pi maybe 0.4 ms.



    Optimizing AOM angle & aperture. No significant reduction in beat-note power.
    Red light blocked: -114.2 dBm (RF feedthrough?)
    Aperture closed: -98 dBm

    Looking at Zeeman AOM transmission purity. Aperture open, 3 orders passing through, beat-note at ~151 MHz (2x f_AOM) at -67 dBM.

    uMotions with new optimal voltages: A: 3%, B: 3.4%, C: 3.2%
    Optimal voltages (Common mode, diff): [ 0.24589062 25.0666472 55.88396938]


    ID18337 is 4x4x4 uMotion minimization - 10 kphotons for speed
    10:24local uMotion A 20%, B 19%, C 19.5% - quite large!?


    ID18333 is OD4 peak search with 30 ms, 2.5 Hz step, 100 Hz span (AOM)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID18331 is OD4 peak search with 40 ms, 2 Hz step, 60 Hz span (AOM)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID 18330 is rabiflop on +C1, OD4. Tau_pi maybe 40 ms.


    Switch to OD4.

    ID18328 is peak search with 120ms, 0.6 Hz step, 20 Hz span (OD4)
    Gaussian fits (Fourier limit 0.8/120ms = 6.7 Hz optical, 3.3 Hz AOM)
    C1+ h=21.54  FWHM 5.23 Hz AOM
    C1- h=29.24  FWHM 5.86 Hz AOM
    C2+ h=28.42  FWHM 7.26 Hz AOM
    C2- h=20.94  FWHM 6.30 Hz AOM
    C4+ h=27.26  FWHM 6.65 Hz AOM
    C4- h=22.54  FWHM 6.52 Hz AOM
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    ID18327 is flop on C1-, less noise? 32 counts around 120ms
    ID18326 is rabiflop on C1+, tau_pi maybe 120ms but flop-peak cut off and noisy counts max 30
    07:26 stop clockrun
    Clock-run notes from 8h plots
    Linecenter vs. AHM3 model -0.327 (54.367/sqrt(27198))    n_samples from comb, 1 sample/s
    EQS C1,2,4 = -0.76 (0.55), -0.17 (0.50), +0.92 (050)  , n_servocycles=1813   EQS-slope +0.278 (0.174) Hz/mjp-sq
    B4 4.820688 uT (u=31 pT)
    B1-B4  -511 (137) pT
    B2-B4  +428 (56) pT
    Dsum -0.013 (2.967) Hz
    g-ratio 1.668077(13)
    S-balance = [ -0.152 (0.142), +0.158 (0.121), +0.191 (0.119) ]
    servo-cycle length 15.013 s
    dead-time 45.111 ms
    Count statists (not very good..)
        R+   B+    R-   B-
    C1  10.0 9.7   13.4 13.2
    C2  18.2 18.0  13.2 13.1
    C4  18.0 17.8  12.3 12.4
    

    2022-01-20

    19:37 restart with comb-linecenter and higher B2 estimate

    C1, C2 now locked OK - C4 linecenter -22.5 Hz ? so one servo locked to 50Hz sideband?
    B1 and B2 estimates agree, B4 is lower
    Dsum ca -50 Hz

    18:42 restart with B4 estimate and line-center from comb
    servos were not locked correctly Dsum rising to 70-80 Hz and B1/B2/B4 differed a lot

    start 80ms clock-run, g1=0.2, g2=0.0002
    servo cycle length 2*6*10 shots 15 s, dead-time 45.1 ms

    Seems like decoherence problems are related to Zeeman aom optics/geometry?? adjusting these a bit and now OD3..4..5 results reasonable again?
    run clock-run until cavity drift settles - then try rabiflops and peak-search spectra?

    Rabiflops on +C1
    OD    tau_pi
    0     0.36 ms   ID 18315
    2     3.2 ms    ID 18316
    3     10 ms     ID 18317
    4     30 ms     ID 18318
    5     80 ms ?   ID 18320 (freq from comb-linecenter, only 21 counts max)
    

    Check peak-height balance with peak-search?
    ID18297 is peak-search with OD1/1.5 ms
    Do we get same C+/C- balance from this as the last clock-run result '271 deg' ?
        h+     h-      S
    C1  42.68  34.73   +0.103 (clockrun +0.109)
    C2  46.16  47.13   -0.010 (clockrun -0.018)
    C4  46.39  39.07   +0.086 (clockrun +0.052)
    
    Waveplate Thorlabs AHWP05M-600
    wavelength  retardation
    422         0.47522
    674         0.5005
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    5-minute mean from clock-run
    S = [ counts(C+) - counts(C-) ] / [ counts(C+) + counts(C-) ]
    waveplate   S1    S2      S4
    274         0.214 -0.165  -0.022 (15min mean)
    272         0.093 -0.082  -0.016
    271         0.109 -0.018  +0.052  (stdevs 0.095, 0.085, 0.086 )
    268 ('vertical')
    260         0.213 -0.040  -0.026
    
    16:06 to 271 deg
    (674nm ion-breadboard servo-electronics removed, no power-stabilization)
    15:22 to 274 (+6 deg)
    15:14 waveplate to 260 (-8 deg from vertical)
    15:04 waveplate to 272 (+4 deg )
    14:59 lower gains g1=0.1, g2=0
    14:55 set to 269 (+1 deg)
    14:50 initial angle for waveplate 268 deg - fast axis on waveplate approximately vertical
    experimental 1..5ms clock run - could we see C+/C- balance from clock-run, and slowly adjust waveplate?
    g1=0.5, g2=0

    Install lambda/2 waveplate in B-beam on breadboard. Use this waveplate to get balanced C+/C- counts?

    fix PMT dark counts with better alu-foil around PMT and camera-cube
    Back to OD1 and peak search with 1.5 ms
    peak heights from gaussian fit, FWHM AOM
         +h     +FWHM   -h    -FWHM    S
    C1   44.4   279 Hz  46.1   256     -0.019
    C2   39.3   260     44.3   271     -0.060
    C4   40.7   266     (something happened during -C4 scan?? too many dark-count photons, ion dark state not detected correctly!)
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    OD2 filter should give ca 10-fold 4-5 ms tau_pi?
    rises to 3-4ms, then noise?


    ID 18271 / 18272 is +C1 flop with OD1 filter
    looks reasonably good? tau_pi as expected 3-fold ca 1.5 ms

    ID 18270 is rabiflop on +C1, OD0 no filter, ca 400-500us tau_pi


    Peak-height asymmetry is now quite big?
    gaussian fits to 500us peak-search
    peak-height in counts / 100 shots
        +     -      S=(b-r)/(b+r) where b/r mean Cn+ and Cn-
    C1  47.6  52.3   -0.048
    C2  50.4  42.8   +0.082
    C4  45.4  43.6   +0.020
    

    ID18268 is 0.5ms peak search. 160Hz step 4000Hz AOM span
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID 18266 rabiflop on +C1, no OD filter, optimized power


    Increase fiber-power servo setpoint 3.0 -> 3.5 V

    ID 18264 rabiflop on +C1, no OD filter, optimized power


    Optimize 674 power to B-beam. 40 -> 450 nW at chamber exit.
    Switch to no OD filter.

    rabiflop looks ok to 3-4 ms - do peak search with 3.3ms probe? ID18257 is 3.3ms peak search. 20Hz step 500Hz AOM span
    C1 peaks:

    C2 peaks:

    C4 peaks:

    ID18257 rabiflop on +C1

    08:12local uMotion A 4.1%, B 4.8%, C 0.1%

    2022-01-19

    ID 18251 Rabi flop with fiber noise + power servos on. Power servo gain adjusted to a low value.


    ID 18250 Rabi flop with fiber noise servo activated


    ID 18249 is rabiflop after uMotion minimization
    fiber-noise from R&S fixed frequency, power-servo off


    ID 18240 run uMotion minization.
    Optimal voltages (Common mode, diff): [ 0.10926128 24.70072518 56.08856022] (outside scan range)
    at this voltage uMotions: A 5.7%, B 6.5%, C 2.8%

    uMotions: A: 19%, B: 14%, C: 19% at [-6.05, 0.2017, 25.0914, 55.8821]
    16:25local A/C back to high setting

    ID 18235 repeat with R&S driving the fiber servo. Power servo off.

    ID 18234 repeat with more points, adjust frequency from comb ( will see large drift as lab temp rises?)

    ID 18232 lowered A/C, power-servo (ion breadboard) off. some oscillation?


    ID 18230 is flop with lowered A/C



    15:53local try lower A/C 25 Hz?

    Try switching back to servo+VCO for fiber-noise?
    ID 18229 is flop with fiber-noise servo + VCO?


    ID 18225: Rabi flop, OD2, 18 dB att.

    ID 18223: Rabi flop, OD2, 18 dB att.


    OD2 filter in place.
    Switch intra-cavity power servo off.


    (Zeeman AOM drive level vs. power meter 2020-04 python script on linux machine)
    What is the effect of Zeeman AOM drive level?
    18 dB attenuation is the usual setting used for now
    ID 18208 is 19 dB
    ID 18209 is 20 dB
    ID 18210 is 21 dB
    ID 18211 is 22 dB
    ID 18212 is 23 dB
    ID 18213 is 24 dB
    ID 18214 is 25 dB
    ID 18215 is 26 dB
    ID 18216 is 27 dB
    ID 18217 is 28 dB












    ID18206 with intra-cavity servo on (should not matter?)

    ID 18205 is same flop R&S frequency offset 10 Hz (should not matter?)


    ID 18204 is flop with R&S fixed frequency and ion-breadboard power-servo on
    (note working in lab now causes laser drift...)


    Replace fiber-noise servo+VCO with fixed frequency from R&S generator, 178.5MHz and +5 dBm. R&S locked to 10 MHz from maser
    ID 18203 is rabiflop with R&S-generator, power-servo still off
    slightly better coherence?


    ID 18202 is rabiflop with power-servo off - this increases 674nm power


    Power-stabilization for 674nm light on ion breadboard
    Out-of-loop measured with Thorlabs Si-photodiode (variable gain) from B-beam on breadboard. Shows 50Hz component but quite weak?



    Fiber-noise error-signal spectrum, with Win-computer and NI DAQ-card:
    50Hz and 150 Hz components

    Peak around 1600Hz is beat-note between VCO (ca 178.5 MHz) and reference signal (artiq DDS 2*278.5 MHz) - includes both VCO instability and the fiber-noise we want to suppress?
    50Hz and 150Hz components

    2022-01-18

    No experiments over night...don't understand decoherence..

    ID 18187 is flop to 100ms
    ID_18184 is same rabiflop with 500 shots for more statistics?

    Seems like we have a lot of decoherence?
    ID_18182 is rabiflop on +C1 with OD0, ca 1000us tau_pi. flop contrast decays to half at maybe 5 ms?


    Try OD3, rabiflop shows risign counts to 40ms, then flat = incoherent??


    Go to OD4 where we expect tau_pi 70ms? run peak-search to see spectrum
    ID18169 is peak-search with OD4/70ms
    1Hz step 50Hz AOM span, initialized with linecenter from comb
    C1 peaks:


    Test new clock-run program with minimized extra dead-time. Get close to 45ms effective dead-time
    Book-keeping etc. database-writes are now in a separate thread - should not contribute to clock-cycle dead time much
    test-run around 16-17 local with OD0/700us (674nm power not as much as before in B-beam), g1=0.2, g2=0.0025
    QPN-limit seems to be easily achieved (clock-laser more stable than ion?) but reaching QPN-limit line takes >10 servo-cycles.
    observations from 700us test-run
    
    - All counts very good C1/2/4 +/- R/B C1=[ (20.6, 20.9), (21.2, 21.2)  ]  C2= [ (21.0, 21.3), (21.2, 21.1)  ] C4 = [ (20.6, 21.3) , (23.0, 22.8) ] (includes short dark-moment also..)
    - Reaching QPN-line takes >10 servo-cycles (higher g1 would make this happen sooner?)
    - Ion data is worse than Maser at around 2-3 servo-cycles already (tau > 10s)
    - 700us gives Fourier limit 1142 Hz
    - D4-2*D2-D1 = -14.276 (212.509) Hz  (can we use abs(Dsum) < FWHM as sanity-check, here FWHM=5*stdev(Dsum) - check if this holds for longer probe length?
    - B-field equivalence from pairs: B1-B4 = -0.28 (8.8) nT and B2-B4 = 0.395 (4.39) nT  - derive sanity check that depends on probe-length from these also?
    - g-ratio sanity check?
    - EQS-slope +0.269 (10.759) Hz/mj-prime-squared, N = 653 - surprisingly similar to runs with long probe?
    - servo-cycle length now very close to programmed pulse-sequence, measured effective dead-time 45.066 ms (check pulse timing, 45 ms programmed?)
    - now using limited sum 2nd-order servo-term, limited to past 25 servo-cycles, can't say much about this from 700us run
    


    Check for 50/100Hz signals:
    - FALC monitor output, no significant 50 Hz component
    - Cavity transmitted power detector, no significant 50 Hz component when power-servo on
    - no large 50Hz component when intra-cavity power servo off
    Fiber-noise servo:
    - error signal shows 50Hz and 150 Hz peaks both when servo on and off?
    - control-signal to VCO is ca 9.8VDC but shows no large 50Hz component
    - 422nm power shows 100Hz component?

    Coolinglaser shutter seems to cause mechanical disturbance, seen as noise on the beat-note spectrum?
    no shutter
    peak-search in progress
    peak-search in progress, waterfall while 10 shots fired spectrum disturbed - short dead-time between bursts shows OK spectrum?
    no shutter, waterfall

    ID_18085 is peak-search with OD2/2.9ms
    C1 peaks:

    C2 peaks:

    C4 peaks:


    check 674nm power?
    ID_18080 is rabiflop on +C1 with OD0 - visually ca 450us or so (same as before?)
    ID_18082 is flop on +C1 with OD2 - expect 10-fold tau_pi? artiq-fit gives 2.9ms
    ID_18084 is OD2 again - to longer taus - drifted off resonance?
    OD0
    OD2

    Temperature now settled higher with A/C at 45Hz?

    ID18077 peak-search OD5/100ms
    C1 peaks:

    C2 peaks:

    C4 peaks:


    Count-statistics and clock-run results do not look that great. Is intra-cavity power-servo limiting power so that ion-breadboard servo not locked?

    2022-01-17

    try g1=0.1, g2=0.0005
    g1=0.1 g2=0.005 seems to oscillate, too high g2?
    try finite sum for Iterm - use previous 25 servo-cycles?
    try higher gain g1=0.4?

    start OD5/100ms clock-run, new artiq-program with less dead-time and constant (?) servo-cycle length?
    g1 = 0.2, g2 = 0.001, n_shots = 10
    Initialize with comb-linecenter and B-value from last clock run

    Test new clock-run program that should have better dead-time and no memory-leak?
    Prome time (ms)     measured dead-time (ms)    measured servo-cycle for 2*6*10 probes (s)
    0.4 ms              46-47 ms                   5.6 - 5.7 s  -> 47.5 ms per shot
    4 ms                46-47 ms                   6.1 s -> 50.8 ms per shot
    40 ms               48 ms                      10.6 s -> 88 ms per shot
    100 ms              48 ms                      17.8 s -> 148.3 ms oer shot
    

    Intra-cavity power stabilization is now on

    Install 10m PM fibre between ECDL and Comb - however 10m is a bit short, order custom longer fibre?

    check if 422nm stays locked to Rb-line?
    ca14:30local set AC up to 45 Hz - check later if 100ms spectra are still ok? and if temperature now more stable?

    Toptica ECDL has been kept locked now by stepping the temperature set-point whenever the FALC slow integrator reaches +2V or -2V
    4V is the limit for FALC, so this could be expanded a bit? Integrator-voltage controls Piezo with -0.5 V/V gain - this could also be increased?
    Now it seems we sit at the +2V or -2V limit mostly?


    uMotion A 10.7%, B 5.8 %, C 4.9%

    Lots of laser drift now... stop clock run.
    Continue 225ms clock-run with lower gains, g1=0.2, g2=0.001

    Preliminary results from previous clock-run
    
    suggested voltages for zero EQS = [-6.5123 0.2017 24.4904 55.8821]
    
    
    B4 = 4.819549 uT  (u=23 pT) 
    EQS-slope = +0.286 ( 0.087/sqrt(N=559) )
    g-ration 1.668078142 (u=0.000005957/sqrt(559))
    Dsum -0.096 Hz (u=1.841 / sqrt(?))
    

    B counts 103.3 / 2.0 in 10ms
    09:34: fluorescence recovered
    09:30 405nm on
    09:28 all beams on, clearout on
    stop spectrum scan.
    09:27 ion dark?

    ID18006 is peak search, OD6/225ms. lineshape should give scale for optimal gain?
    FWHM/3 is optimal gain only if peak-height is 50%?
    0.3 Hz AOM step, 15 Hz AOM span, 51 pts/peak
    C1 peaks:

    C2 peaks:

    C4 peaks:


    07:33local stop clock run. maybe gains were too high, many values oscillate and QPN-limit not reached? gains were g1=0.4 and g2=0.0025

    2022-01-16

    cycle duration ca 43s (ideal 2*6*10*(225ms+50ms) = 33s)
    Try to start 225ms clock run with linecenter from comb and latest B guess.
    latest B guess 4.81928
    Try rabiflop again with B-beam - max at maybe 225ms?
    Ugh, previous spectrum was with C-beam?? (from uMotion measurement)
    Try rabiflop on the one we guess is the carrier.
    Carrier height quite low?
    +C1          75 864 250?? -> B1 =  2*13524 Hz / Z1 = 4.819634 uT ?
    -50Hz SB(?)  75 864 224..6
    -100Hz SB(?) 75 864 203
    linecenter   75 850 726 (while scan ongoing this changed to ..728...727 )
    
    ID17992 is peak-search, intialize with comb linecenter and B = 4.8 uT
    0.4 Hz step to make it go faster (0.27 Hz would give the usual 5 pts/FWHM)
    Try to find at least C1+ and refine B-estimate from this?


    Try new coil-currents with horizontal B-direction, old direction was nearly vertical
    I1 = 0.1319, I2 = -11.4598, I3 = -0.8224
    Previous currents "I(-35deg)" I1 = -2.4623, I2 = -1.668, I3 = -16.5163 - from C4 clock-run B = 4.794473(20) uT (u=20 pT)

    uMotion 100kphot: A 13.4%, B 8.9%, C 7.3%
    B-beam counts now 108 (10) / 2 (1.5) in 10ms
    16:36 fluorescence recovered
    16:35 all beams on, clearout on
    16:30 ion dark?
    300ms clock-run starts after peak-search, maybe ca 14-15 local

    ID17981 is peak search to show carrier-peaks. OD6/300ms, 0.27 Hz AOM step, 20 Hz span, 75 pts/peak
    laser-drift should be low (?), and clock-run peak-centers used to center these scans
    C1 peaks:

    C2 peaks:

    C4 peaks:


    09:23 stop clock run. self-comparison doesn't show much better statistics compared to 100ms or 40ms clock runs? because jump-statistics not very good at 300ms?

    2022-01-15

    17:03 start clock-run, OD6/300ms, g1=0.4, g2=0.0025, n_shots=10. servo-cycle-lenght ca 52s (ideally 2*6*10*(300ms+50ms) = 42s)

    ID17978 is peak-search with OD6/300ms - because why not..
    200 Hz AOM span, 0.27 Hz AOM step, 741 pts/peak. Fourier limit 2.67 Hz optical, 1.33 Hz AOM
    measurement takes ca 100*(300 ms + 50 ms) = 35 s / point. Current laser-drift -5 mHz/s AOM, so 0.175 Hz AOM per point (65% of frequency step!, 13% of FWHM)
    Need to wait for drift compensation to work well (<1 mHz/s AOM ?) before these spectra look OK?


    Even a short visit to the lab to add OD-filter, stop clock-run, and start a peak-search (maybe <5 min?) causes disturbance in air-conditioning and cavity-drift to change
    Vaisala room-temp sensor (at top of clock-laser rack(?)) shows ca +0.2 C change.
    Cavity-drift was stable, and drift-compensated light showed <1 mHz/s AOM for many hours, but then jumps to -5.7 mHz/s AOM suddenly
    AOM-frequency changes by -10 Hz/30min, so -20 Hz optical in 30 min


    ID17977 is peak-search with OD5/100ms
    (right of center peaks compressed, and left of center stretched, due to laser-drift)
    +C1: sidebands not significantly changed by better +5VDC supply to PDH-EOM?
    Fourier-limit 0.8/100ms = 8Hz optical, 4 Hz AOM. AOM-step is 0.8 Hz so 5 steps/FWHM - carrier peak roughly Fourier-limited (some uncompensated laser-drift after visit to lab)


    Menlo comb-control software had crased sometime during the night!?
    This stops plotting etc. programs that ask the menlo-software (over XML-RPC) for the rep-rate frequency
    re-starting comb-control software seems to work - need to manually enable auto-control again

    Memory and load on artiq-machine during clock-run:
    memory use drops when clock-run stopped


    2022-01-14

    18:16 try to start clock-run with OD2/4ms
    Initialize with comb_linecenter and old B-value
    cavity-drift is now quiite high..

    Add Dsum_valid check for the Dsum = D4 - 2*D2 - D1 where Dn is the Zeeman-pair peak separation in optical Hz
    Criterion is abs(Dsum) < 0.8/t_probe - which seems to be valid for the clock-run yesterday?

    Use new Zeeman-shift coefficients based on previous clock-run:
    Z1  5.612044246439378
    Z2  11.188548496368087
    Z4  27.989141239175556
    

    comb-measurement shows linecenter 75.850914 MHz ca 4 kHz lower compared to yesterday

    16:46 ion went dark? Fluorescence recovers after Rb-relock

    PMT counts in 10ms: A 44.8 / 2.4, B 78.6 / 2.2, C 95.5 / 1.8
    uMotion 100kphot A 9.2%, B 4.7%, C 3.6%
    use slot10 OD1 filter for 422nm

    14:37 ion! (beam-config was not optimal, loading time might have been shorter?)
    14:30:30Z load with 1.7A

    Check helres tuning with antenna analyzer (AA-55 Zoom)
    Before tuning VSWR is 1.07 at 14.4245 MHz
    Tune fore resonance at 14.424 MHz

    set Vbias to last uMotion optimum V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.2017, 25.0914, 55.8821]

    16:20 close cavity doors again
    Open cavity door, EOM temperature controller has +5VDC supply (center positive) AC/DC wall-wart - replace with +5VDC from Siglent PSU

    13:25 405nm off - no ion
    13:16 also 405nm on
    13:13local try full power for 422nm and all beams on

    at some point ion went dark? (clock-laser also unlocked)
    trying to fix computer & artiq issues with memorey leak. try latest artiq update?

    9:16L: Stop clock run
    Memory seems to be leaking during clock run.

    2022-01-13

    18:13local re-start once more with line-center from comb and sane B-estimate

    this was not running correctly Dsum (zeemna-shift-sum, not discriminant-sum...) was not close to zero, and B-estimate from C1/2/4 pairs was not consistent.
    18:06 re-start again, with g1=0.2 and g2=0.0005
    After each R/B probing cycle this gives the frequency correction added to the previous peak-center frequency value:
    
    servo_error(R, B):
        if R+B == 0:
            D = 0.0  # avoid divide-by-zero
        else:
            D = ( ( B - R )/( B + R ) )
        Pterm = g1*(FWHM/3.0)*D
        Dsum += D
        Iterm = g2*(FWHM/3.0)*Dsum
        return Pterm + Iterm
    

    17:13 enable 2nd order gain
    17:10 re-initialize, now g=0.5

    17:00 re-initialize again, now smaller gain - since g=1 seemed to be too much? gain now g=0.1
    with this small gain servos are not able to track clock-laser that drifts a lot

    16:53 re-initialize with center/B from C1 (C2/4 was not locked right?)

    16:29 re-initialize again with comb-linecenter, now OD5/100ms
    cycle-time ca 26.4s / 120 shots = 220ms /shot, probe-ratio 45 %

    16:17 Re-initialize with line-center and B from C2 servo

    print:C1 center 75.85415673015876 MHz, B-field 4.781531666067326 uT
    print:C2 center 75.85416877910049 MHz, B-field 4.779032499703762 uT
    print:C4 center 75.8541321613757 MHz, B-field 4.7812504020944635 uT
    C4 is wrong!? (B-field of C1 and C2 probably right - matches OD0/400us peak-search below)
    
    Dsum settles around +100Hz - so some servo(s) not locked right!?
    16:04 switch to OD4/40ms clock-run
    cycle-time ca 18.9s / 120 shots = 157 ms / shot, probe-ratio 40/157 = 25.4%

    15:57 try switching to OD2/4ms clock run
    cycle-time ca 14.2s / 120 shots = 118 ms / shot, probe-ratio 4/118 = 3 %

    15:54local start comb-measurement
    Try to start clock-run with OD4/400us and only 10 shots. cycle-time for 6 peaks * 2 R/B * 10 shots = 120 shots is ca 13s. 108ms/shot - most of this is overhead??

    OD0/400us peak-search
    +C1 Center =  75 867 621 
    -C1 Center =  75 840 789
    C1 Linecenter =  75 854 205
    B1 estimate =  4.781419
    +C2 Center =  75 880 935
    -C2 Center =  75 827 427
    C2 Linecenter =  75 854 181
    B2 estimate =  4.782252
    +C4 Center =  75 921 106
    -C4 Center =  75 787 284
    C4 Linecenter =  75 854195
    B4 estimate =  4.781147 
    

    ID17837 is uMotion optimization 4x4x4 grid, 200kphotons/histogram
    Optimal voltages (Common mode, diff): [ 0.20165608 25.09144063 55.88206965]
    300kphotons: A 2.6%, B 0.9%, C 1.7%
    600kphotons: A 2.3%, B 1.2%, C 1.7% (ca 70-100s per beam)
    Still some problem with very long photon-correlation measurements - now first bin seems overpopulated and sine-fit fails.
    in repeated measurement last bin is overpopulated? check number-type for timestamp - overflow?
    maximum time-stamp seems to be 1e11 (ns) - so duration of measurement 100s ? problems after this, because measurement returns only zeros?




    Fix bug with uMotion detection where sometimes last bin of histogram was incorrect. Bins must be given explicitly to histogram-function, not just number of bins. If just number of bins given, input data ends abruptly at t_mod, but last bin may extend beyond t_mod and becomes underfilled.

    uMotion A 7.9%, B 5.2%, C 6.4%

    2022-01-12

    Leave PMT monitor.



    Back to fiber-noise servo + VCO, compare this to last spectrum?


    relock laser, gain to "4", spectrum again?
    FALC slow integrator output to Fine-in-1 of DLCPro, used with "Analog Remote Control ARC" Piezo gain of -0.5 V/V (use Windows program to set this)
    +C1 now 75 867 129


    ID17802 is scan of +C1 with FALC-gain turned down to "3" - this slightly increases the brightness of the TEM00 mode on the camera, so more light transmitted through cavity?
    laser unlocked?


    ID17799 is peak-search with OD5/100ms. Fiber-noise servo+VCO replaced by fixed-frequency R&S signal source, locked to 10MHz, set at 178.5 MHz and +5 dBm
    servo free-running phase-detector shows oscillations with maybe 1-5s period - so fiber causes 0.2...1 Hz shift for 674nm?
    Noise-spectrum of this signal is rich with various peaks - but not so clearly at 50Hz or 100Hz?
    Initilize peak-search with ca +70 Hz shifted linecenter 75 853 628 - to centralize peaks better after laser drift?
    +C1 75 867 114
    C1 peaks:


    ID17797 is peak-search with OD5/100ms to see level of 50/100Hz sidebands today
    Note FALC gain trimpot at "7" - this could be changed also to see effect on sidebands?
    laser drift causes peaks left of center to be compressed, and right of center to be stretched out? (measuring left->right (or right->left) would compress or stretch all peaks the same?)
    For -C1, reduce FALC gain to "6"
    FALC monitor output does not show any 50 Hz component (?)
    +C1 carrier 75 867 000 (+40Hz from 400us peak-search)
    -C1 carrier 75 840 214 (+60Hz from 400us peak-search)
    +C2 ca 75 880 390 (estimate ends ..323 so +67 Hz from estimate)
    
    C1 peaks:

    +C2 scan interrupted to test fiber-servo off

    ID17795 is peak-search with OD0/400us to find line-center (no comb data now measured..)
    +C1 Center = 75 866 963
    -C1 Center = 75 840 154
    C1 Linecenter = 75 853 558  (ca +400Hz AOM from yesterday)
    B1 estimate =  4.777137
    
    C1 peaks:


    turn the power-servo on
    clock-laser power with power-servo off/on (question from below, is ratio 4.2/3.47 or larger?). at Zeeman AOM power-meter shows 256uW (on) / 164 uW (off) - so photodiode voltage 4.2 V not real, detector saturated!

    ID17789 is uMotion minimization 4x4x4 grid, 100k photons
    Optimal voltages (Common mode, diff): [ 0.20270735 25.11391536 56.15910653]
    now get: A 6.2%, B 5.9%, C 6.4%

    8:30local uMotion A 13%, B 5.5%, C 9.6%

    2022-01-11

    ID17785 is peak-search with OD5/100ms. 50Hz/150Hz peaks of in-loop power servo detector slightly improved with RF-rectifier + DMM off
    comb-measurement not running, so laser drift might be large?
    +C1 measured with power-servo ON
    -C1 power-servo OFF. relative 674nm power rises from 3.4775 V to 4.2135 V - so probe pulse is slightly too long - but should show big changes in 50/100Hz sidebands anyway?
    (check if 4.2 V is saturation of photodiode - or a real measurement?)

    C1 peaks:

    C2 peaks:

    C4 peaks:


    In-loop 674nm power detector PSD measured with Windows computer DAQ + labview program
    power-servo off: no peaks, flat spectrum ca -102 dB
    servo on: rectifier PSU on and close to photodiode-cable. flat level at ca -95 dB. 50Hz peak ca -77dB
    on/improved: rectifier PSU off (or far away from cable) 50 Hz peak -86 dB (reduced by ~10 dB)


    ID17782 is peak-search again, with program outputting 1/2/4 linecenters and B-estimate correctly now?
    make sure the program that sets peak-positions from linecenter and B-estimate also uses the new coefficients!
    +C1 75 866 578
    -C1 75 839 781
    C1 Linecenter =  75 853 179 7
    B1 estimate =  4.774940
    +C2 75 879 871
    -C2 75 826 376
    C2 Linecenter =  75 853 124
    B2 estimate =  4.781101
    +C4 75 920 107
    -C4 75 786 196
    C4 Linecenter =  75 853 151 
    B4 estimate =  4.784308
    comb linecenter at end of search: 75 853 160
    

    Comb linecenter 75 853 156
    ID17779 is OD0/400us peak-search with power-servo on. Initilized with B 4.781518 uT from yesterday (not very precise?)
    +C1 75 866 619 
    -C1 75 839 771
    C1 Linecenter = 75 853 195
    B1  =  4.783925 uT  (this was not using the revised Zeeman-shift coefficients - improve program so it uses the new coefficients!)
    

    ID17772 is 4x4x4 uMotion scan with ranges 0.1...0.3 23...27 50...58 V. 100kphotons/histogram
    Optimal voltages (Common mode, diff): [ 0.17934357 25.1647457 56.70578304]
    A 10.5%, B 7.7%, C 9%

    12local uMotion A 21.5%, B 16.9%, C 14%

    search for 50Hz disturbance of clock-laser.
    intra-cavity power detector seems ok - no large 50 Hz component
    ion-breadboard power-stabilizing detector (in-loop) shows 50 Hz (and odd harmonics 150 Hz etc) component when power servo locked!?
    Try improving power-supply to this detector (and/or servo electronics) - try also spectrum with power-servo off. maybe try also out-of-loop detector to verify?

    2022-01-10

    clock-laser unlocked, run fluorescence monitor instead..

    ID17765 peak search OD5/100ms
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID17763 repeat peak search, OD0/400us
    Initialized with 75.852700 MHz and 4.798 uT
    +C1  75 866 138
    -C1  75 839 323
    C1 Linecenter = 75 852 730 
    B1 estimate =  4.77813 uT
    +C4 75.91967173679576
    -C4 75.78583910623868
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    B-counts now 102.1 / 2.3 in 10ms

    ID17750 is 4x4x4 uMotion optimization
    Optimal voltages (Common mode, diff): [ 0.19533684 25.178081 57.44754843]
    A 4.7%, B 1.3%, C 2%


    At V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 57.0]
    A 10.6%, B 4.8%, C 19.3%

    EMM seems to have changed slightly, but not too much?
    V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 57.45]
    A 17.7%, B 9.8%, C 13.8%

    V34diff   B-contr
    56.45     7.7% (2.02)
    57.0      4% (-0.71)
    57.45     12% (phase -1.18)
    58.45     28.1% (-0.85)
    
    positioning of outer shield is quite sensitive - not to clip either A-beam or B-beam (bigger tolerances for ASE and C? they go better through middle of holes?)
    camera objective alu-foil grounding test. objective alufoil grounded
    install outer shield - outer shield grounded to optical table with clip+lead
    restart artiq with revised beam-names, same as old names, A is horizontal beam, B is used for loading and clock-runs.
    RF-rectifier power-supply on/off seems to raise and lower the 50Hz/150Hz peaks seen in the in-loop power-detector signal
    moving 230VAC wires far away from Photodiode coax seems to lower peaks?

    fits to rabiflops
       A+   A-   R+   R-   Rmean
    C1 0.39 0.38 1060 1056 1058
    C2 0.35 0.42 1101 1061 1081
    C4 0.41 0.44 1252 1173 1213
    
    ID17679 is flop on C1+
    ID17681 is flop on C1-
    ID17683 is flop on C2+
    ID17685 is flop on C2-
    ID17687 is flop on C4+
    ID17689 is flop on C4-


    ID17678 is peak-search again, now with correct beam
    C1+  75 866 153
    C1-  75 839 227
    C1_Linecenter = 75 852 690
    C1_B 4.7980 uT
    C2+  75 879 559
    C2-  75 825 897
    C4+  75 919 823
    C4-  75 785 600
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    ID17676 is peak-search at new suggested I(-35deg) (note this is C beam by accident...)
    I1 = -2.4623, I2 = -1.668, I3 = -16.5163
    C1 peaks:

    C2 peaks:

    C4 peaks:


    ID17674 is spectrum scan with I0 giving zero B
    I0 = [I1 = -4.5801, I2 = -0.3075, I3 = 0.8179 ]

    New coil currents to try I = [ -2.4623 -1.6680 -16.5163 ]
    
    currents that give zero B: I0 = [-4.5801 -0.3075 0.8179]
    
    New Zeeman-sensitivities that give consistent B1, B2, B4 estimates were used since 2021-12-15
    Further revised sensitivities:
    Component   kHz/uT          (2021-12-15)    old NRC values
    C1           5.61195897      5.61204387      5.6124
    C2          11.188804337    11.188552619    11.190
    C3          22.412722269                    22.415
    C4          27.989567640    27.98914811     27.992
    C5          39.213485573                    39.217
    
    C4-2*C2-C1  -4e-9           -0.000001        -0.004
    

    bias-voltage for zero EMM seems to follow double-exponential with 4h and 24h time-constants
    (attempted outer-shield installation on Friday - but could not find ion so went back to inner+mid shield configuration)
    Find some way of grounding extra charges on camera objective after removing&handling??

    10:17local uMotion at V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 57.45]
    A 1.8%, B 9.5%, C 6.8%

    10:11local uMotion A 26%, B 37%, C 19%
    V34diff   A-contr
    58.7      26%
    57.7      5.5% (phase -0.8621)
    57.45     0.7% 
    57.2      5.2 (+2.0)
    56.7      16.3% (+2.1)
    56.2      25.6 (+2.0)
    

    2022-01-09

    Results from fitting:
    B = B0 + A I % uT
    A =   [  0.2668  0.4134  0.2731 ]  % uT/mA  coil-sensitivity
    B0 =  [  1.222   0.1271 -0.2233 ]  % uT     background field
    
    compare to 2021-01-26 results (12 months ago!)
    A = [ 0.2649  0.4196  0.2713 ]
    B0 = [ 1.1847  0.4244  -0.2796 ]
    
    






    27 spectrum scans in queue.
    17:15 first scan done in 2290s = 0.64 hours
    estimate for remaining 26-scans 26*0.64 = 16.5 hours - finish by 10 Monday?

    16local kiky-Sunday: try to measure coil-characterization?
    Comb on
    A-counts now 31.4 / 1.1 in 10ms (lab lights off)
    Set detuning to 208MHz now counts 70.6 / 3.9 in 10ms (back lab lights on)
    uMotion A 3.3%, B 3.5%, C 9.3% at voltages V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 58.7]
    clock-laser on - remove OD-filters, coil characterization with short 400us or so probe ok
    400us peak-search with old coil-currents:
    this gives good enough excitation-levels for peaks to do coil-characterization?
    +C1  75 866 427
    -C1  75 839 037 (separation 27 kHz AOM)
    +C2  75 880 010
    -C2  75 825 402 separation 54.6 kHz AOM
    +C4  75 921 051
    -C4  75 784 370 separation 136 kHz AOM
    now at  I(-35deg) = [-2.1843 -1.3241 -16.8557] mA - gives ca 4.8 uT field
    
    maybe 100 kHz AOM scan range is OK?
    with 200 Hz AOM step this is 501pts/scan - duration ca 2300s per scan.
    
    ID      name                 Coil-currents [mA]
                                 I1        I2        I3
    17602   2022-01-09-coil1     -2.1843   -1.3241   -16.8557
    17605   coil2                0         0         0
    17607   coil3                +2        0         0
    17609   coil4                -2        0         0
    17611   coil5                +4        0         0    done 19:48, all 5 components fit in 100 kHz scan
    17613   coil6                -4        0         0    B close to zero?
    17615   coil7                +6        0         0    done 21:05, C1-C4, not C5
    17617   coil8                -6        0         0
    17619   coil9                0         +2        0
    17621   coil10               0         -2        0
    17623   coil11               0         +4        0
    17625   coil12               0         -4        0
    17627   coil13               0         +6        0
    17629   coil14               0         -6        0
    17631   coil15               0         0         +2
    17633   coil16               0         0         -2
    17635   coil17               0         0         +4
    17637   coil18               0         0         -4
    17639   coil19               0         0         +6
    17641   coil20               0         0         -6
    17643   coil21               +10       0         0
    17645   coil22               -10       0         0
    17647   coil23               0         +10       0
    17649   coil23               0         -10       0
    17651   coil25               0         0         +10      done 08:30?
    17653   coil26               0         0         -10
    17655   coil27               -2.1843   -1.3241   -16.8557
    
    C1 peaks:

    C2 peaks:

    C4 peaks:


    leave A-beam. PMT counts are low because 422-detuning AOM left at ca 200 MHz. counts 29.4 / 1.2 in 10ms
    09:44local leave at V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 59.7]
    here get A 1.3%, B 3.7%, C 3.9%

    V34diff     A contrast
    62.2        38.1%
    61.2        22.7%
    60.2        6.5% (phase -0.29)
    59.2        7.6% (phase +2.47)
    
    09:39 fluorescence restored (7 mins with all beams on + clearout)
    clearout on
    09:32local ion dark? all beams on (was only newB-beam over night)

    2022-01-08

    21:20local optimum now ca V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 62.2] ( ca -3V change on V34diff in 9 hours since 12local)

    V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 65.2]
    gives good minimum, V34diff still +10 V higher compared to previous stable voltages


    ID17532 is 4x4x4 uMotion scan
    suggested optimum V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.252, 25.48, 68.2] A 43%, B 52%, C 42% - not so good minimum..

    ca12 local uMotion:
    V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.2641, 22.5872, 63.0]

    scanning V34diff shows A-minimum ca 63V
    V34diff ca 73V gives sharper image
    11:52 fluorescence restored - lot's of EMM
    leave all beams on
    10:10local ion dark?

    2022-01-07

    17:26local EMM zero roughly at V34diff +91V now - down from 100-102V initially

    Turn off clock-laser and comb

    Use OD1 filter for 422nm

    ID17454 is coarse tickler for z-mode (figure for paper?) 10ms gate

    ID17455 is same tickler but 30ms gate - but we loose all fluorescence at resonance
    ID17459 is repeat 30ms, but more DDS attenuation - see a slight peak upward at resonance
    ID17460 repeat
    ID17461 repeat
    ID17462 is photon-correlation scan, 22 dB attenuation

    ID17463 is photon-correlation scan, 19 dB attenuation

    ID17464 with 17 dB attenuation

    ID17466 coarse tickler, but loose fluorescence at resonance
    ID17472 coarse tickler, all beams, asymmetric resonance, 7 dB att.
    ID17473 coarse tickler, all beams, 8 dB, asymmetric narrower resonance
    ID17474 coarse tickler, all beams, 9 dB, only two points slightly lower fluorescence
    ID17476 more points gives ca 4pts wide resonance

    ID17477 is similar, more points, asymetric resonance


    ID17446 is fourth 4x4x4 scan [0...0.5] [22...24] [100...104]
    double n_photons to 40k
    Optimal voltages (Common mode, diff): [ 0.26413521 22.58716818 100.46592698] - within scan range, so applied by program
    Now uMotion A 9.7%, B 8.6%, C 9.9%


    At V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.2316, 23.0835, 102.3162]
    uMotion A 1.3%, B 22% !??, C 7.9%
    (Note still high 422nm power and non-optimal detuning in use..)


    ID17439 is third uMotion scan 4x4x4 with ranges [0...0.5] [18...20] [99...105]
    Optimal voltages (Common mode, diff): [ 0.2315798 23.08354858 102.31619867] - again outside scan range..

    Try again 4x4x4 uMotion optimization, ranges V12diff= -0.25...+1 V34diff = 100..115
    found Optimal voltages (Common mode, diff): [ 0.27318177 19.10685086 102.01429658] - outside of scan range..

    ID17420 uMotion minimization 4x4x4 points in range: V12diff = [0, 3], V34cm=[24, 26], V34diff=[90, 110]
    (detuning not at 208MHz now - but maybe it doesn't matter for this coarse search?)
    optimal voltage has V34diff as high as +114V...

    Reasonable looking ion-images for all beams found at voltages V12cm , V12diff, V34cm, V34diff = [ -6.05, 1.5, 25.1946, 99.4376]
    uMotion with full power 422nm at this voltage:
    A 59%, B 56%, C 76%


    Try to manually scan voltages to find rough EMM minimum?
    old voltages were [ -6.05, 0.2399, 25.1946, 55.4376] - expect mostly changes in V34diff and possibly V12diff

    install outer shield
    camera focus-micrometer at "7"
    working with camera-objective seems to cause changes in E-field at ion and large EMM?
    Could not find ion-image with outer shield installed, so remove it again and use only inner+mid-shields.
    Now ion-image is blurred with C-beam, and cooling does not work with A or B beam


    11 local work in lab (temperature will change..)

    ID17118 is peak-search to see current level of 50/100 Hz sidebands
    initialized on last clock-run line-center and B
    120 Hz span, 0.8 Hz step, 151 pts/peak
    C1 peaks:

    (stop scan after C1- 50 Hz sideband)

    09:58local stop clock-run

    2022-01-06


    From 8h of clock-run, EQS_slope = -0.045 (0.091/sqrt(146)) Hz/mj-prime-squared
    EQS = (C1, C2, C4) = (+0.12 (0.30), +0.04 (0.29) , -0.15 (0.32) ) Hz, N=146
    Now at V12cm = -6.05 V
    
    2022-01-05 results at V12cm = -6.2635 V were EQS_slope = +0.206 (0.075/sqrt(157))
    
    Linear fit gives zero EQS at V12cm = -6.0884 V (try this for next clock-run?)
    

    19:40local restart clock-run with new 'S-balance' parameter computed & stored - to visualice balance between Cn+ and Cn- peak height
    if Rn+ and Bn+ and the quantum jump count for Cn+ (same for -), then Sn = [ (Rn+ + Bn+) - (Rn- + Bn-) ] / [ (Rn+ + Bn+) + (Rn- + Bn-) ]
    if C- C+ peak-height imbalance is caused by S-state population distribution we expect S1 = -S2 = -S4 (?)

    From 8 hours of clock-run get:
    S1 +0.060 ( 0.116/sqrt(155) = 0.0093)
    S2 -0.065 ( 0.109/sqrt(155) = 0.0088)
    S4 -0.074 ( 0.113/sqrt(155) = 0.0091)
    signs are as expected.
    possibly wave-plate before A-fiber could be used to fine-tune balance (i.e. linear polarization of 422nm beam) ? 
    but data collection is slow - hour(s) per waveplate-setting?
    

    09:00local re-start clock-run with 2nd-order gain set to zero, and increased 1st-order gain to optimal = FWHM/3
    Current run settings:
    P_gain = 1.0 * FWHM/3
    I_gain = 0.0
    
    Previous run settings:
    P_gain = 0.5* FWHM/2
    I_gain = 0.05 * P_gain
    
    Peik2006a uses only 4 shots/RB-side, and k=3e-4...3e-3
    
    

    With oscillating lab-temp 2nd-order peak-servo seems to do more harm than good?
    results show oscillating abs(B) correlated with temperature - not physical/real?

    2022-01-05

    17:08local leave lab
    Leave 100ms clock-run
    Work on computing optical freqs from mean(C1,C2,C4) line-centers

    disconnect 405nm AC/DC powersupply - replace with wires to +12VDC supply powering pulse-stretcher? But 405nm PCB+laser has + as grounded side...

    Note temperature changes in lab from maybe 10:30 local cause large drifts..
    10:22local at start of servo-cycle 294, set new bias voltages for zero EQS, V12cm=-6.05 V
    V12cm , V12diff, V34cm, V34diff = [ -6.05, 0.2399, 25.1946, 55.4376]

    These results at V12cm = -6.2635
    2021-12-16 results show EQS zero-crossing at ca V12cm = -6.05 V
    We are now at [ -6.2635 0.23985782 24.91625425 55.43757527] move along voltage vector dV = 0.2135 * [1 0 1.3037 0]
    New voltages [-6.05 0.23985782 25.1945942 55.43757527 ]

    Now get EQS = [-0.51 (0.23), -0.18 (0.23), +0.69 (0.27)] Hz with N=157 servo-cycles (8 hours), 100ms probe
    Corresponding EQS_slope = 0.206 (0.075) Hz/mj-prime-squared (N=157 servo cycles)
    (NOTE coil currents were changed after 2021-12-16/17 measurements so EQS values not directly comparable - but this should not change zero-crossing V12cm value?)

    2021-12-17 EQS-measurements showed EQS = [C1 C2 C4] = [-0.298 (0.269), -0.040 (0.242), +0.337 (0.262) ] Hz with N=120 servo-cycles, 120ms probe
    10:07local clock-run ongoing, servo index 289

    2022-01-04


    19:25local re-initialize with C1/2 linecenter and B, zero EQS (previous intialization had EQS for V12cm=0!??)
    one peak of C4 servo was probably locked to a 50 Hz sideband??

    Try to start clock-run.
    servo-cycle duration ca 182 seconds
    probing time is 6 peaks * 2 R/B sides * 100 shots * 100 ms probe = 120 seconds. t_probe/t_cycle = 66%, t_dead/t_cycle = 34%. t_dead(effective) = 52 ms
    How does the servo find the peaks and counts settle to half-max of peak?
    n_servo  +C1_R +C1_B -C1_R -C1_B +C2_R +C2_B -C2_B -C2_R +C4_R +C4_B -C4_R -C4_B
    0        6     2     9     7     33    2     4     23    5     28    14    4   (C1 totally lost?)
    1        1     2     30    3     47    3     3     32    1     30    23    1   (found -C1)
    2        3     1     33    3     24    12    4     29    2     28    22    3
    3        17    6     18    12    8     22    4     40    4     32    19    27  (+C1 found?, only -C4 somewhat balanced?)
    4        24    7     12    34    1     38    3     36    7     40    10    17
    5        47    1     10    28    6     26    3     38    6     30    6     14  (+C1 max imbalance)
    6        41    8     3     26    12    16    6     41    8     30    7     25
    7        17    21    13    21    19    32    8     33    10    14    9     23  (somewhat balanced on all peaks? -C2 most off)
    8        23    29    9     20    15    23    24    24    6     21    12    19  ( +C1 ok, -C2 ok)
    9        10    37    11    39    11    36    12    21    24    13    4     26
    and so on...
    servo gains should be checked - doesn't seem to converge very fast (or 2nd order servo takes long time to find true uncompensated drift value?)
    

    +C1 measured three times today: with PDH/EOM transformers, with PMT Siglent supply, with Pulse-Stretcher Siglent supply


    ID17079 is 100ms peak search with also pulse-stretcher powered from Siglent PSU (+12 V, ca 20 mA or less)

    abort scan and try 100ms clock-run for evening/night

    ID17076 is 100ms peak-search with new PMT supply

    stop scan after -C1

    Power PMT from +5 V output of Siglent powersupply (this powers also +24 V for Zeeman AOM RF amplifier)

    PMT +5 VDC
    Pulse-stretcher +12 VDC
    both PMT and pulse-stretcher have simple AC/DC powersupplies - possibly need to change both?
    reconnect camera

    Try to repeat 100ms spectrums, now with transformers on EOM/PDH signals
    ID17069 is 100ms peak-search
    clock-laser and comb only running for 10 minutes before scan - so drift may not be well compensated..
    C1 peaks:
    12:14local stop camera program and disconnect USB3 cable from camera (at camera/ion breadboard) - this might have caused a false 60-70 jump poin in dataset!?
    camera disconnected before any 50/100Hz sideband on -C1?

    C2:

    abort scan at end of +C2 - try improving PMT power-source?

    comb data shows linecenter 75 852 850 (ca -20 Hz from yesterday)
    Install FTB-1-6 RF transformers on cables from Toptica to EOM and from PDH-detector to Toptica

    Clock-laser relocking - simple on/off of the FALC fast-servo locks laser to TEM00 again, but at one mode too low beat-note frequency
    falc-OFF, then Lowering ECDL piezo voltage by ca 1V, then falc-ON, locks to correct TEM00 mode again
    this could maybe be automated? (does FALC have computer control of servo-on/off, or just mechanical switch?)

    check uMotion 300kphot A 1.1%, B ~0% !, C 1.4%


    back to OD1 for 422nm laser
    10:24:20 ion fluorescence recovered!
    to max 422nm power (this helped previously, 422nm also useful for dissociating molecule?)
    all beams ABC on
    10:21 turn on 405nm laser
    10:20local ion now dark? molecule?
    clock-laser jumped to TEM01 mode at some point during +C4 scan

    2022-01-03

    17:05local leave lab

    RID17060 is peak-search with OD3+OD2 / 100 ms probe
    Fourier limit 4 Hz AOM, 8 Hz optical - this should clearly separate Rabi-sidebands from 50/100Hz sidebands
    0.8 Hz AOM step, 150 Hz AOM span, 187 pts/peak
    Comb data shows about 2.4 mHz/s AOM of residual drift - each peak takes maybe 15 min (??) to scan - laser drift is 2.16 Hz AOM during 15 min (should not obscure/distort 50/100Hz sidebands much?)
    C1 peaks:

    C2 peaks:

    C4 peaks:


    Go to OD3+OD2, expect 120ms tau_pi!?
    ID17058 is flop on +C1. visually ca 100-110ms tau_pi



    ID17057 is peak-search with OD3+OD1 / 40 ms probe
    Fourier limit 10 Hz AOM, 20 Hz optical - this should show 50/100 Hz sidebands? (but Rabi-sidebands may overlap..)
    Span is 100 Hz AOM, 200 Hz optical - should show +/- 50 Hz sidebands at least, possibly only one +/- 100 Hz sideband if carrier not well centered
    step is 2 Hz AOM, so 51 points/peak

    Gaussian fit to spectrum fails..

    C1 peaks:

    abort scan and go to OD5!?
    OD3+OD1=OD4 for 674nm - expect ca 40ms tau_pi?
    RID17055 is flop on +C1 - shows approx 40ms tau_pi


    OD2 filter for 674nm - expect ca 4ms tau_pi
    ID17054 is peak-search with OD2/4 ms probe
    +C1 75 866 543 ( -9 Hz from OD0 result)
    -C1 75 839 201 (+7 Hz from OD0 result)
    C1_center 75 852 872 (+3 Hz from OD0 result)
    B1 4.8720215
    comb_center 75 852 871
    abort this scan and go to OD4
    
    C1 peaks:



    ID17053 is peak-search with OD0 / 400us probe. old abs(B)-estimate and latest line-center from comb
    +C1 75 866 552
    -C1 75 839 195
    C1_center 75 852 874  
    B1 4.874725 uT  (old estimate 4.87098 OK?)
    comb_center 75 852 869
    +C2 75 880 073
    -C2 75 825 641
    +C4 75 921 075
    -C4 75 784 690
    
    C1 peaks:

    C2 peaks:

    C4 peaks:

    background/dark counts quite high with back lab-lights on? threshold now 26 / 10 ms
    ID17052 is rabiflop on +C1 - same 400us tau_pi as before

    comb line-center prediction 75 852 865
    start automatic drift-comp
    beat-note ca 15.538 MHz (15.547 MHz on 2021-12-21)
    start comb and clock-laser. clock-laser beat-note was maybe 10 modes too low (ca 11 MHz) - ECDL piezo reduced from 56 V down to ca 49 V to find right TEM00 mode


    Photon collection efficiency bounded 295..305.

    RID17038 is A-beam detuning scan, 1000ms gate
    RID17039 is A-beam counts 9952.6 / 722.1 bright/dark in 1000ms
    RID17041 is B-beam detuning scan
    RID17042 is B-beam counts 10020.7 / 718.5
    RID17044 is C-beam detuning scan
    RID17045 is C-beam counts 7240.3 / 745.6


    RID 17033 is uMotion optimization in 4x4x4 grid - now with fixed shutters!?
    reduce photon count to 25 k per histogram to make it faster
    Optimal voltages (Common mode, diff): [ 0.23985782 24.91625425 55.43757527]
    Now uMotion 300k:
    A 0.92 %
    B 0.5 %
    C 1.1 %
    


    uMotion 300kphotons again with fixed shutters:
    A 11.6% in 42s
    B 7.4% in 41s
    C 12.8% in 37s
    

    back to original voltages V12cm , V12diff, V34cm, V34diff = [ -6.2635, 0.238, 25.0238, 54.9503]
    Shutters were not working properly - abort run. Press and hold "enable" on C-shutter should fix things?
    ID 17019 is 4x4x4 uMotion minimization scan
    Voltage ranges [[0.22 0.26], [24 26], [54 56]] - (more intellingent scan-range would be better, based on current uMotion measured, and known voltage sensitivities from previous scans)

    reduce 422nm power to slot10/OD1

    uMotion with 422nm OD1, 300k photons (these are probably WRONG, because C-shutter was permanently on??)
    A 2.2% in 28.45 s
    B 2.8% in 27.4 s
    C 11.8% in 37.4 s !!
    repeat C 11.7% in 39.8s
    
    100ms counts with 422nm OD1
    A 137.0 / 7.2
    B 152.3 / 7.7
    C 73.8 / 7.3
    

    100ms Counts with 422nm OD0/full power
    A 342 / 11  "newB"
    B 357 / 12  "oldA"
    C 334 / 10 
    

    100ms PMT counts while loading is RID17001

    count rise to ca 4000 / 100ms when ion appears
    13:50:40 ion! loading time 5min 40s
    PMT counts ca 120-170 / 100ms while loading, lab back-lights on (front lights off)
    assume old current-setting for 461nm laser is ok(?)
    13:45local start load with 1.7A dispenser. 422nm laser detuning 208MHz, OD0/full-power, all beams
    empty trap with RF on/off

    Install FTB-1-6 RF transformers on 100MHz and 10MHz cables from AHM3. Installed before distribution-amp inputs, PRI side towards maser, SEC side towards distribution amplifier

    11:00 No sign of ion in 15min - lost?

    422nm OD0 max power
    check if ion still stored? 422nm laser on, all beams, repumper+clearout, 405nm on

    10:45local turn down AC to 30Hz - to stop oscillation?

    Copy this html page to create 2021 archive index_2021
    Remove 2021 content from this main page for clarity


    For 2021 results see index_2021

    For 2020 results see index_2020

    For 2019 results see index_2019